### v1

# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

Abstract.
First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD) have been proven effective in training neural networks. In the setting of over-parameterization, there is a line of work demonstrating that randomly initialized (stochastic) gradient descent converges to a globally optimal solution at a linear convergence rate for the quadratic loss function. However, the learning rate of GD in training two-layer neural networks has a poor dependence on the sample size and the Gram matrix, resulting in a slow training process. In this paper, we show that for the $L^{2}$ regression problems, the learning rate can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$ to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, which implies that GD enjoys a faster convergence rate. Moreover, we further generalize the method for GD in training two-layer Physics-Informed Neural Networks (PINNs), showing a similar improvement for the learning rate. Although the improved learning rate depends mildly on the Gram matrix, we still need to set it small enough in practice due to the agnostic eigenvalues of the Gram matrix. More importantly, the convergence rate relies on the least eigenvalue of the Gram matrix, leading to slow convergence. In this work, we provide the convergence analysis of natural gradient descent (NGD) in training two-layer PINNs. We show that the learning rate can be $\mathcal{O}(1)$ and at this time, the convergence rate is independent of the Gram matrix.
* Corresponding author.
1. Introduction
In recent years, neural networks have achieved remarkable breakthroughs in the fields of image recognition [1], natural language procssing [2], reinforcement learning [3] and so on. Moreover, due to the flexibility and scalability of neural networks, researchers are paying much attention in exloring new methods involving neural networks for handling problems in scientific computing, where one long-standing and essential problem is solving partial differntial equations (PDEs) numerically. Classical numerical methods, such as finite difference, finite volume and finite elements methods, suffer the curse of dimensionality in solving high-dimensional PDEs. Due to this drawback, various methods involving neural networks have been proposed for solving different type PDEs [4, 5, 6, 7, 8]. Among them, the most representative approach is Physics-Informed Neural Networks (PINNs) [5]. In the framework of PINNs, one incorporate PDE constraints into the loss function and train the neural network with it. With the use of automatic differentiation, the neural network can be efficiently trained by first-order or second-order methods.
In the applications of neural networks, one inevitable issue is the selection of the optimization method. First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD), are widely used in optimizing neural networks as they only calculate the gradient and therefore are computationally efficient. Beside the first-order methods, there has been much interest in utilizing second-order optimization methods to speed up the training, not only for the regression problems [9] but also for the problems related to PDEs [4, 5].
For the convergence part of the optimization method, it has been shown that gradient descent algorithm can even achieve zero training loss under the setting of over-parametrization, which refers to a situation where a model has more parameters than necessary to fit the data [10, 11, 12, 13, 14, 15]. These works are based on the idea of neural tangent kernel (NTK), which shows that training multi-layer fully-connected neural networks via gradient descent is equivalent to performing a certain kernel method as the width of every layer goes to infinity. As for the finite width neural networks, with more refined analysis, it can be shown that the parameters are closed to the initialization throughout the entire training process when the width is large enough. It directly leads to the linear convergence for GD. Despite these beautiful convergence results, the learning rate depends on the sample size and the Gram matrix. Thus, it need to be sufficiently small to guarantee the convergence in practice. But doing so results in a slow training process. In contrast to the first-order methods, the second-order methods enjoy fast convergence due to the absence of requirements for the learning rate [16].
1.1. Contributions
The main contributions of our work can are summarized as follows:
  • For the $L^{2}$ regression problems, we demonstrate that the learning rate $\eta$ of gradient descent can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$, as shown in [10], to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the ReLU activation function and the random initialization, $\lambda_{0}$ is the least eigenvalue of $\bm{H}^{\infty}$. [17] has also shown the same improvement of the learning rate, but it requires that $m=\Omega(\frac{n^{6}}{\lambda_{0}^{4}\delta^{3}})$. Moreover in [18], the dependence on $n$, i.e. $\Omega(n^{6})$, is necessary and can not be improved due to the requirement of the proof method. Different from the method in [17], our method comes from a new recursion formula of gradient descent and it can be easily generalized to PINNs. Compared to [17], we only require that $m=\Omega(\frac{n^{4}}{\lambda_{0}^{4}}(\log(\frac{n}{\delta}))^{2})$.
  • For the PINNs, we simultaneously improve both the learning rate $\eta$ of gradient descent and the requirement of the width $m$. The improvements rely on a new recursion formula of gradient descent, which is similar to that for regression problems. Specifically, we can improve the learning rate $\eta=\mathcal{O}(\lambda_{0})$ required in [19] to $\eta=\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$ and the requirement for the width $m$, i.e. $m=\tilde{\Omega}(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}})$, can be improved to $m=\tilde{\Omega}(\frac{1}{\lambda_{0}^{4}}(\log(\frac{n_{1}+n_{2}}{\delta})))$.
  • We provide the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. Compared to gradient descent, the convergence rate of NGD is independent of $\lambda_{0}$, which results in a faster convergence. Meanwhile, we show that the learning rate can be even $\mathcal{O}(1)$. When the activation function is smooth, NGD enjoys quadratic convergence rate.
1.2. Related Works
First-order methods. There are mainly two ways to study the optimization of neural networks and try to understand why first-order methods can find a global minimum. One way is to analyse the optimization landscape [15, 20]. It has been shown that gradient descent can find a global minimum in polynomial time if the optimization landscape has some good geometric properties. However, some unrealistic assumptions in these works make it difficult to generalize to practical neural networks. Another way to understand the optimization of neural networks is to analyze the optimization dynamics of first-order methods. For the two-layer ReLU neural networks, as shown in [10], as long as the width $m$ is large enough and no two inputs are parallel, randomly initialized gradient descent converges to a globally optimal solution at a linear rate. Later, the results have been generalized to deep neural networks with smooth activation functions [11]. Both rely on that the Gram matrices are stable throughout the training process, which leads to the convergence to the global minimum. Besides regression and classification problems, [19] demonstrated the convergence of the gradient descent for two-layer PINNs by the similar analysis of optimization dynamics. However, both [10] and [19] require that the learning rate be sufficiently small to converge. In this work, we conduct a refined analysis of gradient descent for $L^{2}$ regression problems and PINNs, leading to a milder requirement for the learning rate.
Second-order methods. Although second-order methods possess better convergence rate, they are rarely used in training deep neural networks due to the prohibitive computational cost. As a variant of the original Gauss-Newton method, natural gradient descent (NGD) is more efficient in practice. Meanwhile, as shown in [21] and [16], it also enjoys faster convergence rate for the $L^{2}$ regression problems than of gradient descent. In this paper, we provide the convergence analysis for NGD in training two-layer PINNs.
1.3. Notations
We denote $[n]=\{1,2,\cdots,n\}$ for $n\in\mathbb{N}$. Given a set $S$, we denote the uniform distribution on $S$ by $Unif\{S\}$. We use $I\{E\}$ to denote the indicator function of the event $E$. For variables $a$ and $b$, $a\lesssim b$ is equivalent to $a\leq Cb$ for some universal constant $C$. For two functions $f_{1}(n)$ and $f_{2}(n)$, the relation $f_{1}(n)=\mathcal{O}(f_{2}(n))$ is equivalent to $f_{1}(n)\leq Cf_{2}(n)$ for some universal constant $C$. And similarly, $f_{1}(n)=\Omega(f_{2}(n))$ is equivalent to $f_{1}(n)\geq Cf_{2}(n)$. Throughout the paper, we use boldface denote vectors. Given $x_{1},\cdots,x_{d}\in\mathbb{R}$, we use $(x_{1},\cdots,x_{d})$ or $[x_{1},\cdots,x_{d}]$ to denote a row vector with $i$-th component $x_{i}$ for $i\in[d]$ and then $(x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d}$ is a column vector.
1.4. Organization of this Paper
We present the improvements of the learning rate of gradient descent for $L^{2}$ regression problems and PINNs in Section 2 and Section 3 respectively. In Section 4, we show the convergence of natural gradient descent for training PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. We conclude in Section 5 and the detailed proofs are put in the Appendix for readability and brevity.
2. Improved Learning Rate of Gradient Descent for $L^{2}$ Regression Problems
2.1. Problem Setup
In this section, we consider a two-layer neural network $f$ with the following form.
$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\tilde{\bm{x}}),$ (1)
where $\bm{x}\in\mathbb{R}^{d}$ is the input, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$, $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T},\bm{a}=(a_{1},\cdots,a_{m})^{T}$, for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+1}$ is the weight vector of the first layer, $a_{r}\in\mathbb{R}$ is the output weight and $\sigma(\cdot)$ is the ReLU activation function. Different from the setup in [10], the two-layer neural network that we consider has bias term. In the following, we assume that each input vector $\bm{x}\in\mathbb{R}^{d}$ has been augmented to $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$. With a little abuse of notation, we write $\bm{x}$ for $\tilde{\bm{x}}$ and write $f$ as
$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x}).$
For the $L^{2}$ regression problem, given training data set $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, the aim is to minimize the loss function
$L(\bm{w},\bm{a}):=\sum\limits_{i=1}^{n}\frac{1}{2}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})^{2}.$ (2)
Before training, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$ and output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$. In the training process, we fix the second layer and optimize the first layer by gradient descent (GD), i.e., we update the weights by the following formulation.
$\bm{w}(k+1)=\bm{w}(k)-\eta\frac{\partial L(\bm{w}(k))}{\partial\bm{w}},$ (3)
where $\eta>0$ is the learning rate, $k\in\mathbb{N}$ and $L(\bm{w}(k))$ is an abbreviation of $L(\bm{w}(k),\bm{a})$. From the form of $L(\bm{w})$, we can deduce that
$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{m}}\sum\limits_{i=1}^{n}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})\bm{x}_{i}I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}.$ (4)
We denote $u_{i}(k)=f(\bm{x}_{i};\bm{w}(k),\bm{a})$ the prediction on input $\bm{x}_{i}$ at $k$-th iteration and we let $\bm{u}(k)=(u_{1}(k),\cdots,u_{n}(k))^{T}\in\mathbb{R}^{n}$ be the prediction vector at $k$-th iteration. For the labels, we define $\bm{y}:=(y_{1},\cdots,y_{n})^{T}\in\mathbb{R}^{n}$.
According to [10], in the continuous setting, the dynamics of predictions can be written as
$\frac{d}{dt}\bm{u}(t)=\bm{H}(t)(\bm{y}-\bm{u}(t)),$ (5)
where $H(t)\in\mathbb{R}^{n\times n}$ is the Gram matrix at time $t$ with $(i,j)$-entry
$H_{ij}(t)=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}(t)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(t)^{T}\bm{x}_{j}\geq 0\right\}.$ (6)
In the setting of over-parameterization and randomly initialization, [10] has shown that (1) $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(\sqrt{1/m})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the initialization with $(i,j)$-th entry
$H_{ij}^{\infty}=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}^{T}\bm{x}_{i}\geq 0,\bm{w}^{T}\bm{x}_{j}\geq 0\right\}\right]$ (7)
and (2) $\|\bm{H}(t)-\bm{H}(0)\|_{2}=\mathcal{O}(\sqrt{1/m})$ for all $t>0$. Therefore, as $m\to\infty$, the dynamics of the predictions are charactered by $H^{\infty}$, leading to the linear convergence.
2.2. Main Results
To simplify the analysis, we make the following assumptions on the training data.
Assumption 1.
For $i\in[n]$, $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$ and $|y_{i}|\leq 1$, where $\bm{x}_{i}\in\mathbb{R}^{d+1}$ is the augmented input.
Assumption 2.
No two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are parallel, i.e., for any $\bm{x}_{t},\bm{x}_{s}\in\{\bm{x}_{i}\}_{i=1}^{n}$ and any $\alpha\in\mathbb{R}$, we have $\bm{x}_{t}\neq\bm{x}_{s}$.
Since inputs are all augmented, Assumption 2 is equivalent to that no two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are equal, which holds naturally. Under Assumption 2, Theorem 3.1 in [10] implies that $\lambda_{0}:=\lambda_{min}(\bm{H}^{\infty})>0$, which is crucial in the convergence analysis.
As stated before, there are two improtant facts that determine the optimization dynamics, one is that at initialization $\bm{H}(0)$ is closed to $\bm{H}^{\infty}$ and another is that $\bm{H}(k)$ does not go far away from the initialization $\bm{H}(0)$ for all $k\in\mathbb{N}$. These facts are supported by the following two lemmas.
Lemma 1 (Lemma 3.1 in [10]).
If $m=\Omega\left(\frac{n^{2}}{\lambda_{0}^{2}}\log\left(\frac{n}{\delta}\right)\right)$, we have with probability at least $1-\delta$, $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\frac{\lambda_{0}}{4}$ and $\lambda_{min}(\bm{H}(0))\geq\frac{3\lambda_{0}}{4}$.
Lemma 2.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-n^{2}e^{-mR/4}$, the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then the matrix $\bm{H}(\bm{w})\in\mathbb{R}^{n\times n}$ defined by
$H(\bm{w})_{ij}=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\right\}.$
satisfies
$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}<4nR.$ (8)
Remark 1.
Although Lemma 2 looks almost same to Lemma 3.1 in [18], the proof of Lemma 3.1 in [18] lacks an important part, i.e., the proof of the measurability for the related variable. In fact, Lemma 2 ensures us to bound the change of $\bm{H}(\bm{w})$ when $\bm{w}$ in a small ball. However, the set consisting of vectors in the small ball is uncountable, thus the related variable may not be measurable due to the discontinuity of the indicator function appeared in the definition of $\bm{H}(\bm{w})$. In the proof, we borrow the concept of pointwise measurability in the empirical process (see Chapter 8 in [22]) to bridge the gap. Then we can bound the supremum with a new random variable. By applying a concentration inequality to this new random variable, we can reach the conclusion.
Abount the convergence of gradient descent, [10] has demonstrated that if the learning rate $\eta=\mathcal{O}(\lambda_{0}/n^{2})$, then randomly initialized gradient descent converges to a globally optimal solution at a linear convergence rate when $m$ is large enough. The requirement of $\eta$ is from the decomposition for the residual in the $(k+1)$-th iteration, i.e.,
$\bm{y}-\bm{u}(k+1)=\bm{y}-\bm{u}(k)-(\bm{u}(k+1)-\bm{u}(k)).$ (9)
Instead of decomposing the residual into the two terms as above, we write it as follows, which serves as a recursion formula.
Lemma 3.
For all $k\in\mathbb{N}$, we have
$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta H(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k),$ (10)
where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n}(k))^{T}\in\mathbb{R}^{n}$ with $i$-th
$I_{1}^{i}(k):=u_{i}(k+1)-u_{i}(k)-\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$ (11)
After [10], [17] improved the network size $m$ to $m=\mathcal{O}(\lambda_{0}^{-4}n^{4}\log^{3}(n/\delta)\log(m/\delta))$, but it still requires $\eta=\mathcal{O}(\lambda_{0}/n^{2})$. Another work [17] provided a refined analysis over [10], which shows that $\eta=\frac{1}{C_{1}\|\bm{H}^{\infty}\|_{2}}$ is enough for the convergence rate $1-\frac{\lambda_{0}C_{2}}{C_{1}\|\bm{H}{\infty}\|_{2}}$, but the constants $C_{1},C_{2}$ may depend on the parameters $\lambda_{0},n$ and $\delta$. It requires the network size $m$ to be $\Omega(\lambda_{0}^{-4}n^{6}\delta^{-3})$ to make $C_{1},C_{2}$ independent of $\lambda_{0},n$ and $\delta$. In this paper, we improve both the learning rate and network size and the main reslut is the following theorem.
Theorem 1.
Under Assumption 1 and Assumption 2, if we set the number of hidden nodes $m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}}\log\left(\frac{n}{\delta}\right)\right)$ and the learning rate $\eta=\mathcal{O}\left(\frac{1}{\|\bm{H}^{\infty}\|_{2}}\right)$, then with probability at least $1-\delta$ over the random initialization, the gradient descent algorithm satisfies
$\|\bm{y}-\bm{u}(k)\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ (12)
for all $k\in\mathbb{N}$.
Remark 2.
Our proof method requires that $\bm{I}-\eta\bm{H}(k)$ is positive definite for all $k\in\mathbb{N}$. As $\bm{H}(k)$ is closed to $\bm{H}^{\infty}$, the requirement is satisfied if $\eta\|\bm{H}^{\infty}\|_{2}\lesssim 1$. Note that $\|\bm{H}^{\infty}\|_{2}\leq n$, it is sufficient to set $\eta=\mathcal{O}(1/n)$, which results in an improvement of $\mathcal{O}(\lambda_{0}/n)$.
3. Improved Learning Rate of Gradient Descent for Two-Layer Physics-Informed Neural Networks
3.1. Problem Setup
In this section, we consider the same setup as [19], i.e., we focus on the PDE with the following form.
$&\frac{\partial u}{\partial x_{0}}(\bm{x})-\sum\limits_{i=1}^{d}\frac{\partial^{2}u}{\partial x_{i}^{2}}(\bm{x})=f(\bm{x}),\,\bm{x}\in(0,T)\times\Omega,\\ &u(\bm{x})=g(\bm{x}),\ \bm{x}\in\{0\}\times\Omega\cup[0,T]\times\partial\Omega,\end{aligned}\right.$ (13)
where $\bm{x}=(x_{0},x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d+1}$ and $x_{0}\in[0,T]$ is the time variable. In the following, we assume that $\|\bm{x}\|_{2}\leq 1$ for $\bm{x}\in[0,T]\times\bar{\Omega}$ and $f,g$ are bounded continuous functions.
Moreover, we consider a two-layer neural network of the following form.
$\phi(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\tilde{\bm{x}}),$ (14)
where $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}\in\mathbb{R}^{m(d+2)}$, $\bm{a}=(a_{1},\cdots,a_{m})^{T}\in\mathbb{R}^{m}$ and for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+2}$ is the weight vector of the first layer, $a_{r}$ is the output weight and $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function. Similar as that in Section 2, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+2}$ is the augmented vector from $\bm{x}$ and in the following, we write $\bm{x}$ for $\tilde{\bm{x}}$ for brevity.
In the framework of PINNs, given training samples $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ and $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ that are from interior and boundary respectively, we aim to minimize the following empirical loss function.
$L(\bm{w},\bm{a})$ $:=\sum\limits_{p=1}^{n_{1}}\frac{1}{2n_{1}}\left(\frac{\partial\phi}{\partial x_{0}}(\bm{x}_{p};\bm{w},\bm{a})-\sum\limits_{i=1}^{d}\frac{\partial^{2}\phi}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w},\bm{a})-f(\bm{x}_{p})\right)^{2}$ (15)
$\quad+\sum\limits_{j=1}^{n_{2}}\frac{1}{2n_{2}}\left(\phi(\bm{y}_{j};\bm{w},\bm{a})-g(\bm{y}_{j})\right)^{2}.$
Similar as that for the $L^{2}$ regression problems, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$, output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$ and fix the output weights. Then the gradient descent updates the hidden weights by the following formulations:
$\bm{w}_{r}(k+1)$ $=\bm{w}_{r}(k)-\eta\frac{\partial L(\bm{w}(k),\bm{a})}{\partial\bm{w}_{r}}$ (16)
for all $r\in[m]$ and $k\in\mathbb{N}$, where $\eta>0$ is the learning rate. For brevity, we write $L(\bm{w})$ for $L(\bm{w},\bm{a})$.
To simplify the notations, for the residuals of interior and boundary, we denote them by $s_{p}(\bm{w})$ and $h_{j}(\bm{w})$ respectively, i.e.,
$s_{p}(\bm{w})=\frac{1}{\sqrt{n_{1}}}\left(\frac{\partial u}{\partial x_{0}}(\bm{x}_{p};\bm{w})-\sum\limits_{i=1}^{d}\frac{\partial^{2}u}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w})-f(\bm{x}_{p})\right)$ (17)
and
$h_{j}(\bm{w})=\frac{1}{\sqrt{n_{2}}}(u(\bm{y}_{j};\bm{w})-g(\bm{y}_{j})).$ (18)
Then the empirical loss function can be written as
$L(\bm{w})=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right),$ (19)
where
$\bm{s}(\bm{w})=(s_{1}(\bm{w}),\cdots,s_{n_{1}}(\bm{w}))^{T}\in\mathbb{R}^{n_{1}}$ (20)
and
$\bm{h}(\bm{w})=(h_{1}(\bm{w}),\cdots,h_{n_{2}}(\bm{w}))^{T}\in\mathbb{R}^{n_{2}}.$ (21)
At this time, we have
$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\sum\limits_{p=1}^{n_{1}}s_{p}(\bm{w})\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}+\sum\limits_{j=1}^{n_{2}}h_{j}(\bm{w})\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}$ (22)
and the Gram matrix $\bm{H}(\bm{w})$ is defined as
$\bm{H}(\bm{w})=\bm{D}^{T}\bm{D},\ \bm{D}:=\left[\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right].$ (23)
3.2. Main Results
First, we make the following assumptions about the training samples, which are similar to those for the $L^{2}$ regression problems.
Assumption 3.
For $p\in[n_{1}]$ and $j\in[n_{2}]$, $\|\bm{x}_{p}\|_{2}\leq\sqrt{2},\|\bm{y}_{j}\|_{2}\leq\sqrt{2}$, where all inputs have been augmented.
Assumption 4.
No two samples in $\{\bm{x}_{p}\}_{p=1}^{n_{1}}\cup\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ are parallel.
Under Assumption 4, Lemma 3.3 in [19] implies that the Gram matrix $\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\bm{H}(\bm{w})$ is strictly positive definite and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$. Similarly, $\bm{H}^{\infty}$ plays an important role in the optimization process. The following two lemmas indicate that $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(1/\sqrt{m})$ and $\|\bm{H}(k)-\bm{H}(0)\|_{2}=\mathcal{O}(1/\sqrt{m})$, leading to the linear convergence.
Lemma 4.
If $m=\Omega\left(\frac{d^{4}}{\lambda_{0}^{2}}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$, we have that with probability at least $1-\delta$, $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\frac{\lambda_{0}}{4}$ and $\lambda_{min}(\bm{H}(0))\geq\frac{3}{4}\lambda_{0}$.
Lemma 5.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-\delta-(n_{1}+n_{2})^{2}e^{-mR}$, the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
$\|H(\bm{w})-H(0)\|_{F}<CM^{2}R,$ (24)
where $M=2(d+2)\log(2m(d+2)/\delta)$ and $C$ is a universal constant.
In [19], the decomposition for the residual in the $(k+1)$-th iteration is same as the one in [10], i.e.,
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\left(\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right),$ (25)
which leads to the requirement $\eta=\mathcal{O}(\lambda_{0})$. Instead, we can generalize easily the method used in the $L^{2}$ regression problems to PINNs and obtain the following recursion formula.
Lemma 6.
For all $k\in\mathbb{N}$, we have
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k),$ (26)
where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}}(k),I_{1}^{n_{1}+1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$ and for $p\in[n_{1}]$,
$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle,$ (27)
for $j\in[n_{2}]$,
$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle.$ (28)
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
By using the recursion formula, we arrive at our main result.
Theorem 2.
Under Assumption 3 and Assumption 4, if we set the number of hidden nodes
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
and the learning rate $\eta=\mathcal{O}\left(\frac{1}{\|\bm{H}^{\infty}\|_{2}}\right)$, then with probability at least $1-\delta$ over the random initialization, the gradient descent algorithm satisfies
$\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\left\|\begin{bmatrix}\bm{s}(0)\\ \bm{h}(0)\end{bmatrix}\right\|_{2}^{2}$ (29)
for all $k\in\mathbb{N}$.
Remark 3.
It may be confused that [19] has uesd the same method in [10], but it only requires $\eta=\mathcal{O}(\lambda_{0})$. Actually, it is due to that the loss function of PINN has been normalized. If we let $n_{1}=n_{2}=n$ and $\widetilde{\bm{H}}^{\infty}$ be the Gram matrix induced by unnormalized loss function of PINN, then $\lambda_{min}(\bm{H}^{\infty})=\lambda_{min}(\widetilde{\bm{H}}^{\infty})/n$, leading to the convergence rate similar to that of regression problem. Due to the normalization of loss function, $\|\bm{H}^{\infty}\|_{2}$ is independent of the sample size $n$, but may depends on the dimension $d$.
4. Convergence of Natural Gradient Descent for Two-Layer Physics-Informed Neural Networks
4.1. Problem Setup
Although we have improved the learning rates of gradient descent for $L^{2}$ regression problems and PINNs, one needs to set the learning rates to be small enough due to the agnostic $\|\bm{H}^{\infty}\|_{2}$. For instance, we may need to set it to be $\mathcal{O}(1/n)$ for the $L^{2}$ regression problem. Moreover, the convergence rate $1-\frac{\eta\lambda_{0}}{2}$ also depends on $\lambda_{0}$, which may be slow with small $\lambda_{0}$. [21] and [16] have provided the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer neural networks for $L^{2}$ regression problems. They showed that the maximal learning rate can be $\mathcal{O}(1)$ and the convergence rate is independent of $\lambda_{0}$, which result in a faster convergence rate. In the section, we conduct the convergence analysis of NGD for PINNs and show that NGD indeed leads to an even faster convergence rate.
We consider the same setup as Section 3 and aim to minimize the following empirical loss function via NGD.
$L(\bm{w}):=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right).$ (30)
The NGD gives the following update rule:
$\bm{w}(k+1)=\bm{w}(k)-\eta\bm{J}(k)^{T}\left(\bm{J}(k)\bm{J}(k)^{T}\right)^{-1}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$ (31)
where $\bm{J}(k)=\left(\bm{J}_{1}(k)^{T},\cdots,\bm{J}_{n_{1}+n_{2}}(k)^{T}\right)^{T}\in\mathbb{R}^{(n_{1}+n_{2})\times m(d+2)}$ is the Jacobian matrix for the whole dataset and $\eta>0$ is the learning rate. Specifically, for $p\in[n_{1}]$,
$\bm{J}_{p}(k)=\left[\left(\frac{\partial s_{p}(k)}{\partial\bm{w}_{1}}\right)^{T},\cdots,\left(\frac{\partial s_{p}(k)}{\partial\bm{w}_{m}}\right)^{T}\right]\in\mathbb{R}^{1\times m(d+2)}$ (32)
and for $j\in[n_{2}]$,
$\bm{J}_{n_{1}+j}(k)=\left[\left(\frac{\partial h_{j}(k)}{\partial\bm{w}_{1}}\right)^{T},\cdots,\left(\frac{\partial h_{j}(k)}{\partial\bm{w}_{m}}\right)^{T}\right]\in\mathbb{R}^{1\times m(d+2)}.$ (33)
For the activation function of the two-layer neural network
$\phi(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x}),$ (34)
we consider both settings that $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function or a smooth activation function that satisfies the following assumption. From Lemma 3.3 in [19] and Lemma 2 in [23], we know that $\bm{H}^{\infty}$ is strictly positive definite in both settings and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$.
Assumption 5.
There exists a constant $c>0$ such that $|\sigma(0)|\leq c$ and for any $z,z^{{}^{\prime}}\in\mathbb{R}$,
$|\sigma^{(k)}(z)-\sigma^{(k)}(z^{{}^{\prime}})|\leq c|z-z^{{}^{\prime}}|,$ (35)
where $k\in\{0,1,2,3\}$. Moreover, $\sigma(\cdot)$ is analytic and is not a polynomial function.
Unlike the approach for gradient descent, [21] and [16] focus on the change of the Jacobian matrix for NGD, rather than focusing on the Gram matrix. More precisely, they show that $\bm{J}(\bm{w})$ is stable with respect to $\bm{w}$, where $\bm{J}(\bm{w})$ is the Jacobian matrix with weight vector $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}$. However, this approach does not work for PINNs. Instead, we consider the stability of $\bm{J}(\bm{w})$ with respect to every single weight vector $\bm{w}_{r}$.
Lemma 7 (Stable Jacobian).
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-P(\delta,m,R)$ the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+2}$ that satsify for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
(1) when $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, we have that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}\leq CM\sqrt{R},$ (36)
where $C$ is a universal constant, $M=2(d+2)\log(2m(d+2)/\delta)$ and
$P(\delta,m,R)=1-\delta-n_{1}e^{-mR};$ (37)
(2) when $\sigma(\cdot)$ satisfies Assumption 5, we have that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}\leq CdR$ (38)
for $m\geq\log^{2}(1/\delta)$, where $C$ is a universal constant and $P(\delta,m,R)=\delta$.
With the stability of Jacobian matrix, we can derive the following main result.
Theorem 3 (Convergence of NGD).
Let $L(k)=L(\bm{w}(k))$, then the following conclusions hold.
(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, under Assumption 4, we set
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
and $\eta\in(0,1)$, then with probability at least $1-\delta$ over the random initialization for all $k\in\mathbb{N}$
$L(k)\leq(1-\eta)^{k}L(0).$ (39)
(2) When $\sigma(\cdot)$ satisfies Assumption 5, under Assumption 4, we set
$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
and $\eta\in(0,1)$, then with probability at least $1-\delta$ over the random initialization for all $k\in\mathbb{N}$
$L(k)\leq(1-\eta)^{k}L(0).$ (40)
In fact, NGD can enjoy a second-order convergence rate when $\sigma(\cdot)$ satisfies Assumption 5.
Corollary 1.
Under the setting of Theorem 3(2), in addition to (40), we also have that
$\left\|\begin{pmatrix}\bm{s}(t+1)\\ \bm{h}(t+1)\end{pmatrix}\right\|_{2}\leq\frac{CB^{4}}{\sqrt{m\lambda_{0}^{3}}}\left\|\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}\right\|_{2}^{2},$
where $C$ is a universal constant and $B=\sqrt{2(d+2)\log(2m(d+2)/\delta)}+1$.
5. Conclusion and Discussion
In this paper, we have improved the learning rate of gradient descent for both $L^{2}$ regression problems and PINNs, indicating that gradient descent enjoys better convergence. Furthermore, we demonstrate that natural gradient descent can find the global optima of two-layer PINNs with $\text{ReLU}^{3}$ or smooth activation functions for a class of second-order linear PDEs. Compared to gradient descent, natural gradient descent possesses faster convergence rate and the maximal learning rate is $\mathcal{O}(1)$. Despite this, natural gradient descent is quite expensive in terms of cumputation and memory in training neural networks. Therefore, several cost-effective variants have been proposed, such as K-FAC [9] and mini-batch natural gradient descent [16]. It would be interesting to investigate the convergence of these methods for PINNs. Additionally, generalizing the convergence analysis to deep neural networks is an important direction for future work.
Appendix
6. Proof of Section 2
Before the proofs, we first define the event
$A_{ir}:=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{i}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\}\}$ (41)
for all $i\in[n]$.
Note that the event happens if and only if $|\bm{w}_{r}(0)^{T}\bm{x}_{i}|<R$, thus by the anti-concentration inequality of Gaussian distribution, we have
$P(A_{ir})=P_{z\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2}^{2})}(|z|<R)=P_{z\sim\mathcal{N}(0,1)}\left(|z|<\frac{R}{\|\bm{x}_{i}\|_{2}}\right)\leq\frac{2R}{\sqrt{2\pi}\|\bm{x}_{i}\|_{2}}\leq\frac{2R}{\sqrt{2\pi}},$ (42)
as $\|\bm{x}_{i}\|_{2}\geq 1$ for all $i\in[n]$.
6.1. Proof of Lemma 2
Proof.
Recall that
$\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2}$ (43)
$=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}(\bm{x}_{i}^{T}\bm{x}_{j})^{2}\left(\frac{1}{m}\sum\limits_{r=1}^{m}(I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\})\right)^{2}.$
To prove the measurability, we will show that $I\left\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\right\}$ can be approximated by the sequence $\{I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0,\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\}_{k\in\mathbb{N}}$, where $\tilde{\bm{w}}_{k}$ is a rational vector in $\mathbb{Q}^{d+1}$ for each $k\in\mathbb{N}$, with $\mathbb{Q}$ denoting the set of all rational numbers. Therefore, taking the supremum of $\bm{w}$ within the ball is equivalent to taking the supremum over all rational vectors within this ball, which implies that the random variable is measurable.
We first focus on $I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}$ and let $\bm{w}_{r}=(\bm{w}_{r1}^{T},w_{r0})^{T},\bm{x}_{i}=(\bm{x}_{i1}^{T},1)^{T}$ with $w_{r0}\in\mathbb{R}$ and $\bm{w}_{r1},\bm{x}_{i1}\in\mathbb{R}^{d}$. For each $k\in\mathbb{N}$, we can take $\bm{\beta}_{k}\in\mathbb{Q}^{d}$ such that $\|\bm{\beta}_{k}-\bm{w}_{r1}\|_{2}\leq\frac{1}{2k}$. Then choose $t_{k}\in\mathbb{Q}$ such that $t\in(w_{r0}+\frac{1}{2k},w_{r0}+\frac{1}{k}]$ and let $\tilde{\bm{w}}_{k}=(\bm{\beta}_{k}^{T},t_{k})^{T}$. From the construction, we have
$I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0\}=I\{\bm{\beta}_{k}^{T}\bm{x}_{i1}+t_{k}\geq 0\}=I\{\bm{w}_{r}^{T}\bm{x}_{i}+t_{k}-w_{r0}+(\bm{\beta}_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}\geq 0\},$
as $\bm{w}_{r}^{T}\bm{x}_{i}=\bm{w}_{r1}^{T}\bm{x}_{i1}+w_{r0}$. Define $r_{k}=t_{k}-w_{r0}+(\beta_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}$, then $r_{k}>0$ and $r_{k}\to 0$, as $|(\beta_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}|\leq\|\beta_{k}-\bm{w}_{r1}\|_{2}\leq\frac{1}{2k}$ and $t-w_{r0}\in(\frac{1}{2k},\frac{1}{k}]$. Since the function $u\rightarrow I\{u\geq x\}$ is right-continuous for any $x\in\mathbb{R}$. Thus $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0\}=I\{\bm{w}_{k}^{T}\bm{x}_{i}+r_{k}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{i}\geq 0\}$ as $k\to\infty$. Meanwhile, when $k$ is large enough, $\tilde{\bm{w}}_{k}$ is in the small ball centered at $\bm{w}_{r}(0)$, i.e., $\|\tilde{\bm{w}}_{k}-\bm{w}_{r}(0)\|_{2}<R$. Similarly, we have that $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{j}\geq 0\}$ as $k\to\infty$. Thus, we can deduce that $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0,\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{i}\geq 0,\bm{w}_{k}^{T}\bm{x}_{j}\geq 0\}$ as $k\to\infty$.
Therefore,
$\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2}=\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\bm{w}_{1}\in\mathbb{Q}^{d+1}\\ \cdots\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R,\bm{w}_{m}\in\mathbb{Q}^{d+1}\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2},$
which leads to the measurability.
From the definition of $A_{ir}$, i.e., (41), and the fact $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, we can deduce that
$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\}|\leq I\{A_{ir}\}.$
Thus, when both $A_{ir}$ and $A_{jr}$ do not happen, we have
$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|$
$\leq|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\geq 0\}|+|I\{\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\geq 0\}|$
$\leq I\{A_{ir}\}+I\{A_{jr}\}=0,$
which imples
$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|\leq I\{A_{ir}\lor A_{jr}\}.$ (44)
Therefore, combining (43) and (44) yields that
$\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots,\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(w)_{ij}-H(0)_{ij}|^{2}$ (45)
$\leq 4\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots,\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\left(\frac{1}{m}\sum\limits_{r=1}^{m}(I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\})\right)^{2}$
$\leq 4\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\left(\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\right)^{2}.$
It remains only to bound the term $\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}$. Note that
$\mathbb{E}[I\{A_{ir}\lor A_{jr}\}]\leq P(A_{ir})+P(A_{jr})\leq\frac{4R}{\sqrt{2\pi}}$
and then
$Var(I\{A_{ir}\lor A_{jr}\})\leq\mathbb{E}[I\{A_{ir}\lor A_{jr}\}\leq\frac{4R}{\sqrt{2\pi}}.$
Thus, applying Bernstein’s inequality for the random variable $I\{A_{ir}\lor A_{jr}\}-\mathbb{E}[I\{A_{ir}\lor A_{jr}\}]$ yields that with probability at least $1-e^{-t}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\leq\frac{4R}{\sqrt{2\pi}}+\sqrt{\frac{2t}{m}\frac{4R}{\sqrt{2\pi}}}+\frac{t}{3m}.$
Choosing $t=mR$ and taking a union bound for $i,j\in[n]$ yields that with probability at least $1-n^{2}e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\leq 4R$ (46)
holds for all $i,j\in[n]$.
Plugging (46) into (45) leads to the conclusion. ∎
6.2. Proof of Lemma 3
Proof.
First, we can decompose $u_{i}(k+1)-u_{i}(k)$ as follows.
$u_{i}(k+1)-u_{i}(k)$ $=u_{i}(k+1)-u_{i}(k)-\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle+\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$
$:=I_{1}^{i}(k)+I_{2}^{i}(k).$
For $I_{2}^{i}(k)$, we have
$I_{2}^{i}(k)$ $=\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$
$=\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle$
$=\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\rangle$
$=\sum\limits_{r=1}^{m}(-\eta)\sum\limits_{j=1}^{n}(u_{j}-y_{j})\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\frac{\partial u_{j}(k)}{\partial\bm{w}_{r}}\rangle$
$=\eta\sum\limits_{j=1}^{n}(\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\frac{\partial u_{j}(k)}{\partial\bm{w}_{r}}\rangle)(y_{j}-u_{j})$
$=\eta[\bm{H}(k)]_{i}(y-u(k)),$
where $[\bm{H}(k)]_{i}$ denotes the $i$-row of the matrix $\bm{H}(k)$.
Thus,
$u_{i}(k+1)-u_{i}(k)=\eta[\bm{H}(k)]_{i}(y-u(k))+I_{1}^{i}(k)$
and then
$(y_{i}-u_{i}(k))-(y_{i}-u_{i}(k+1))=\eta[\bm{H}(k)]_{i}(y-u(k))+I_{1}^{i}(k),$
which yields
$(\bm{y}-\bm{u}(k))-(\bm{y}-\bm{u}(k+1))=\eta\bm{H}(k)(\bm{y}-\bm{u}(k))+I_{1}(k).$
A simple transformation leads to the conclusion
$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k).$
6.3. Proof of Theorem 1
Proof.
The proof follows a similar induction to that in [10]. Our induction hypothesis is the following condition.
Condition 1.
At the $t$-th iteration, we have
$\|\bm{y}-\bm{u}(t)\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{t}\|\bm{y}-\bm{u}(0)\|_{2}^{2}.$ (47)
Suppose that Condition 1 holds for $t=0,\cdots,k$, then Corollary 4.1 in [10] implies that for every $r\in[m]$ and $t=0,\cdots,k$
$\|\bm{w}_{r}(t+1)-\bm{w}_{r}(0)\|_{2}\leq\frac{4\sqrt{n}\|\bm{y}-\bm{u}(0)\|_{2}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}}.$ (48)
In the following, we aim to show that Condition 1 also holds for $t=k+1$, thus the conclusion of Theorem 1 holds directly.
Let $S_{i}=\{r\in[m]:I\{A_{ir}\}=0\}$ and $S_{i}^{\perp}=[m]\textbackslash S_{i}$.
Recall that
$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k).$
As Condition 1 holds for $t=0,\cdots,k$, from (48), we have that for any $r\in[m]$,
$\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}.$
By taking $R=\frac{\lambda_{0}}{32n}$ in Lemma 2 and $R^{{}^{\prime}}\leq R$, we have that
$\|\bm{H}(k)-\bm{H}(0)\|_{2}\leq\frac{\lambda_{0}}{4},$
which implies that $\lambda_{min}(\bm{H}(k))\geq\lambda_{min}(\bm{H}(0))-\frac{\lambda_{0}}{4}\geq\frac{\lambda_{0}}{2}$ and
$\|\bm{H}(k)\|_{2}\leq\|\bm{H}(0)\|_{2}+\frac{\lambda_{0}}{4}\leq\|\bm{H}^{\infty}\|_{2}+\frac{\lambda_{0}}{2}\leq\frac{3}{2}\|\bm{H}^{\infty}\|_{2}.$
Thus, when $\eta\leq\frac{2}{3\|H^{\infty}\|_{2}}$, $\bm{I}-\eta\bm{H}(k)$ is positive definite and meanwhile $\|\bm{I}-\eta\bm{H}(k)\|_{2}\leq 1-\frac{\eta\lambda_{0}}{2}$. Combining with the recursion formula (10) yields that
$\|\bm{y}-\bm{u}(k+1)\|_{2}^{2}$ $=\|(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k)\|_{2}^{2}$ (49)
$=\|(I-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}-2\langle(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k)),\bm{I}_{1}(k)\rangle$
$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\|\bm{I}_{1}(k)\|_{2}\|(I-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))\|_{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\|\bm{I}_{1}(k)\|_{2}\|\bm{y}-\bm{u}(k)\|_{2}.$
Then, it remains only to bound the term $\|\bm{I}_{1}(k)\|_{2}$.
Recall that for $i\in[n]$,
$I_{1}^{i}(k)=\sum\limits_{r=1}^{m}\frac{a_{r}}{\sqrt{m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\right].$
Note that for $r\in S_{i}$, we can deduce that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{i}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$, as $\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$ and $\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$. Thus, for $r\in S_{i}$, we have
$\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}i)I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{i}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}]$
$\quad-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}]$
$\quad-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
$=0.$
It implies that
$I_{1}^{i}(k)=\sum\limits_{r\in S_{i}^{\perp}}\frac{a_{r}}{\sqrt{m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\right].$
Therefore,
$|I_{1}^{i}(k)|$ $\leq\frac{1}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}2\sqrt{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}$ (50)
$=\frac{2\sqrt{2}}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}\|-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\|_{2}$
$\leq\frac{4}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}\frac{\eta\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{\sqrt{m}}$
$=\frac{4\eta\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\},$
where the second inequality follows from that
$\frac{\partial L(k)}{\partial\bm{w}_{r}}$ $=\sum\limits_{i=1}^{n}(u_{i}(k)-y_{i})\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}}$
$=\sum\limits_{i=1}^{n}(u_{i}(k)-y_{i})\frac{a_{r}}{\sqrt{m}}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\bm{x}_{i}$
and then
$\left\|\frac{\partial L(k)}{\partial\bm{w}_{r}}\right\|_{2}\leq\frac{\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{\sqrt{m}},$
as $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$.
Since $S_{i}=\{r\in[m]:I\{A_{ir}\}=0\}$, it follows that $I\{r\in S_{i}^{\perp}\}=I\{A_{ir}\}$, then applying Berstein’s inequality yields that with probability at least $1-ne^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\}\leq 4R,$ (51)
holds for all $i\in[n]$.
Thus from (50) and (51), we have
$\|I_{1}(k)\|_{2}$ $=\sqrt{\sum\limits_{i=1}^{n}|I_{1}^{i}(k)|^{2}}$ (52)
$\leq 4\eta\sqrt{n}\|y-u(k)\|_{2}\sqrt{\sum\limits_{i=1}^{n}(\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\})^{2}}$
$\leq 8\eta nR\|y-u(k)\|_{2}$
$=\frac{\eta\lambda_{0}}{8}\|y-u(k)\|_{2},$
where the last inequality is due to $R=\frac{\lambda_{0}}{64n}$.
Plugging (52) into (49) yields that
$\|y-u(k+1)\|_{2}^{2}$ $\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\|\bm{I}_{1}(k)\|_{2}\|\bm{y}-\bm{u}(k)\|_{2}$
$\leq\left[(1-\frac{\eta\lambda_{0}}{2})^{2}+\frac{\eta^{2}\lambda_{0}^{2}}{64}+2(1-\frac{\eta\lambda_{0}}{2})\frac{\eta\lambda_{0}}{8}\right]\|\bm{y}-\bm{u}(k)\|_{2}^{2}$
$\leq\left[1-\eta\lambda_{0}+\frac{\eta^{2}\lambda_{0}^{2}}{4}+\frac{\eta^{2}\lambda_{0}^{2}}{64}+\frac{\eta\lambda_{0}}{4}\right]\|\bm{y}-\bm{u}(k)\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})\|\bm{y}-\bm{u}(k)\|_{2}^{2},$
where the last inequality follows from the fact that
$\eta\leq\frac{2}{3\|\bm{H}^{\infty}\|_{2}}\leq\frac{2}{3\lambda_{0}},$
which imples that
$\frac{\eta^{2}\lambda_{0}^{2}}{4}+\frac{\eta^{2}\lambda_{0}^{2}}{64}\leq\frac{\eta\lambda_{0}}{4}.$
Finally, we need consider the requirement for $m$. First, Lemma 1 shows that $m=\Omega\left(\frac{n^{2}}{\lambda_{0}^{2}}\log\left(\frac{n}{\delta}\right)\right)$. Second, from $R^{{}^{\prime}}\leq R=\frac{\lambda_{0}}{32n}$ and $R^{{}^{\prime}}=\frac{4\sqrt{n}\|\bm{y}-\bm{u}(0)\|_{2}}{\sqrt{m}\lambda_{0}}$, we know
$m=\Omega\left(\frac{n^{3}\|\bm{y}-\bm{u}(0)\|_{2}^{2}}{\lambda_{0}^{4}}\right).$
Then from the estimation in Lemma 10 for the initial prediction $\|\bm{y}-\bm{u}(0)\|_{2}$, we obtain that
$m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}}\log\left(\frac{n}{\delta}\right)\right).$
7. Proof of Section 3
Before the proofs, we first recall that
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{1}}}\left(\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}\right)$ (53)
and
$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{2}}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j}.$ (54)
7.1. Proof of Lemma 4
Proof.
In the following, we aim to bound $\|\bm{H}(0)-\bm{H}^{\infty}\|_{F}$, as $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\|\bm{H}(0)-\bm{H}^{\infty}\|_{F}$. Note that the entries of $\bm{H}(0)-\bm{H}^{\infty}$ have three forms as follows.
$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle,$ (55)
$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle$ (56)
and
$\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle.$ (57)
For the first form (55), to simplify the analysis, we let
$\bm{Z}_{r}(i)$ $=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})w_{r0}(0)\bm{x}_{i}+\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}$
$\quad-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\bm{x}_{p}-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})\begin{pmatrix}0\\ \bm{w}_{r1}(0)\end{pmatrix}$
and
$X_{r}(ij)=\langle\bm{Z}_{r}(i),\bm{Z}_{r}(j)\rangle,$
then
$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle=\frac{1}{n_{1}m}\sum\limits_{r=1}^{m}(X_{r}(ij)-\mathbb{E}X_{r}(ij)).$
Note that $|X_{r}(ij)|\lesssim 1+\|\bm{w}_{r}(0)\|_{2}^{4}$, thus
$\|X_{r}(ij)\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{4}\right\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{2}\right\|_{\psi_{1}}^{2}\lesssim d^{2}.$
For the centered random variable, the property of $\psi_{\frac{1}{2}}$ quasi-norm implies that
$\|X_{r}(ij)-\mathbb{E}[X_{r}(ij)]\|_{\psi_{\frac{1}{2}}}\lesssim\|X_{r}(ij)\|_{\psi_{\frac{1}{2}}}+\|\mathbb{E}[X_{r}(ij)]\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}.$
Therefore, applying Lemma 8 yields that with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\frac{1}{m}(X_{r}(ij)-\mathbb{E}[X_{r}(ij)])|\lesssim\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2},$
which means
$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}w_{r}}\rangle|\lesssim\frac{d^{2}}{n_{1}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{1}m}(\log(\frac{1}{\delta}))^{2}.$ (58)
Similarly, for the second form (56) and third form (57), we can deduce that
$\left\|\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}$
and
$\left\|\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{n_{2}m}.$
Thus applying Lemma 8 yields that with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}\log(\frac{1}{\delta})$ (59)
and with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{n_{2}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{2}m}\log(\frac{1}{\delta}).$ (60)
Combining (58), (59) and (60), we can deduce that with probability at least $1-\delta$,
$\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}^{2}$
$\leq\|\bm{H}(0)-\bm{H}^{\infty}\|_{F}^{2}$
$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})+\frac{d^{4}}{m^{2}}(\log(\frac{n_{1}+n_{2}}{\delta}))^{4}$
$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta}).$
Thus when $\sqrt{\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})}\lesssim\frac{\lambda_{0}}{4}$, i.e.,
$m=\Omega(\frac{d^{4}}{\lambda_{0}^{2}}\log(\frac{n_{1}+n_{2}}{\delta})),$
we have $\lambda_{min}(\bm{H}(0))\geq\frac{3}{4}\lambda_{0}$.
7.2. Proof of Lemma 5
Proof.
We first reformulate the term $\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ in (53) as follows.
$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}w_{r0}(k)x_{p0}\\ w_{r0}(k)\bm{x}_{p1}-2\bm{w}_{r1}(k)\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}.$
Similar to Lemma 4, it suffices to bound $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}$, which can in turn allows us to bound each entry of $\bm{H}(\bm{w})-\bm{H}(0)$.
For $i\in[n_{1}]$ and $j\in[n_{1}]$, we have
$H_{ij}(\bm{w})$ $=\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}}{\partial\bm{w}_{r}},\frac{\partial s_{j}}{\partial\bm{w}_{r}}\rangle$
$=\frac{1}{n_{1}m}\sum\limits_{r=1}^{m}\left\langle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}w_{r0}x_{i0}\\ w_{r0}\bm{x}_{i1}-2\bm{w}_{r1}\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{i},\right.$
$\qquad\left.\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}w_{r0}x_{j0}\\ w_{r0}\bm{x}_{j1}-2\bm{w}_{r1}\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{j}\right\rangle$
After expanding the inner product term, we can find that although it has nine terms, it only consists of six classes. For simplicity, we use the following six symbols to represent the corresponding classes.
$\sigma^{{}^{\prime\prime}}\sigma^{{}^{\prime\prime}},\sigma^{{}^{\prime\prime}}\sigma^{{}^{\prime}},\sigma^{{}^{\prime}}\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}},\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime\prime}}.$
For instance, $\sigma^{{}^{\prime\prime}}\sigma^{{}^{\prime}}$ represents
$\left\langle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}w_{r0}x_{i0}\\ w_{r0}\bm{x}_{i1}-2\bm{w}_{r1}\end{pmatrix},\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}\right\rangle,\left\langle\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix},\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}w_{r0}x_{j0}\\ w_{r0}\bm{x}_{j1}-2\bm{w}_{r1}\end{pmatrix}\right\rangle.$
In fact, when bounding the corresponding terms for $H_{ij}(\bm{w})-H_{ij}(0)$, the first four classes can be grouped into one category. They are of the form $f_{1}(\bm{w})f_{2}(\bm{w})f_{3}(\bm{w})f_{4}(\bm{w})$, where for each $1\leq i\leq 4$, $f_{i}(\bm{w})$ is Lipschitz continuous with respect to $\|\cdot\|_{2}$ and $|f_{i}(\bm{w})|\lesssim\|\bm{w}\|_{2}$ (Note that $\sigma^{{}^{\prime}}(\cdot)=(\sigma^{{}^{\prime\prime}}(\cdot))^{2}$). On the other hand, when $\|\bm{w}_{1}-\bm{w}_{2}\|_{2}\leq R\leq 1$, we can deduce that
$|f_{1}(\bm{w}_{1})f_{2}(\bm{w}_{1})f_{3}(\bm{w}_{1})f_{4}(\bm{w}_{1})-f_{1}(\bm{w}_{2})f_{2}(\bm{w}_{2})f_{3}(\bm{w}_{2})f_{4}(\bm{w}_{2})|\lesssim R(\|\bm{w}_{1}\|_{2}^{3}+1).$
Thus, for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the first four classes, we can deduce that they are less than $CR(\|\bm{w}_{r}(0)\|_{2}^{3}+1)$, where $C$ is a universal constant.
For the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are both involving $\sigma^{{}^{\prime\prime\prime}}$ that is not Lipschitz continuous. To make it precise, we write the class $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ explicitly as follows.
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{2}\begin{pmatrix}w_{r0}x_{i0}\\ w_{r0}\bm{x}_{i1}-2\bm{w}_{r1}\end{pmatrix}^{T}\bm{x}_{j}.$
Note that when $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, we have that
$|\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{j})|=|I\{\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|\leq I\{A_{jr}\},$
where the definition of $A_{jr}$ is same as (44) for the regression problems.
Thus, we can deduce that for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are less than
$C\left[(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{3}+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1)\right],$
where $C$ is a universal constant.
Similarly, for the last class $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime\prime}}$ that are of the form
$\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j},$
we can deduce that
$|\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{j})\|\bm{w}_{r1}(0)\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j}|$
$\lesssim I\{A_{ir}\lor A_{jr}\}\|\bm{w}_{r}(0)\|_{2}^{4}+R\|\bm{w}_{r}(0)\|_{2}^{3}+R.$
Combining the upper bounds for the six classes, we have that
$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{3})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right]$ (61)
$\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right],$
where the last inequality follows from that $\|\bm{w}_{r}(0)\|_{2}^{3}\lesssim\|\bm{w}_{r}(0)\|_{2}^{4}+1$ due to Young’s inequality for products.
Now, we focus on the term $\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}$.
Since
$P(|w_{ri}(0)|^{2}\geq 2\log(\frac{2}{\delta}))\leq\delta$
and then
$P(\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log(\frac{2(d+2)}{\delta}))\leq\delta,$
we obtain that
$P(\exists r\in[m],\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\frac{2m(d+2)}{\delta})\leq\delta.$ (62)
Let $M=2(d+2)\log(\frac{2m(d+2)}{\delta})$, then
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}$
$=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}\leq M\}+\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}$
$\leq\frac{M^{2}}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}.$
Applying Bernstein’s inequality for the first term yields that with probability at least $1-e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\leq 4R.$
Moreover, with probability at least $1-\delta$, $I\{\|w_{r}(0)\|_{2}^{2}>M\}=0$ holds for all $r\in[m]$.
Thus from (61), with probability at least $1-\delta-n_{1}e^{-mR}$, we have that for any $i\in[n_{1}]$ and $j\in[n_{1}]$,
$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[RM^{2}+RM^{2}+R\right]$
$\lesssim\frac{1}{n_{1}}M^{2}R.$
For $i\in[n_{1}],j\in[n_{1}+2,n_{2}]$ and $i\in[n_{1}+1,n_{2}],j\in[n_{2}]$, from the form of $\frac{\partial h_{j}}{\partial\bm{w}_{r}}$, i.e.,
$\frac{\partial h_{j}}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j},$
we can obtain similar results for the terms $\langle\frac{\partial s_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$ and $\langle\frac{\partial h_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$.
With all results above, we have that with probability at least $1-\delta-(n_{1}+n_{2})e^{-mR}$,
$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}\lesssim M^{2}R.$
7.3. Proof of Lemma 6
Proof.
Similar to the proof of Lemma 3, we have
$s_{p}(k+1)-s_{p}(k)$ $=\left[s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (63)
$:=I_{1}^{p}(k)+I_{2}^{p}(k).$
For the second term $I_{2}^{p}(k)$, we have that
$I_{2}^{p}(k)$ $=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (64)
$=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},-\eta\frac{\partial L(k)}{\partial\bm{w}}\rangle$
$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial L(k)}{\partial\bm{w}_{r}}\rangle$
$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\sum\limits_{t=1}^{n_{1}}s_{t}(k)\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}+\lambda\sum\limits_{j=1}^{n_{2}}h_{j}(k)\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\rangle$
$=-\eta[\bm{H}(k)]_{p}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$
where $[\bm{H}(k)]_{p}$ denotes the $p$-row of $\bm{H}(k)$.
Similarly, for $h(k)$, we have
$h_{j}(k+1)-h_{j}(k)$ $=\left[h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (65)
$:=I_{1}^{n_{1}+j}(k)+I_{2}^{n_{1}+j}(k)$
and
$I_{2}^{n_{1}+j}(k)=-\eta[\bm{H}(k)]_{n_{1}+j}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$ (66)
Combining (63), (64), (65) and (66) yields that
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}$ $=\bm{I}_{1}(k)+\bm{I}_{2}(k)$
$=\bm{I}_{1}(k)-\eta\bm{H}(k)\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$
A simple transformation directly leads to
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
7.4. Proof of Theorem 2
Proof.
The proof strategy is similar to that for Theorem 1. Our induction hypothesis is the following boundedness of the hidden weights and convergence rate of the empirical loss.
Condition 2.
At the $t$-th iteration, we have that for each $r\in[m]$, $\|\bm{w}_{r}(t)\|_{2}\leq B$ and
$L(t)\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{t}L(0),$ (67)
where $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1$.
From (62), we know that with probability at least $1-\delta$, $\|\bm{w}_{r}(0)\|_{2}\leq\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}$ holds for any $r\in[m]$. Thus, if we can prove that $\bm{w}_{r}(t)$ is closed enough to $\bm{w}_{r}(0)$, then $\|\bm{w}_{r}(t)\|_{2}\leq B$.
Corollary 2 (Lemma 4.1 in [19]).
If Condition 2 holds for $t=0,\cdots,k$, then we have for every $r\in[m]$,
$\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}},$ (68)
where $C$ is a universal constant.
Corollary 2 implies that when $m$ is large enough, we have $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq 1$ and then $\|\bm{w}_{r}(k+1)\|_{2}\leq B$. Thus, in induction, we only need to prove that (67) also holds for $t=k+1$, which relies on the recursion formula (26).
Recall that
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
From Corollary 2 and Lemma 5, taking $CM^{2}R<\frac{\lambda_{0}}{4}$ in (24) and $R^{{}^{\prime}}\leq R$ in (68) yields that $\lambda_{min}(\bm{H}(k))\geq\frac{\lambda_{0}}{2}$ and
$\|\bm{H}(k)\|_{2}\leq\|\bm{H}(0)\|_{2}+\frac{\lambda_{0}}{4}\leq\|\bm{H}^{\infty}\|_{2}+\frac{\lambda_{0}}{2}\leq\frac{3}{2}\|\bm{H}^{\infty}\|_{2}.$
Therefore, if we take $\eta\leq\frac{2}{3}\frac{1}{\|\bm{H}^{\infty}\|_{2}}$, then $\bm{I}-\eta\bm{H}(k)$ is positive definite and $\|I-\eta\bm{H}(k)\|_{2}\leq 1-\frac{\eta\lambda_{0}}{2}$.
Combining these facts with the recursion formula, we have that
$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$ (69)
$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left\langle(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},\bm{I}_{1}(k)\right\rangle$
$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}.$
Thus, it remains to bound $\|\bm{I}_{1}(k)\|_{2}$.
For $\bm{I}_{1}(k)$, recall that $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}}(k),I_{1}^{n_{1}+1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$ and for $p\in[n_{1}]$,
$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle,$
for $j\in[n_{2}]$,
$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$
Recall that
$s_{p}(k)=\frac{1}{\sqrt{n_{1}}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}x_{p})w_{r0}(k)-a_{r}\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}x_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}-f(x_{p})\right)$
and
$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+2}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}$
$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}(k)\end{pmatrix}].$
Define $\chi_{pr}^{1}(k):=\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$ and $\chi_{pr}^{2}(k):=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$, i.e., $\chi_{pr}^{1}(k)$ and $\chi_{pr}^{2}(k)$ are related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ respectively.
Then define
$\hat{\chi}_{pr}^{1}(k)=\chi_{pr}^{1}(k+1)-\chi_{pr}^{1}(k)-\langle\frac{\partial\chi_{pr}^{1}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle$
and
$\hat{\chi}_{pr}^{2}(k)=\chi_{pr}^{2}(k+1)-\chi_{pr}^{2}(k)-\langle\frac{\partial\chi_{pr}^{2}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle.$
At this time, we have
$I_{1}^{p}(k)=\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}a_{r}\left[\hat{\chi}_{pr}^{1}(k)-\hat{\chi}_{pr}^{2}(k)\right].$
The purpose of defining them in this way is to enable us to handle the terms related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ separately.
Similar to that for regression problems with ReLU activation function, we first define some events. For $p\in[n_{1}]$, define
$A_{p,r}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}\}$
and let $S_{p}^{\perp}=[n_{1}]\textbackslash S_{p}$.
In the following, we are going to show that $|\hat{\chi}_{pr}^{1}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for every $r\in[m]$ and $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for $r\in S_{p}$, $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2})$ for $r\in S_{p}^{\perp}$. Thus, we can prove that $\|\bm{I}_{1}\|_{2}=\mathcal{O}\left(\frac{\sqrt{L(k)}}{\sqrt{m}}\right)$. Then combining with (69) leads to a milder requirement for $\eta$.
For $\hat{\chi}_{pr}^{1}(k)$, we have that
$\hat{\chi}_{pr}^{1}(k)$ $=\sigma^{{}^{\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})w_{r0}(k+1)-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)-(w_{r0}(k+1)-w_{r0}(k))\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})$
$=(\sigma^{{}^{\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p}))w_{r0}(k+1)-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k).$
From the mean value theorem, we can deduce that there exists $\zeta(k)\in\mathbb{R}$ such that
$\sigma^{{}^{\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})=\sigma^{{}^{\prime\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle$
and
$|\sigma^{{}^{\prime\prime}}(\zeta(k))-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})|$ $\leq|\zeta(k)-\bm{w}_{r}(k)^{T}\bm{x}_{p}|$
$\leq\sqrt{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
Then, for $\hat{\chi}_{pr}^{1}(k)$, we can rewrite it as follows.
$\hat{\chi}_{pr}^{1}(k)$ $=\sigma^{{}^{\prime\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle w_{r0}(k+1)-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$
$=\left[\left(\sigma^{{}^{\prime\prime}}(\zeta(k))-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\right)\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle w_{r0}(k+1)\right]$
$\quad+\left[\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})(w_{r0}(k+1)-w_{r0}(k))\right].$
It implies that
$|\hat{\chi}_{pr}^{1}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}.$
For $\hat{\chi}_{pr}^{2}(k)$, we write it as follows explicitly.
$\hat{\chi}_{pr}^{2}(k)$ $=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k+1)\|_{2}^{2}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$ (70)
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$
$\quad-2\langle\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k),\bm{w}_{r1}(k)\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p}).$
Note that for the term $\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{w}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$, we can rewrite it as follows.
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$ (71)
$=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)-\bm{w}_{r1}(k+1)+\bm{w}_{r1}(k+1)\|_{2}^{2}$
$=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})[\|\bm{w}_{r1}(k)-\bm{w}_{r1}(k+1)\|_{2}^{2}+\|\bm{w}_{r1}(k+1)\|_{2}^{2}-2\langle\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k),\bm{w}_{r1}(k+1)\rangle],$
where the first term $\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)-\bm{w}_{r1}(k+1)\|_{2}^{2}=\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$.
plugging (71) into (70) yields that
$\hat{\chi}_{pr}^{2}(k)$ $=[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})]\|\bm{w}_{r1}(k+1)\|_{2}^{2}$ (72)
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$
$\quad+2\langle\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k),\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k)\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})+\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$
$=[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})]\|\bm{w}_{r1}(k+1)\|_{2}^{2}$
$\quad+\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})(\|\bm{w}_{r1}(k+1)\|_{2}^{2}-\|\bm{w}_{r1}(k)\|_{2}^{2})$
$\quad+\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$
$=\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\right]\|\bm{w}_{r1}(k+1)\|_{2}^{2}$
$\quad+\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}).$
Thus, we only need to consider the term
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p}).$
For $r\in S_{p}$, as $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq R,\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R$, we have that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$, which yields that
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})$ (73)
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}]$
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}]$
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$
$=0.$
For $r\in S_{p}^{\perp}$, the Lipschitz continuity of $\sigma^{{}^{\prime\prime}}$ implies that
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}).$ (74)
Combining (72), (73) and (74), we can deduce that for $r\in S_{p}$,
$|\hat{\chi}_{pr}^{2}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2},$
for $r\in S_{p}^{\perp}$,
$|\hat{\chi}_{pr}^{2}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}+B^{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
With the estimations for $\hat{\chi}_{pr}^{1}(k)$ and $\hat{\chi}_{pr}^{2}(k)$, we have
$|I_{1}^{p}(k)|$ $\leq\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}(|\hat{\chi}_{pr}^{1}(k)|+|\hat{\chi}_{pr}^{2}(k)|)$ (75)
$\lesssim\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}+\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r\in S_{p}^{\perp}}B^{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
For $j\in[n_{2}]$, we consider $I_{1}^{n_{1}+j}(k)$, which can be written as follows.
$I_{1}^{n_{1}+j}(k)$ $=h_{j}(k+1)-h_{j}(k)-\langle\bm{w}(k+1)-\bm{w}(k),\frac{\partial h_{j}(k)}{\partial\bm{w}}\rangle$
$=\sum\limits_{r=1}^{m}\frac{a_{r}}{\sqrt{n_{2}m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{y}_{j})-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})\right].$
From the mean value theorem, we have that there exists $\zeta(k)\in\mathbb{R}$ such that
$\sigma(\bm{w}_{r}(k+1)^{T}\bm{y}_{j})-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})=\sigma^{{}^{\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle$
and
$|\sigma^{{}^{\prime}}(\zeta(k))-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})|$ $\leq 2B|\zeta(k)-\bm{w}_{r}(k)^{T}\bm{y}_{j}|$
$\leq 2\sqrt{2}B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
Thus,
$|\sigma(\bm{w}_{r}(k+1)^{T}\bm{y}_{j})-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})|$
$=|\sigma^{{}^{\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})|$
$=|(\sigma^{{}^{\prime}}(\zeta(k))-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j}))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle|$
$\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
Therefore, for $j\in[n_{2}]$,
$|I_{1}^{n_{1}+j}(k)|\lesssim\frac{B}{\sqrt{n_{2}m}}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}.$ (76)
From the updating rule of gradient descent, we can deduce that for every $r\in[m]$,
$\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}=\left\|-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{\eta B^{2}}{\sqrt{m}}\sqrt{L(k)}.$ (77)
Plugging (77) into (75) and (76), we can deduce that
$|I_{1}^{p}(k)|$ $\lesssim\frac{B}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}+\frac{B^{2}}{\sqrt{n_{1}m}}\sum\limits_{r\in S_{p}^{\perp}}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}$ (78)
$\lesssim\frac{B}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}\frac{\eta^{2}B^{4}}{m}L(k)+\frac{B^{2}}{\sqrt{n_{1}m}}\sum\limits_{r\in S_{p}^{\perp}}\frac{\eta B^{2}}{\sqrt{m}}\sqrt{L(k)}$
$=\frac{\eta^{2}B^{5}L(k)}{\sqrt{n_{1}m}}+\frac{\eta B^{4}\sqrt{L(k)}}{\sqrt{n_{1}}}\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{p}^{\perp}\}$
$\leq\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{1}m}}+\frac{\eta B^{4}\sqrt{L(k)}}{\sqrt{n_{1}}}\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{p}^{\perp}\}$
and
$|I_{1}^{n_{1}+j}(k)|$ $\lesssim\frac{B}{\sqrt{n_{2}m}}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}$ (79)
$\lesssim\frac{B}{\sqrt{n_{2}m}}\sum\limits_{r=1}^{m}\frac{\eta^{2}B^{4}}{m}L(k)$
$\leq\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{2}m}}.$
Recall that
$A_{pr}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}^{T}\bm{x}_{p}\geq 0\}\}$
and
$P(A_{p,r})\leq\frac{R}{\sqrt{2\pi}},\ S_{p}=\{r\in[m]:I\{A_{p,r}\}=0\}.$
Thus, from Bernstein’s inequality, we have that with probability at least $1-e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{p}^{\perp}\}=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{pr}\}\lesssim 4R.$
Then the inequality holds for all $p\in[n_{1}]$ with probability at least $1-n_{1}e^{-mR}$. Thus from (78), we have that for every $p\in[n_{1}]$
$|I_{1}^{p}(k)|\lesssim\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{1}m}}+\frac{\eta B^{4}\sqrt{L(k)}}{\sqrt{n_{1}}}R.$ (80)
Combining (79) and (80), we have that
$\|\bm{I}_{1}(k)\|_{2}$ $=\sqrt{\sum\limits_{p=1}^{n_{1}}|I_{1}^{p}(k)|^{2}+\sum\limits_{j=1}^{n_{2}}|I_{1}^{n_{1}+j}(k)|^{2}}$
$\lesssim\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{m}}+\eta B^{4}\sqrt{L(k)}R.$
Plugging it into (69) yields that
$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}$
$\leq[(1-\frac{\eta\lambda_{0}}{2})^{2}+C^{2}(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)^{2}+2C(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)]\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2},$
where the last inequality requires that
$\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}\lesssim\eta\lambda_{0},\ \eta B^{4}R\lesssim\eta\lambda_{0}.$
Recall that we also require that $CM^{2}R<\frac{\lambda_{0}}{4}$ and
$R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}<R.$
Finally, with $R=\mathcal{O}(\frac{\lambda_{0}}{B^{4}})$ and Lemma 11 for the upper bound of $L(0)$, $m$ needs to satisfies that
$m=\Omega\left(\frac{B^{12}L(0)}{\lambda_{0}^{4}}\right)=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
8. Proof of Section 4
8.1. Proof of Lemma 7
Proof.
Recall that
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}x_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}$
$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}]$
and
$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j}.$
(1) For $p\in[n_{1}]$, define the event
$A_{p,r}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}\}.$
Similar to (44), we can deduce that for any $p\in[n_{1}]$,
$P(A_{p,r})\leq\frac{2R}{\sqrt{2\pi}}.$
From the form of $\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$, we can deduce that
$\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}$ (81)
$\lesssim\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+|I\{\bm{w}_{r}^{T}\bm{x}_{p}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}|(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right]$
$\leq\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right],$
where the second inequality follows from the fact $\|\bm{w}-\bm{w}_{r}(0)\|_{2}<R\leq 1$ and the definition of $A_{p,r}$.
Similarly, we have that
$\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{2}m}}R(\|\bm{w}_{r}(0)\|_{2}+1).$ (82)
Combining (81) and (82), we can deduce that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}^{2}$
$\leq\|\bm{J}(\bm{w})-\bm{J}(0)\|_{F}^{2}$
$=\sum\limits_{i=1}^{n_{1}+n_{2}}\|\bm{J}_{i}(\bm{w})-\bm{J}_{i}(0)\|_{2}^{2}$
$=\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}\right)$
$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}(R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1))^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
$=\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)$
$\quad+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}\left(\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}+1\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{M^{2}}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\},$
where $M=2(d+2)\log(2m(d+2)/\delta)$. Note that from (62), we have
$P\left(\exists r\in[m],\|\bm{w}_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)\right)\leq\delta.$
On the other hand, applying Bernstein’s inequality yields that with probability at least $1-n_{1}e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{p,r}\}<4R$
holds for all $p\in[n_{1}]$.
Therefore, we have that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}^{2}\lesssim MR^{2}+R^{2}+M^{2}R\lesssim M^{2}R$
holds with probability at least $1-\delta-n_{1}e^{-mR}$.
(2) Note that when $\sigma$ satisfies the assumption 5, $\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}$ are all Lipschitz continuous. Thus we can obtain that
$\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}R(\|\bm{w}_{r}(0)\|_{2}^{2}+\|\bm{w}_{r}(0)\|_{2}+1)\lesssim\frac{1}{\sqrt{n_{1}m}}R(\|\bm{w}_{r}(0)\|_{2}^{2}+1),$ (83)
where the second inequality is from Young’s inequality.
Similarly, we have
$\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{2}m}}R(\|\bm{w}_{r}(0)\|_{2}+1).$ (84)
Combining (83) and (84) yields that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}^{2}$
$\leq\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}\right)$
$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}(R\|\bm{w}_{r}(0)\|_{2}^{2}+R)^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
$\lesssim R^{2}\left[d^{2}+\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2}\right],$
where the last inequality follows from the fact that $\|\|\bm{w}_{r}(0)\|_{2}^{4}\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}$ and Lemma 8. ∎
8.2. Proof of Theorem 3
We prove Theorem 3 by induction. Our induction is the following condition for the hidden weights.
Condition 3.
At the $t$-th iteration, we have $\|\bm{w}_{r}(t)\|_{2}\leq B$ and
$\|\bm{w}_{r}(t)-\bm{w}_{r}(0)\|_{2}\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}}$
for all $r\in[m]$, where $C$ is a universal constant and $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1.$
Instead of inducing on the convergence rate of the empirical loss function, as shown in Condition 2, we perform induction on the movements of the hidden weights. Because with Condition 3, we can directly derive the following convergence rate of the empirical loss function.
Corollary 3.
If Condition 3 holds for $t=0,\cdots,k$ and $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$, then
$L(t)\leq(1-\eta)^{t}L(0),$
holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 7 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ in (38) when $\sigma$ satisfies Asssumption 5.
Thanks to Corollary 3, it is sufficient to prove that Condition 3 also holds for $t=k+1$. For readability, we defer the proof of Corollary 3 to the end of this section. In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary 3 leads to Theorem 3.
Proof of Theorem 3.
We let $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function and let $R^{{}^{\prime\prime}}=CdR$ in (37) when $\sigma$ satisfies Assumption 5.
First, set $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$, then from Lemma 7 we have $\|\bm{J}(t)-\bm{J}(0)\|_{2}\leq\frac{\sqrt{\lambda_{0}}}{3}$, thus $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and $\lambda_{min}(\bm{H}(t))\geq\frac{4\lambda_{0}}{9}$ for $t=0,\cdots,k$.
From the updating role of NGD, we have
$\bm{w}_{r}(t+1)=\bm{w}_{r}(t)-\eta\left[\bm{J}(t)^{T}\right]_{r}(\bm{H}(t))^{-1}\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix},$
where
$\left[\bm{J}(t)^{T}\right]_{r}=\left(\frac{\partial s_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial s_{n_{1}}(t)}{\partial\bm{w}_{r}},\frac{\partial h_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial h_{n_{2}}(t)}{\partial\bm{w}_{r}}\right).$
Therefore, for $t=0,\cdots,k$ and any $r\in[m]$, we have
$\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}$ (85)
$\leq\eta\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\|\bm{H}(t)^{-1}\|_{2}\sqrt{L(t)}$
$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\sqrt{L(t)}$
$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{F}\sqrt{L(t)}$
$=\frac{9\eta}{4\lambda_{0}}\sqrt{\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}}\sqrt{L(t)}$
$\lesssim\frac{4\eta}{9\lambda_{0}}\sqrt{\frac{B^{4}+1}{m}}\sqrt{L(t)}$
$\lesssim\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
$\leq\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}(1-\eta)^{t/2}\sqrt{L(0)},$
where the last inequality is due to Corollary 3.
Summing $t$ from 00 to $k$ yields that
$\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}$
$\leq\sum\limits_{t=0}^{k}\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}$
$\leq C\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}\sum\limits_{t=0}^{k}(1-\eta)^{t/2}\sqrt{L(0)}$
$\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}.$
Note that although $C$ is a universal constant, it still depends on the activation function.
Now, when $R^{{}^{\prime}}\leq 1$, we can deduce that Condition 3 also holds for $t=k+1$. Thus, it remains only to derive the requirement for $m$.
Recall that we need $m$ to satisfy that $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$.
(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CM\sqrt{R}\leq\frac{\sqrt{\lambda_{0}}}{3}$, wihch implies that $R\lesssim\frac{\lambda_{0}}{M^{2}}$. Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ imples that
$m=\Omega\left(\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right).$
From Lemma 11, we can deduce that
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
(2) When $\sigma$ satisfies Assumption 5, we have that
$R\lesssim\frac{\sqrt{\lambda_{0}}}{d},R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R.$
From Lemma 5 in [23], we know
$L(0)\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right).$
Thus, we can deduce that
$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
Proof of Corollary 3.
Similar as before, when $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$ , we have $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and then $\lambda_{min}(\bm{H}(t))\geq\frac{4}{9}\lambda_{0}$ for $t=0,\cdots,k$.
Let $\bm{u}(t)=\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}$, then
$\bm{u}(t+1)-\bm{u}(t)$ (86)
$=\bm{u}\left(\bm{w}(t)-\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(\bm{w}(t))\right)-\bm{u}(\bm{w}(t))$
[…]
$:=\bm{I}_{1}(t)+\bm{I}_{2}(t),$
where the second equality is from fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging it into (85) yields that
$\bm{u}(t+1)=(1-\eta)\bm{u}(t)+\bm{I}_{2}(t).$ (87)
Therefore, it remains only to bound $\|\bm{I}_{2}(t)\|_{2}$.
[…]
$\lesssim\frac{\eta\sqrt{L(t)}}{\sqrt{\lambda_{0}}}R^{{}^{\prime\prime}}$
where the last inequality follows from the fact that
$\|\bm{w}_{r}(s)-\bm{w}_{r}(0)\|_{2}\leq s\|\bm{w}_{r}(t+1)-\bm{w}_{r}(0)\|_{2}+(1-s)\|\bm{w}_{r}(t)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$
and Lemma 7.
Plugging (88) into the recursion formula (87) yields that
$\|\bm{u}(t+1)\|_{2}^{2}$ $=\|(1-\eta)\bm{u}(t)+\bm{I}_{2}(t)\|_{2}^{2}$
$=(1-\eta)^{2}\|\bm{u}(t)\|_{2}^{2}+\|\bm{I}_{2}(t)\|_{2}^{2}+2\langle(1-\eta)\bm{u}(t),\bm{I}_{2}(t)\rangle$
$\leq(1-\eta)^{2}\|\bm{u}(t)\|_{2}^{2}+\|\bm{I}_{2}(t)\|_{2}^{2}+2(1-\eta)\|\bm{u}(t)\|_{2}\|\bm{I}_{2}(t)\|_{2}$
$\leq\left[(1-\eta)^{2}+\frac{C^{2}\eta^{2}(R^{{}^{\prime\prime}})^{2}}{\lambda_{0}}+2(1-\eta)\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\right]\|u(t)\|_{2}^{2}$
$\leq(1-\eta)\|\bm{u}(t)\|_{2}^{2},$
where the last inequality follows from the requirement
$2\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\leq\frac{\eta}{4},$
i.e., $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$.
Therefore, we can deduce that $\|\bm{u}(t)\|_{2}^{2}\leq(1-\eta)^{t}\|\bm{u}(0)\|_{2}^{2}$ holds for $t=0,\cdots,k$.
8.3. Proof of Corollary 1
Proof.
Setting $\eta=1$ in (85) yields that
$\bm{u}(t+1)=\bm{I}_{2}(t).$
From (88), we have that
[…]
As $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)$, then for any $r\in[m]$, we have $\|\bm{w}_{r}(s)\|_{2}\leq s\|\bm{w}_{r}(t+1)\|_{2}+(1-s)\|\bm{w}_{r}(t)\|_{2}\leq B$.
When $\sigma$ is smooth, we can deduce that for any $r\in[m]$,
$\left\|\frac{\partial s_{p}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}(B^{2}+1)\|\bm{w}_{r}(s)-\bm{w}_{r}(t)\|_{2}\leq\frac{1}{\sqrt{n_{1}m}}(B^{2}+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}$
and
$\left\|\frac{\partial h_{j}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}(B+1)\|\bm{w}_{r}(s)-\bm{w}_{r}(t)\|_{2}\leq\frac{1}{\sqrt{n_{1}m}}(B+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}.$
From (85), we know that for any $r\in[m]$,
$\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}\lesssim\frac{B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}.$
Thus for any $s\in[0,1]$, we have
$\|\bm{J}(\bm{w}(s))-\bm{J}(\bm{w}(t))\|_{2}^{2}$
$\leq\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\left\|\frac{\partial h_{j}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}^{2}\right)$
$\lesssim\frac{1}{m}\sum\limits_{r=1}^{m}\left((B^{4}+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}^{2}+(B^{2}+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}^{2}\right)$
$\lesssim B^{4}\left(\frac{B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}\right)^{2}.$
Plugging it into (89), we have
[…]
$\lesssim\frac{\sqrt{L(t)}}{\sqrt{\lambda_{0}}}\frac{\eta B^{4}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
$=\frac{B^{4}}{\sqrt{m\lambda_{0}^{3}}}L(t).$
Combining with the fact $\bm{u}(t+1)=\bm{I}_{2}(t)$ yields that
$\left\|\begin{pmatrix}\bm{s}(t+1)\\ \bm{h}(t+1)\end{pmatrix}\right\|_{2}\leq\frac{CB^{4}}{\sqrt{m\lambda_{0}^{3}}}\left\|\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}\right\|_{2}^{2}$
holds for $t\in\mathbb{N}$, where $C$ is a universal constant. ∎
9. Auxiliary Lemmas
Lemma 8 (Theorem 3.1 in [24]).
If $X_{1},\cdots,X_{n}$ are independent mean zero random variables with $\|X_{i}\|_{\psi_{\alpha}}<\infty$ for all $1\leq i\leq n$ and some $\alpha>0$, then for any vector $a=(a_{1},\cdots,a_{n})\in\mathbb{R}^{n}$, the following holds ture:
$P(|\sum\limits_{i=1}^{n}a_{i}X_{i}|\geq 2eC(\alpha)\|b\|_{2}\sqrt{t}+2eL_{n}^{*}(\alpha)t^{1/\alpha}\|b\|_{\beta(\alpha)})\leq 2e^{-t},\ for\ all\ t\geq 0,$
where $b=(a_{1}\|X_{1}\|_{\psi_{\alpha}},\cdots,a_{n}\|X_{n}\|_{\psi_{\alpha}})\in\mathbb{R}^{n}$,
$C(\alpha):=\max\{\sqrt{2},2^{1/\alpha}\}\left\{\begin{aligned} \sqrt{8}(2\pi)^{1/4}e^{1/24}(e^{2/e}/\alpha)^{1/\alpha}&,&if\ \alpha<1,\\ 4e+2(\log 2)^{1/\alpha}&,&if\ \alpha\geq 1.\end{aligned}\right.$
and for $\beta(\alpha)=\infty$ when $\alpha\geq 1$ and $\beta(\alpha)=\alpha/(\alpha-1)$ when $\alpha>1$,
$L_{n}(\alpha):=\frac{4^{1/\alpha}}{\sqrt{2}\|b\|_{2}}\times\left\{\begin{aligned} &\|b\|_{\beta(\alpha)},&if\ \alpha<1,\\ &4e\|b\|_{\beta(\alpha)}/C(\alpha),&if\alpha\geq 1.\end{aligned}\right.$
and $L^{*}_{n}(\alpha)=L_{n}(\alpha)C(\alpha)\|b\|_{2}/\|b\|_{\beta(\alpha)}$.
Lemma 9 (Bernstein inequality, Theorem 3.1.7 in [25]).
Let $X_{i}$, $1\leq i\leq n$ be independent centered random variables a.s. bounded by $c<\infty$ in absolute value. Set $\sigma^{2}=1/n\sum_{i=1}^{n}\mathbb{E}X_{i}^{2}$ and $S_{n}=1/n\sum_{i=1}^{n}X_{i}$. Then, for all $t\geq 0$,
$P\left(S_{n}\geq\sqrt{\frac{2\sigma^{2}t}{n}}+\frac{ct}{3n}\right)\leq e^{-u}.$
Lemma 10.
For $10<\delta<$ 1, with probability at least $1-\delta$, we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}=\mathcal{O}\left(n\log\left(\frac{n}{\delta}\right)\right).$
Proof.
First, from Cauchy’s inequality, we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ $=\sum\limits_{i=1}^{n}\left(y_{i}-\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
$\leq\sum\limits_{i=1}^{n}2y_{i}^{2}+2\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
$=2\sum\limits_{i=1}^{n}y_{i}^{2}+2\sum\limits_{i=1}^{n}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}.$
Since $\bm{w}_{r}(0)^{T}\bm{x}_{i}\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2})$ and $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$, we can deduce that
$\|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\|_{\psi_{2}}\leq\|\bm{w}_{r}(0)^{T}\bm{x}_{i}\|_{\psi_{2}}=\mathcal{O}(1)$
holds for all $r\in[m]$ and $i\in[n]$.
Thus for any fixed $i\in[n]$, applying Lemma 8 yields that with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right|\leq C\sqrt{t},$
where $C$ is a universal constant. By taking union bound on all $i\in[n]$, the above inequality holds for all $i\in[n]$ with probability at least $1-2ne^{-t}$.
Let $2ne^{-t}=\delta$, i.e., $t=\log\left(\frac{2n}{\delta}\right)$, then we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}\lesssim n+nt\lesssim n\log\left(\frac{n}{\delta}\right).$
Lemma 11.
For $10<\delta<$ 1, with probability at least $1-\delta$, we have that when $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$,
$L(0)=\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}=\mathcal{O}\left(d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
Proof.
Recall that for $p\in[n_{1}]$,
$s_{p}(0)=\frac{1}{\sqrt{n_{1}}}\left[\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)-f(x_{p})\right]$
and for $j\in[n_{2}]$,
$h_{j}(0)=\frac{1}{\sqrt{n_{2}}}\left[\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})-g(\bm{y}_{j})\right].$
Then
$L(0)$ $=\sum\limits_{p=1}^{n_{1}}\frac{1}{2}(s_{p}(0))^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{2}(h_{j}(0))^{2}$
$\leq\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right)+\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}f^{2}(x_{p})$
$\quad+\frac{1}{n_{2}}\sum\limits_{j=1}^{n_{2}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right)^{2}+\frac{1}{n_{2}}\sum\limits_{j=1}^{n_{2}}g^{2}(\bm{y}_{j}).$
Note that
$|a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}$
and $|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}.$
Since $\|\|\bm{w}_{r}(0)\|_{2}^{3}\|_{\psi_{\frac{2}{3}}}\leq(\|\|\bm{w}_{r}(0)\|_{2}^{2}\|_{\psi_{1}})^{\frac{3}{2}}\lesssim d^{3}$, from Lemma 8, we have that for fixed $i\in[n_{1}]$ and $j\in[n_{2}]$ with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}$
and with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}.$
Then taking a union bound for all $i\in[n_{1}]$ and $j\in[n_{2}]$ with $2(n_{1}+n_{2})e^{-t}=\delta$ yields that
$L(0)$ $\lesssim(d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}})^{2}$
$\lesssim d^{6}t+\frac{d^{6}t^{3}}{m}$
$=d^{6}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{1}{m}\log^{3}\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
$\lesssim d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
as $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$.

### v3

@@
 Abstract.
-First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD) have been proven effective in training neural networks.
-In the setting of over-parameterization, there is a line of work demonstrating that randomly initialized (stochastic) gradient descent converges to a globally optimal solution at a linear convergence rate for the quadratic loss function.
-However, the learning rate of GD in training two-layer neural networks has a poor dependence on the sample size and the Gram matrix, resulting in a slow training process.
-In this paper, we show that for the $L^{2}$ regression problems, the learning rate can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$ to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, which implies that GD enjoys a faster convergence rate.
-Moreover, we further generalize the method for GD in training two-layer Physics-Informed Neural Networks (PINNs), showing a similar improvement for the learning rate.
-Although the improved learning rate depends mildly on the Gram matrix, we still need to set it small enough in practice due to the agnostic eigenvalues of the Gram matrix.
-More importantly, the convergence rate relies on the least eigenvalue of the Gram matrix, leading to slow convergence.
-In this work, we provide the convergence analysis of natural gradient descent (NGD) in training two-layer PINNs.
-We show that the learning rate can be $\mathcal{O}(1)$ and at this time, the convergence rate is independent of the Gram matrix.
+In the context of over-parameterization, there is a line of work demonstrating that randomly initialized (stochastic) gradient descent (GD) converges to a globally optimal solution at a linear convergence rate for the quadratic loss function.
+However, the learning rate of GD for training two-layer neural networks exhibits poor dependence on the sample size and the Gram matrix, leading to a slow training process.
+In this paper, we show that for training two-layer $\text{ReLU}^{3}$ Physics-Informed Neural Networks (PINNs), the learning rate can be improved from $\mathcal{O}(\lambda_{0})$ to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, implying that GD actually enjoys a faster convergence rate.
+Despite such improvements, the convergence rate is still tied to the least eigenvalue of the Gram matrix, leading to slow convergence.
+We then develop the positive definiteness of Gram matrices with general smooth activation functions and provide the convergence analysis of natural gradient descent (NGD) in training two-layer PINNs, demonstrating that the learning rate can be $\mathcal{O}(1)$ and at this rate, the convergence rate is independent of the Gram matrix.
+In particular, for smooth activation functions, the convergence rate of NGD is quadratic.
+Numerical experiments are conducted to verify our theoretical results.
 * Corresponding author.
 1. Introduction
-In recent years, neural networks have achieved remarkable breakthroughs in the fields of image recognition [1], natural language procssing [2], reinforcement learning [3] and so on.
-Moreover, due to the flexibility and scalability of neural networks, researchers are paying much attention in exloring new methods involving neural networks for handling problems in scientific computing, where one long-standing and essential problem is solving partial differntial equations (PDEs) numerically.
-Classical numerical methods, such as finite difference, finite volume and finite elements methods, suffer the curse of dimensionality in solving high-dimensional PDEs.
+In recent years, neural networks have achieved remarkable breakthroughs in the fields of image recognition [1], natural language processing [2], reinforcement learning [3], and so on.
+Moreover, due to the flexibility and scalability of neural networks, researchers are paying much attention in exploring new methods involving neural networks for handling problems in scientific computing.
+One long-standing and essential problem in this area is solving partial differential equations (PDEs) numerically.
+Classical numerical methods, such as finite difference, finite volume and finite elements methods, suffer from the curse of dimensionality when solving high-dimensional PDEs.
 Due to this drawback, various methods involving neural networks have been proposed for solving different type PDEs [4, 5, 6, 7, 8].
@@
 With the use of automatic differentiation, the neural network can be efficiently trained by first-order or second-order methods.
-In the applications of neural networks, one inevitable issue is the selection of the optimization method.
-First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD), are widely used in optimizing neural networks as they only calculate the gradient and therefore are computationally efficient.
-Beside the first-order methods, there has been much interest in utilizing second-order optimization methods to speed up the training, not only for the regression problems [9] but also for the problems related to PDEs [4, 5].
-For the convergence part of the optimization method, it has been shown that gradient descent algorithm can even achieve zero training loss under the setting of over-parametrization, which refers to a situation where a model has more parameters than necessary to fit the data [10, 11, 12, 13, 14, 15].
-These works are based on the idea of neural tangent kernel (NTK), which shows that training multi-layer fully-connected neural networks via gradient descent is equivalent to performing a certain kernel method as the width of every layer goes to infinity.
-As for the finite width neural networks, with more refined analysis, it can be shown that the parameters are closed to the initialization throughout the entire training process when the width is large enough.
-It directly leads to the linear convergence for GD.
-Despite these beautiful convergence results, the learning rate depends on the sample size and the Gram matrix.
-Thus, it need to be sufficiently small to guarantee the convergence in practice.
-But doing so results in a slow training process.
-In contrast to the first-order methods, the second-order methods enjoy fast convergence due to the absence of requirements for the learning rate [16].
+In the applications of neural networks, one inevitable issue is the selection of the optimization methods.
+First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD), are widely used in optimizing neural networks as they only calculate the gradient, making them computationally efficient.
+In addition to first-order methods, there has been significant interest in utilizing second-order optimization methods to accelerate training.
+These methods have proven to be applicable not only to regression problems, as demonstrated in [9], but also to problems related to PDEs, as shown in [4, 5].
+As for the convergence aspect of the optimization methods, it has been shown that gradient descent algorithm can even achieve zero training loss under the setting of over-parameterization, which refers to a situation where a model has more parameters than necessary to fit the data [10, 11, 12, 13, 14, 15, 16, 17].
+These works are based on the idea of neural tangent kernel (NTK)[18], which shows that training multi-layer fully-connected neural networks via gradient descent is equivalent to performing a certain kernel method as the width of every layer goes to infinity.
+As for the finite width neural networks, with more refined analysis, it can be shown that the parameters are closed to the initializations throughout the entire training process when the width is large enough.
+This directly leads to the linear convergence for GD.
+Despite these attractive convergence results, the learning rate depends on the sample size and the Gram matrix, so it needs to be sufficiently small to guarantee convergence in practice.
+However, doing so results in a slow training process.
+In contrast to first-order methods, the second-order method natural gradient descent (NGD) has been shown to enjoy fast convergence for the $L^{2}$ regression problems, as demonstrated in [19, 20].
+However, the convergence of NGD in the context of training PINNs is still an open problem.
+In this paper, we demonstrate that when training PINNs, NGD indeed enjoys a faster convergence rate.
 1.1. Contributions
 The main contributions of our work can are summarized as follows:
-  • For the $L^{2}$ regression problems, we demonstrate that the learning rate $\eta$ of gradient descent can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$, as shown in [10], to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the ReLU activation function and the random initialization, $\lambda_{0}$ is the least eigenvalue of $\bm{H}^{\infty}$.
-[17] has also shown the same improvement of the learning rate, but it requires that $m=\Omega(\frac{n^{6}}{\lambda_{0}^{4}\delta^{3}})$.
-Moreover in [18], the dependence on $n$, i.e.
-$\Omega(n^{6})$, is necessary and can not be improved due to the requirement of the proof method.
-Different from the method in [17], our method comes from a new recursion formula of gradient descent and it can be easily generalized to PINNs.
-Compared to [17], we only require that $m=\Omega(\frac{n^{4}}{\lambda_{0}^{4}}(\log(\frac{n}{\delta}))^{2})$.
-  • For the PINNs, we simultaneously improve both the learning rate $\eta$ of gradient descent and the requirement of the width $m$.
-The improvements rely on a new recursion formula of gradient descent, which is similar to that for regression problems.
+  • For the PINNs, we simultaneously improve both the learning rate $\eta$ of gradient descent and the requirement for the width $m$.
+The improvements rely on a new recursion formula for gradient descent, which is similar to that for regression problems.
 Specifically, we can improve the learning rate $\eta=\mathcal{O}(\lambda_{0})$ required in [21] to $\eta=\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$ and the requirement for the width $m$, i.e.
-$m=\tilde{\Omega}(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}})$, can be improved to $m=\tilde{\Omega}(\frac{1}{\lambda_{0}^{4}}(\log(\frac{n_{1}+n_{2}}{\delta})))$.
+$m=\widetilde{\Omega}\left(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}}\right)$, can be improved to $m=\widetilde{\Omega}\left(\frac{1}{\lambda_{0}^{4}}(\log(\frac{n_{1}+n_{2}}{\delta}))\right)$, where $\widetilde{\Omega}$ indicates that some terms involving $\log(m)$ are omitted.
+  • We present a framework for demonstrating the positive definiteness of Gram matrices for a variety of commonly used smooth activation functions, including the logistic function, softplus function, hyperbolic tangent function, and others.
+This conclusion is not only applicable to the PDE we have considered but can also be naturally extended to other forms of PDEs.
   • We provide the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions.
-Compared to gradient descent, the convergence rate of NGD is independent of $\lambda_{0}$, which results in a faster convergence.
-Meanwhile, we show that the learning rate can be even $\mathcal{O}(1)$.
-When the activation function is smooth, NGD enjoys quadratic convergence rate.
+Due to the distinct optimization dynamics of NGD compared to GD, the learning rate can be $\mathcal{O}(1)$.
+Consequently, the convergence rate is independent of $n$ and $\lambda_{0}$, leading to faster convergence.
+Moreover, when the activation function is smooth, NGD can achieve a quadratic convergence rate.
 1.2. Related Works
-First-order methods.
-There are mainly two ways to study the optimization of neural networks and try to understand why first-order methods can find a global minimum.
-One way is to analyse the optimization landscape [15, 20].
-It has been shown that gradient descent can find a global minimum in polynomial time if the optimization landscape has some good geometric properties.
-However, some unrealistic assumptions in these works make it difficult to generalize to practical neural networks.
-Another way to understand the optimization of neural networks is to analyze the optimization dynamics of first-order methods.
-For the two-layer ReLU neural networks, as shown in [10], as long as the width $m$ is large enough and no two inputs are parallel, randomly initialized gradient descent converges to a globally optimal solution at a linear rate.
-Later, the results have been generalized to deep neural networks with smooth activation functions [11].
-Both rely on that the Gram matrices are stable throughout the training process, which leads to the convergence to the global minimum.
-Besides regression and classification problems, [19] demonstrated the convergence of the gradient descent for two-layer PINNs by the similar analysis of optimization dynamics.
-However, both [10] and [19] require that the learning rate be sufficiently small to converge.
-In this work, we conduct a refined analysis of gradient descent for $L^{2}$ regression problems and PINNs, leading to a milder requirement for the learning rate.
-Second-order methods.
-Although second-order methods possess better convergence rate, they are rarely used in training deep neural networks due to the prohibitive computational cost.
-As a variant of the original Gauss-Newton method, natural gradient descent (NGD) is more efficient in practice.
-Meanwhile, as shown in [21] and [16], it also enjoys faster convergence rate for the $L^{2}$ regression problems than of gradient descent.
-In this paper, we provide the convergence analysis for NGD in training two-layer PINNs.
+First-order optimizers.
+There are mainly two approaches to studying the optimization of neural networks and understanding why first-order methods can find a global minimum.
+One approach is to analyze the optimization landscape, as demonstrated in [22, 23].
+It has been shown that gradient descent can find a global minimum in polynomial time if the optimization landscape possesses certain favorable geometric properties.
+However, some unrealistic assumptions in these works make it challenging to generalize the findings to practical neural networks.
+Another approach to understand the optimization of neural networks is by analyzing the optimization dynamics of first-order methods.
+For the two-layer ReLU neural networks, as shown in [10], randomly initialized gradient descent converges to a globally optimal solution at a linear rate, provided that the width $m$ is sufficiently large and no two inputs are parallel.
+Later, these results were extended to deep fully-connected feedforward neural networks and ResNet with smooth activation functions [11].
+Results for both shallow and deep neural networks depend on the stability of the Gram matrices throughout the training process, which is crucial for convergence to the global minimum.
+In addition to regression and classification problems, [21] demonstrated the convergence of the gradient descent for two-layer PINNs through a similar analysis of optimization dynamics.
+However, both [10] and [21] require a sufficiently small learning rate and a large enough network width to achieve convergence.
+In this work, we conduct a refined analysis of gradient descent for PINNs, resulting in milder requirements for the learning rate and network width.
+[…]
 1.3. Notations
@@
 We use $I\{E\}$ to denote the indicator function of the event $E$.
-For variables $a$ and $b$, $a\lesssim b$ is equivalent to $a\leq Cb$ for some universal constant $C$.
-For two functions $f_{1}(n)$ and $f_{2}(n)$, the relation $f_{1}(n)=\mathcal{O}(f_{2}(n))$ is equivalent to $f_{1}(n)\leq Cf_{2}(n)$ for some universal constant $C$.
-And similarly, $f_{1}(n)=\Omega(f_{2}(n))$ is equivalent to $f_{1}(n)\geq Cf_{2}(n)$.
-Throughout the paper, we use boldface denote vectors.
+For two positive functions $f_{1}(n)$ and $f_{2}(n)$, we use $f_{1}(n)=\mathcal{O}(f_{2}(n))$, $f_{2}(n)=\Omega(f_{1}(n))$ or $f_{1}(n)\lesssim f_{2}(n)$ to represent $f_{1}(n)\leq Cf_{2}(n)$, where $C$ is a universal constant $C$.
+A universal constant means a constant independent of any variables.
+Throughout the paper, we use boldface to denote vectors.
 Given $x_{1},\cdots,x_{d}\in\mathbb{R}$, we use $(x_{1},\cdots,x_{d})$ or $[x_{1},\cdots,x_{d}]$ to denote a row vector with $i$-th component $x_{i}$ for $i\in[d]$ and then $(x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d}$ is a column vector.
 1.4. Organization of this Paper
-We present the improvements of the learning rate of gradient descent for $L^{2}$ regression problems and PINNs in Section 2 and Section 3 respectively.
-In Section 4, we show the convergence of natural gradient descent for training PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions.
-We conclude in Section 5 and the detailed proofs are put in the Appendix for readability and brevity.
-2. Improved Learning Rate of Gradient Descent for $L^{2}$ Regression Problems
-2.1. Problem Setup
-In this section, we consider a two-layer neural network $f$ with the following form.
-$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\tilde{\bm{x}}),$ (1)
-where $\bm{x}\in\mathbb{R}^{d}$ is the input, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$, $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T},\bm{a}=(a_{1},\cdots,a_{m})^{T}$, for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+1}$ is the weight vector of the first layer, $a_{r}\in\mathbb{R}$ is the output weight and $\sigma(\cdot)$ is the ReLU activation function.
-Different from the setup in [10], the two-layer neural network that we consider has bias term.
-In the following, we assume that each input vector $\bm{x}\in\mathbb{R}^{d}$ has been augmented to $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$.
-With a little abuse of notation, we write $\bm{x}$ for $\tilde{\bm{x}}$ and write $f$ as
-$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x}).$
-For the $L^{2}$ regression problem, given training data set $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, the aim is to minimize the loss function
-$L(\bm{w},\bm{a}):=\sum\limits_{i=1}^{n}\frac{1}{2}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})^{2}.$ (2)
-Before training, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$ and output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$.
-In the training process, we fix the second layer and optimize the first layer by gradient descent (GD), i.e., we update the weights by the following formulation.
-$\bm{w}(k+1)=\bm{w}(k)-\eta\frac{\partial L(\bm{w}(k))}{\partial\bm{w}},$ (3)
-where $\eta>0$ is the learning rate, $k\in\mathbb{N}$ and $L(\bm{w}(k))$ is an abbreviation of $L(\bm{w}(k),\bm{a})$. From the form of $L(\bm{w})$, we can deduce that
-$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{m}}\sum\limits_{i=1}^{n}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})\bm{x}_{i}I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}.$ (4)
-We denote $u_{i}(k)=f(\bm{x}_{i};\bm{w}(k),\bm{a})$ the prediction on input $\bm{x}_{i}$ at $k$-th iteration and we let $\bm{u}(k)=(u_{1}(k),\cdots,u_{n}(k))^{T}\in\mathbb{R}^{n}$ be the prediction vector at $k$-th iteration.
-For the labels, we define $\bm{y}:=(y_{1},\cdots,y_{n})^{T}\in\mathbb{R}^{n}$.
-According to [10], in the continuous setting, the dynamics of predictions can be written as
-$\frac{d}{dt}\bm{u}(t)=\bm{H}(t)(\bm{y}-\bm{u}(t)),$ (5)
-where $H(t)\in\mathbb{R}^{n\times n}$ is the Gram matrix at time $t$ with $(i,j)$-entry
-$H_{ij}(t)=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}(t)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(t)^{T}\bm{x}_{j}\geq 0\right\}.$ (6)
-In the setting of over-parameterization and randomly initialization, [10] has shown that (1) $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(\sqrt{1/m})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the initialization with $(i,j)$-th entry
-$H_{ij}^{\infty}=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}^{T}\bm{x}_{i}\geq 0,\bm{w}^{T}\bm{x}_{j}\geq 0\right\}\right]$ (7)
-and (2) $\|\bm{H}(t)-\bm{H}(0)\|_{2}=\mathcal{O}(\sqrt{1/m})$ for all $t>0$.
-Therefore, as $m\to\infty$, the dynamics of the predictions are charactered by $H^{\infty}$, leading to the linear convergence.
-2.2. Main Results
-To simplify the analysis, we make the following assumptions on the training data.
-Assumption 1.
-For $i\in[n]$, $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$ and $|y_{i}|\leq 1$, where $\bm{x}_{i}\in\mathbb{R}^{d+1}$ is the augmented input.
-Assumption 2.
-No two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are parallel, i.e., for any $\bm{x}_{t},\bm{x}_{s}\in\{\bm{x}_{i}\}_{i=1}^{n}$ and any $\alpha\in\mathbb{R}$, we have $\bm{x}_{t}\neq\bm{x}_{s}$.
-Since inputs are all augmented, Assumption 2 is equivalent to that no two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are equal, which holds naturally.
-Under Assumption 2, Theorem 3.1 in [10] implies that $\lambda_{0}:=\lambda_{min}(\bm{H}^{\infty})>0$, which is crucial in the convergence analysis.
-As stated before, there are two improtant facts that determine the optimization dynamics, one is that at initialization $\bm{H}(0)$ is closed to $\bm{H}^{\infty}$ and another is that $\bm{H}(k)$ does not go far away from the initialization $\bm{H}(0)$ for all $k\in\mathbb{N}$.
-These facts are supported by the following two lemmas.
-Lemma 1 (Lemma 3.1 in [10]).
-If $m=\Omega\left(\frac{n^{2}}{\lambda_{0}^{2}}\log\left(\frac{n}{\delta}\right)\right)$, we have with probability at least $1-\delta$, $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\frac{\lambda_{0}}{4}$ and $\lambda_{min}(\bm{H}(0))\geq\frac{3\lambda_{0}}{4}$.
-Lemma 2.
-Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d.
-generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-n^{2}e^{-mR/4}$, the following holds.
-For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then the matrix $\bm{H}(\bm{w})\in\mathbb{R}^{n\times n}$ defined by
-$H(\bm{w})_{ij}=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\right\}.$
-satisfies
-$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}<4nR.$ (8)
-Remark 1.
-Although Lemma 2 looks almost same to Lemma 3.1 in [18], the proof of Lemma 3.1 in [18] lacks an important part, i.e., the proof of the measurability for the related variable.
-In fact, Lemma 2 ensures us to bound the change of $\bm{H}(\bm{w})$ when $\bm{w}$ in a small ball.
-However, the set consisting of vectors in the small ball is uncountable, thus the related variable may not be measurable due to the discontinuity of the indicator function appeared in the definition of $\bm{H}(\bm{w})$.
-In the proof, we borrow the concept of pointwise measurability in the empirical process (see Chapter 8 in [22]) to bridge the gap.
-Then we can bound the supremum with a new random variable.
-By applying a concentration inequality to this new random variable, we can reach the conclusion.
-Abount the convergence of gradient descent, [10] has demonstrated that if the learning rate $\eta=\mathcal{O}(\lambda_{0}/n^{2})$, then randomly initialized gradient descent converges to a globally optimal solution at a linear convergence rate when $m$ is large enough.
-The requirement of $\eta$ is from the decomposition for the residual in the $(k+1)$-th iteration, i.e.,
-$\bm{y}-\bm{u}(k+1)=\bm{y}-\bm{u}(k)-(\bm{u}(k+1)-\bm{u}(k)).$ (9)
-Instead of decomposing the residual into the two terms as above, we write it as follows, which serves as a recursion formula.
-Lemma 3.
-For all $k\in\mathbb{N}$, we have
-$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta H(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k),$ (10)
-where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n}(k))^{T}\in\mathbb{R}^{n}$ with $i$-th
-$I_{1}^{i}(k):=u_{i}(k+1)-u_{i}(k)-\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$ (11)
-After [10], [17] improved the network size $m$ to $m=\mathcal{O}(\lambda_{0}^{-4}n^{4}\log^{3}(n/\delta)\log(m/\delta))$, but it still requires $\eta=\mathcal{O}(\lambda_{0}/n^{2})$.
-Another work [17] provided a refined analysis over [10], which shows that $\eta=\frac{1}{C_{1}\|\bm{H}^{\infty}\|_{2}}$ is enough for the convergence rate $1-\frac{\lambda_{0}C_{2}}{C_{1}\|\bm{H}{\infty}\|_{2}}$, but the constants $C_{1},C_{2}$ may depend on the parameters $\lambda_{0},n$ and $\delta$.
-It requires the network size $m$ to be $\Omega(\lambda_{0}^{-4}n^{6}\delta^{-3})$ to make $C_{1},C_{2}$ independent of $\lambda_{0},n$ and $\delta$.
-In this paper, we improve both the learning rate and network size and the main reslut is the following theorem.
-Theorem 1.
-Under Assumption 1 and Assumption 2, if we set the number of hidden nodes $m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}}\log\left(\frac{n}{\delta}\right)\right)$ and the learning rate $\eta=\mathcal{O}\left(\frac{1}{\|\bm{H}^{\infty}\|_{2}}\right)$, then with probability at least $1-\delta$ over the random initialization, the gradient descent algorithm satisfies
-$\|\bm{y}-\bm{u}(k)\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ (12)
-for all $k\in\mathbb{N}$.
-Remark 2.
-Our proof method requires that $\bm{I}-\eta\bm{H}(k)$ is positive definite for all $k\in\mathbb{N}$.
-As $\bm{H}(k)$ is closed to $\bm{H}^{\infty}$, the requirement is satisfied if $\eta\|\bm{H}^{\infty}\|_{2}\lesssim 1$.
-Note that $\|\bm{H}^{\infty}\|_{2}\leq n$, it is sufficient to set $\eta=\mathcal{O}(1/n)$, which results in an improvement of $\mathcal{O}(\lambda_{0}/n)$.
-3. Improved Learning Rate of Gradient Descent for Two-Layer Physics-Informed Neural Networks
-3.1. Problem Setup
-In this section, we consider the same setup as [19], i.e., we focus on the PDE with the following form.
+In Section 2, we provide the problem setup for training two-layer PINNs.
+We then present the improved convergence results of gradient descent for PINNs in Section 3.
+In Section 4, we analyze the convergence of natural gradient descent in training two-layer PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions.
+In Section 5, we conduct experiments to verify the theoretical results.
+The limitations are briefly discussed in Section 6 and we conclude in Section 7.
+All the detailed proofs and experiments are provided in the Appendix for readability and brevity.
+2. Problem Setup
+In this section, we consider the same setup as [21], focusing on the PDE with the following form.
 $&\frac{\partial u}{\partial x_{0}}(\bm{x})-\sum\limits_{i=1}^{d}\frac{\partial^{2}u}{\partial x_{i}^{2}}(\bm{x})=f(\bm{x}),\,\bm{x}\in(0,T)\times\Omega,\\ &u(\bm{x})=g(\bm{x}),\ \bm{x}\in\{0\}\times\Omega\cup[0,T]\times\partial\Omega,\end{aligned}\right.$ (1)
-where $\bm{x}=(x_{0},x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d+1}$ and $x_{0}\in[0,T]$ is the time variable.
+where $\Omega\subset\mathbb{R}^{d}$ is an open and bounded domain, $\bm{x}=(x_{0},x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d+1}$ and $x_{0}\in[0,T]$ is the time variable.
 In the following, we assume that $\|\bm{x}\|_{2}\leq 1$ for $\bm{x}\in[0,T]\times\bar{\Omega}$ and $f,g$ are bounded continuous functions.
@@
 where $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}\in\mathbb{R}^{m(d+2)}$, $\bm{a}=(a_{1},\cdots,a_{m})^{T}\in\mathbb{R}^{m}$ and for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+2}$ is the weight vector of the first layer, $a_{r}$ is the output weight and $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function.
-Similar as that in Section 2, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+2}$ is the augmented vector from $\bm{x}$ and in the following, we write $\bm{x}$ for $\tilde{\bm{x}}$ for brevity.
-In the framework of PINNs, given training samples $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ and $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ that are from interior and boundary respectively, we aim to minimize the following empirical loss function.
-$L(\bm{w},\bm{a})$ $:=\sum\limits_{p=1}^{n_{1}}\frac{1}{2n_{1}}\left(\frac{\partial\phi}{\partial x_{0}}(\bm{x}_{p};\bm{w},\bm{a})-\sum\limits_{i=1}^{d}\frac{\partial^{2}\phi}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w},\bm{a})-f(\bm{x}_{p})\right)^{2}$ (15)
-$\quad+\sum\limits_{j=1}^{n_{2}}\frac{1}{2n_{2}}\left(\phi(\bm{y}_{j};\bm{w},\bm{a})-g(\bm{y}_{j})\right)^{2}.$
-Similar as that for the $L^{2}$ regression problems, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$, output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$ and fix the output weights.
-Then the gradient descent updates the hidden weights by the following formulations:
-$\bm{w}_{r}(k+1)$ $=\bm{w}_{r}(k)-\eta\frac{\partial L(\bm{w}(k),\bm{a})}{\partial\bm{w}_{r}}$ (16)
-for all $r\in[m]$ and $k\in\mathbb{N}$, where $\eta>0$ is the learning rate. For brevity, we write $L(\bm{w})$ for $L(\bm{w},\bm{a})$.
-To simplify the notations, for the residuals of interior and boundary, we denote them by $s_{p}(\bm{w})$ and $h_{j}(\bm{w})$ respectively, i.e.,
-$s_{p}(\bm{w})=\frac{1}{\sqrt{n_{1}}}\left(\frac{\partial u}{\partial x_{0}}(\bm{x}_{p};\bm{w})-\sum\limits_{i=1}^{d}\frac{\partial^{2}u}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w})-f(\bm{x}_{p})\right)$ (17)
+Here, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+2}$ is the augmented vector from $\bm{x}$ and in the following, we write $\bm{x}$ for $\tilde{\bm{x}}$ for brevity.
+Similar to that for the $L^{2}$ regression problems, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$, output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$ and fix the output weights.
+In the framework of PINNs, given training samples $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ and $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ that are from interior and boundary respectively, we denote $s_{p}(\bm{w})$ and $h_{j}(\bm{w})$ by
+$s_{p}(\bm{w})=\frac{1}{\sqrt{n_{1}}}\left(\frac{\partial\phi}{\partial x_{0}}(\bm{x}_{p};\bm{w})-\sum\limits_{i=1}^{d}\frac{\partial^{2}\phi}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w})-f(\bm{x}_{p})\right)$ (3)
 and
-$h_{j}(\bm{w})=\frac{1}{\sqrt{n_{2}}}(u(\bm{y}_{j};\bm{w})-g(\bm{y}_{j})).$ (18)
+$h_{j}(\bm{w})=\frac{1}{\sqrt{n_{2}}}(\phi(\bm{y}_{j};\bm{w})-g(\bm{y}_{j})).$ (4)
 Then the empirical loss function can be written as
 $L(\bm{w})=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right),$ (5)
-where
-$\bm{s}(\bm{w})=(s_{1}(\bm{w}),\cdots,s_{n_{1}}(\bm{w}))^{T}\in\mathbb{R}^{n_{1}}$ (20)
-and
-$\bm{h}(\bm{w})=(h_{1}(\bm{w}),\cdots,h_{n_{2}}(\bm{w}))^{T}\in\mathbb{R}^{n_{2}}.$ (21)
-At this time, we have
-$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\sum\limits_{p=1}^{n_{1}}s_{p}(\bm{w})\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}+\sum\limits_{j=1}^{n_{2}}h_{j}(\bm{w})\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}$ (22)
-and the Gram matrix $\bm{H}(\bm{w})$ is defined as
-$\bm{H}(\bm{w})=\bm{D}^{T}\bm{D},\ \bm{D}:=\left[\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right].$ (23)
-3.2. Main Results
-First, we make the following assumptions about the training samples, which are similar to those for the $L^{2}$ regression problems.
-Assumption 3.
+where $\bm{s}(\bm{w})=(s_{1}(\bm{w}),\cdots,s_{n_{1}}(\bm{w}))^{T}\in\mathbb{R}^{n_{1}}$ and $\bm{h}(\bm{w})=(h_{1}(\bm{w}),\cdots,h_{n_{2}}(\bm{w}))^{T}\in\mathbb{R}^{n_{2}}$.
+The gradient descent updates the hidden weights by the following formulations:
+$\bm{w}_{r}(k+1)$ $=\bm{w}_{r}(k)-\eta\frac{\partial L(\bm{w}(k))}{\partial\bm{w}_{r}}$ (6)
+for all $r\in[m]$ and $k\in\mathbb{N}$, where $\eta>0$ is the learning rate. The Gram matrix $\bm{H}(\bm{w})$ is defined as $\bm{H}(\bm{w})=\bm{D}^{T}\bm{D}$, where
+$\bm{D}:=\left(\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right).$ (7)
+3. Improved Results of GD for Two-Layer PINNs
+To simplify the analysis, we make the following assumptions on the training data.
+Assumption 3.1.
 For $p\in[n_{1}]$ and $j\in[n_{2}]$, $\|\bm{x}_{p}\|_{2}\leq\sqrt{2},\|\bm{y}_{j}\|_{2}\leq\sqrt{2}$, where all inputs have been augmented.
-Assumption 4.
+Assumption 3.2.
 No two samples in $\{\bm{x}_{p}\}_{p=1}^{n_{1}}\cup\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ are parallel.
-Under Assumption 4, Lemma 3.3 in [19] implies that the Gram matrix $\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\bm{H}(\bm{w})$ is strictly positive definite and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$.
-Similarly, $\bm{H}^{\infty}$ plays an important role in the optimization process.
-The following two lemmas indicate that $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(1/\sqrt{m})$ and $\|\bm{H}(k)-\bm{H}(0)\|_{2}=\mathcal{O}(1/\sqrt{m})$, leading to the linear convergence.
-Lemma 4.
+Under Assumption 3.2, Lemma 3.3 in [21] implies that the Gram matrix $\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}[\bm{H}(\bm{w})]$ is strictly positive definite and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$.
+Similar to the case of the regression problem in [10], $\bm{H}^{\infty}$ plays an important role in the optimization process.
+Specifically, under over-parameterization and random initialization, we have two facts that (1) at initialization $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(1/\sqrt{m})$ and (2) for any iteration $k\in\mathbb{N}$, $\|\bm{H}(k)-\bm{H}(0)\|_{2}=\mathcal{O}(1/\sqrt{m})$.
+The following two lemmas can be used to verify these two facts, which are crucial in the convergence analysis.
+Lemma 3.3.
 If $m=\Omega\left(\frac{d^{4}}{\lambda_{0}^{2}}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$, we have that with probability at least $1-\delta$, $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\frac{\lambda_{0}}{4}$ and $\lambda_{min}(\bm{H}(0))\geq\frac{3}{4}\lambda_{0}$.
-Lemma 5.
+Remark 3.4.
+Under the premise of deriving the same conclusion as Lemma 3.3, Lemma 3.5 in [21] requires that $m=\tilde{\Omega}\left(\frac{(n_{1}+n_{2})^{4}}{(n_{1}n_{2})^{2}\lambda_{0}^{2}}\left(\log(\frac{1}{\delta})\right)^{7}\right)$, where some terms involving $\log(m)$ are omitted.
+In contrast, on one hand, our conclusion is independent of $n_{1}$ and $n_{2}$, and on the other hand, our conclusion exhibits a clear dependence on $d$.
+Moreover, the method in [21] involves truncating the Gaussian distribution and then applying Hoeffding’s inequality, which is quite complicated.
+In contrast, we utilize the concentration inequality for sub-Weibull random variables, which serves as a simple framework for this class of problems.
+Lemma 3.5.
 Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d.
-generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-\delta-(n_{1}+n_{2})^{2}e^{-mR}$, the following holds.
+generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-\delta-n_{1}e^{-mR}$, the following holds.
 For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
-$\|H(\bm{w})-H(0)\|_{F}<CM^{2}R,$ (24)
+$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}<CM^{2}R,$ (8)
 where $M=2(d+2)\log(2m(d+2)/\delta)$ and $C$ is a universal constant.
-In [19], the decomposition for the residual in the $(k+1)$-th iteration is same as the one in [10], i.e.,
-$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\left(\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right),$ (25)
-which leads to the requirement $\eta=\mathcal{O}(\lambda_{0})$.
-Instead, we can generalize easily the method used in the $L^{2}$ regression problems to PINNs and obtain the following recursion formula.
-Lemma 6.
-For all $k\in\mathbb{N}$, we have
-$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k),$ (26)
-where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}}(k),I_{1}^{n_{1}+1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$ and for $p\in[n_{1}]$,
-$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle,$ (27)
-for $j\in[n_{2}]$,
-$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle.$ (28)
-$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
-By using the recursion formula, we arrive at our main result.
-Theorem 2.
-Under Assumption 3 and Assumption 4, if we set the number of hidden nodes
-$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
+Remark 3.6.
+Lemma 3.6 in [21] shows that when $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}\leq R=\tilde{\mathcal{O}}\left(\frac{\lambda_{0}\delta}{(n_{1}+n_{2})(\log m)^{3}}\right)$ holds for all $r\ in[m]$, then $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{2}\leq\frac{\lambda_{0}}{4}$.
+In contrast, Lemma 3.5 only requires $R=\mathcal{O}\left(\frac{\lambda_{0}}{d^{2}(\log(m/\delta)^{2}}\right)$ to reach same result.
+For the $L^{2}$ regression problem, as shown in [10], the convergence of gradient descent requires that the learning rate $\eta=\mathcal{O}(\lambda_{0}/n^{2})$, where $n$ is the sample size of the regression problem.
+It is evident that this requirement on the learning rate is difficult to satisfy in practical scenarios, since $\lambda_{0}$ is unknown and $n^{2}$ is too large .
+For PINNs, [21] follows the methodology of [10], thus inheriting similarly stringent requirements on the learning rate.
+By investigating a new decomposition method for the residual, we arrive at our main result.
+Theorem 3.7.
+Under Assumption 3.1 and Assumption 3.2, if we set the number of hidden nodes
+$m=\Omega\left(\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
 and the learning rate $\eta=\mathcal{O}\left(\frac{1}{\|\bm{H}^{\infty}\|_{2}}\right)$, then with probability at least $1-\delta$ over the random initialization, the gradient descent algorithm satisfies
-$\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\left\|\begin{bmatrix}\bm{s}(0)\\ \bm{h}(0)\end{bmatrix}\right\|_{2}^{2}$ (29)
+$\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}$ (9)
 for all $k\in\mathbb{N}$.
-Remark 3.
-It may be confused that [19] has uesd the same method in [10], but it only requires $\eta=\mathcal{O}(\lambda_{0})$.
-Actually, it is due to that the loss function of PINN has been normalized.
+Remark 3.8.
+It may be confusing that [21] has used the same method in [10], yet it only requires $\eta=\mathcal{O}(\lambda_{0})$.
+Actually, it is because that the loss function of PINN has been normalized.
 If we let $n_{1}=n_{2}=n$ and $\widetilde{\bm{H}}^{\infty}$ be the Gram matrix induced by unnormalized loss function of PINN, then $\lambda_{min}(\bm{H}^{\infty})=\lambda_{min}(\widetilde{\bm{H}}^{\infty})/n$, leading to the convergence rate similar to that of regression problem.
-Due to the normalization of loss function, $\|\bm{H}^{\infty}\|_{2}$ is independent of the sample size $n$, but may depends on the dimension $d$.
-4. Convergence of Natural Gradient Descent for Two-Layer Physics-Informed Neural Networks
-4.1. Problem Setup
-Although we have improved the learning rates of gradient descent for $L^{2}$ regression problems and PINNs, one needs to set the learning rates to be small enough due to the agnostic $\|\bm{H}^{\infty}\|_{2}$.
-For instance, we may need to set it to be $\mathcal{O}(1/n)$ for the $L^{2}$ regression problem.
-Moreover, the convergence rate $1-\frac{\eta\lambda_{0}}{2}$ also depends on $\lambda_{0}$, which may be slow with small $\lambda_{0}$.
+At this point, due to the normalization of loss function, $\|\bm{H}^{\infty}\|_{2}$ can be bounded by the trace of $\bm{H}^{\infty}$, which is an explicit constant that is independent of the sample size $n_{1},n_{2}$.
+Therefore, in practice, we can set the learning rate to satisfy the theoretical convergence requirement, bridging the gap between theory and practice.
+Remark 3.9.
+Regarding the impact of dimensionality on the convergence of GD for PINNs, Theorem 3.7 consists of two parts: one is explicit, namely $d^{8}$, and the other is implicit, specifically $\lambda_{0}^{4}$, whose relationship with dimensionality remains unclear.
+The explicit impact arises from the form of the PDE.
+For instance, the PDE (1) we consider contains $\mathcal{O}(d)$ terms.
+In concurrent work, [26] investigated the impact of dimensionality and the order of PDEs on convergence under the setting of continuous gradient flow.
+Specifically, for PDEs of order 2, the form they considered includes $\mathcal{O}(d^{2})$ terms, and the explicit dependence on dimensionality is $d^{28}$.
+On the one hand, [26] only addressed the continuous gradient flow case, while the discrete case requires more refined analysis as stated in [10].
+On the other hand, our results can naturally extend to the PDEs they considered, where the explicit dependence on dimensionality is $d^{16}$, which is better than the result in [26].
+Investigating the lower bounds of the smallest eigenvalue of the NTK for PINNs, similar to what has been done for deep ReLU neural networks in [27], represents a promising direction for future research.
+4. Convergence of NGD for Two-Layer PINNs
+Although we have improved the learning rate of gradient descent for PINNs, it may still be necessary to set the learning rates to be sufficiently small for some complex PDEs.
+Because, although for all PDEs, $\bm{tr}(\bm{H}^{\infty})$ is an explicit constant, it depends on the form of the PDE, and for complex PDEs, it may be quite large.
+Moreover, the convergence rate $1-\frac{\eta\lambda_{0}}{2}$ also depends on $\lambda_{0}$, which depends on the sample size and may be extremely small.
 [19] and [20] have provided the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer neural networks for $L^{2}$ regression problems.
 They showed that the maximal learning rate can be $\mathcal{O}(1)$ and the convergence rate is independent of $\lambda_{0}$, which result in a faster convergence rate.
-In the section, we conduct the convergence analysis of NGD for PINNs and show that NGD indeed leads to an even faster convergence rate.
-We consider the same setup as Section 3 and aim to minimize the following empirical loss function via NGD.
-$L(\bm{w}):=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right).$ (30)
+However, their methods cannot generalize directly to PINNs.
+In the section, we conduct the convergence analysis of NGD for PINNs and demonstrate that it results in a faster convergence rate for PINNs compared to gradient descent.
+In this section, we consider the same setup as described in Section 2.
+Specifically, we focus on the PDE of the form given in (1) and follow the same initialization as described in Section 2.
+During the training process, we fix the output weight $\bm{a}$ and update the hidden weights via NGD.
+The optimization objective is the empirical loss function presented in (5), which is defined as follows:
+$L(\bm{w})=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right),$ (10)
 The NGD gives the following update rule:
-$\bm{w}(k+1)=\bm{w}(k)-\eta\bm{J}(k)^{T}\left(\bm{J}(k)\bm{J}(k)^{T}\right)^{-1}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$ (31)
-where $\bm{J}(k)=\left(\bm{J}_{1}(k)^{T},\cdots,\bm{J}_{n_{1}+n_{2}}(k)^{T}\right)^{T}\in\mathbb{R}^{(n_{1}+n_{2})\times m(d+2)}$ is the Jacobian matrix for the whole dataset and $\eta>0$ is the learning rate.
-Specifically, for $p\in[n_{1}]$,
+$\bm{w}(k+1)=\bm{w}(k)-\eta\bm{J}(k)^{T}\left(\bm{J}(k)\bm{J}(k)^{T}\right)^{-1}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix},$ (11)
+where
+$\bm{J}(k)=\left(\bm{J}_{1}(k)^{T},\cdots,\bm{J}_{n_{1}+n_{2}}(k)^{T}\right)^{T}\in\mathbb{R}^{(n_{1}+n_{2})\times m(d+2)}$
+is the Jacobian matrix for the whole dataset and $\eta>0$ is the learning rate. Specifically, for $p\in[n_{1}]$,
 $\bm{J}_{p}(k)=\left[\left(\frac{\partial s_{p}(k)}{\partial\bm{w}_{1}}\right)^{T},\cdots,\left(\frac{\partial s_{p}(k)}{\partial\bm{w}_{m}}\right)^{T}\right]\in\mathbb{R}^{1\times m(d+2)}$ (12)
@@
 $\bm{J}_{n_{1}+j}(k)=\left[\left(\frac{\partial h_{j}(k)}{\partial\bm{w}_{1}}\right)^{T},\cdots,\left(\frac{\partial h_{j}(k)}{\partial\bm{w}_{m}}\right)^{T}\right]\in\mathbb{R}^{1\times m(d+2)}.$ (13)
+Remark 4.1.
+We note that [19] and [20] have independently and concurrently established the convergence of NGD in the context of regression problems.
+The difference lies in the fact that [19] focused on ReLU activation functions, whereas [20] considered smooth activation functions and consistently set the learning rate to $1$.
+Here, following [19], we refer to this approach as NGD.
+In [20], the authors derived this method based on NTK kernel regression and termed it the Gram-Gauss-Newton (GGN) method.
+Remark 4.2.
+The classical Gauss-Newton method is given by
+$\bm{w}(k+1)=\bm{w}(k)-\left(\bm{J}(k)^{T}\bm{J}(k)\right)^{-1}\bm{J}(k)^{T}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k),\end{pmatrix}$
+which is different from our NGD method (11).
+The Gram matrix $\bm{J}(k)^{T}\bm{J}(k)$ of classical Gauss-Newton method will be extremely high-dimensional and singular for large network parameters, while our NGD $\bm{J}(k)\bm{J}(k)^{T}$ won’t.
 For the activation function of the two-layer neural network
 $\phi(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x}),$ (14)
-we consider both settings that $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function or a smooth activation function that satisfies the following assumption.
-From Lemma 3.3 in [19] and Lemma 2 in [23], we know that $\bm{H}^{\infty}$ is strictly positive definite in both settings and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$.
-Assumption 5.
-There exists a constant $c>0$ such that $|\sigma(0)|\leq c$ and for any $z,z^{{}^{\prime}}\in\mathbb{R}$,
+we consider settings where $\sigma(\cdot)$ is either the $\text{ReLU}^{3}$ activation function or a smooth activation function satisfying the following assumption.
+Assumption 4.3.
+There exists a constant $c>0$ such that $\sup_{z\in\mathbb{R}}|\sigma^{(3)}(z)|\leq c$ and for any $z,z^{{}^{\prime}}\in\mathbb{R}$,
 $|\sigma^{(k)}(z)-\sigma^{(k)}(z^{{}^{\prime}})|\leq c|z-z^{{}^{\prime}}|,$ (15)
-where $k\in\{0,1,2,3\}$. Moreover, $\sigma(\cdot)$ is analytic and is not a polynomial function.
-Unlike the approach for gradient descent, [21] and [16] focus on the change of the Jacobian matrix for NGD, rather than focusing on the Gram matrix.
-More precisely, they show that $\bm{J}(\bm{w})$ is stable with respect to $\bm{w}$, where $\bm{J}(\bm{w})$ is the Jacobian matrix with weight vector $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}$.
-However, this approach does not work for PINNs.
-Instead, we consider the stability of $\bm{J}(\bm{w})$ with respect to every single weight vector $\bm{w}_{r}$.
-Lemma 7 (Stable Jacobian).
+where $k\in\{0,1,2,3\}$.
+Moreover, $\sigma(\cdot)$ is analytic and is not a polynomial function.
+We also assume that for any positive integer $n\geq 2$, $\lim\limits_{x\to+\infty}\sigma^{(n)}(x)/\phi(x)=c_{n}\neq 0$, where the function $\phi(\cdot)$ satisfies that
+$\lim\limits_{x\to+\infty}\phi(x)=0,\lim\limits_{x\to+\infty}x\cdot\frac{\phi(bx)}{\phi(x)}=0$ (16)
+holds for any constant $b>1$.
+Lemma 4.4.
+If no two samples in $\{\bm{x}_{p}\}_{p=1}^{n_{1}}\cup\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ are parallel, then the Gram matrix $\bm{H}^{\infty}$ is strictly positive definite for activation functions that satisfy Assumption 4.2, i.e., $\lambda_{0}:=\lambda_{min}(\bm{H}^{\infty})>0$.
+Remark 4.5.
+Assumption 4.3 holds for various commonly used activation functions, including logistic function $\sigma(z)=1/(1+e^{-z})$ (with $\phi(z)=e^{-z}$), softplus function $\sigma(z)=\log(1+e^{z})$ (with $\phi(z)=e^{-z}$), hyperbolic tangent function $\sigma(z)=(e^{z}-e^{-z})/(e^{z}+e^{-z})$ (with $\phi(z)=e^{-2z}$), swish function $\sigma(z)=z/(1+e^{-z})$ (with $\phi(z)=ze^{-z}$) and others.
+Compared to the $\text{ReLU}^{3}$ activation function, these smooth activation functions are more popular for PINNs because solving PDEs typically requires high-order derivatives.
+Unlike the approach for gradient descent, [19] focus on the change of the Jacobian matrix for NGD rather than the Gram matrix.
+More precisely, they demonstrate that $\bm{J}(\bm{w})$ is stable with respect to $\bm{w}$, where $\bm{J}(\bm{w})$ is the Jacobian matrix with weight vector $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}$.
+Roughly speaking, they show that when $\|\bm{w}-\bm{w}(0)\|_{2}$ is small, then $\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}$ is also proportionately small.
+However, this approach is not applicable to PINNs, because the loss function involves derivatives.
+Roughly speaking, the stability considered in [19] is more global in nature, whereas ours is local.
+Since the subsequent conclusions require the boundedness of local weights, we do not use this stability.
+Moreover, from Theorem 1 in [19], we can see that this stability imposes additional constraints on the learning rate.
+Therefore, we instead focus on the stability of $\bm{J}(\bm{w})$ with respect to each individual weight vector $\bm{w}_{r}$, which provides a more targeted approach.
+Lemma 4.6.
 Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d.
 generated $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-P(\delta,m,R)$ the following holds.
-For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+2}$ that satsify for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
+For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+2}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
 (1) when $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, we have that
@@
 where $C$ is a universal constant, $M=2(d+2)\log(2m(d+2)/\delta)$ and
-$P(\delta,m,R)=1-\delta-n_{1}e^{-mR};$ (37)
-(2) when $\sigma(\cdot)$ satisfies Assumption 5, we have that
+$P(\delta,m,R)=\delta+n_{1}e^{-mR};$ (18)
+(2) when $\sigma(\cdot)$ satisfies Assumption 4.3, we have that
 $\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}\leq CdR$ (19)
 for $m\geq\log^{2}(1/\delta)$, where $C$ is a universal constant and $P(\delta,m,R)=\delta$.
-With the stability of Jacobian matrix, we can derive the following main result.
-Theorem 3 (Convergence of NGD).
+With the stability of Jacobian matrix, we can derive the following convergence results.
+Theorem 4.7.
 Let $L(k)=L(\bm{w}(k))$, then the following conclusions hold.
-(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, under Assumption 4, we set
-$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
+(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, under Assumption 3.2, we set
+$m=\Omega\left(\frac{1}{(1-\eta)^{2}}\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
 and $\eta\in(0,1)$, then with probability at least $1-\delta$ over the random initialization for all $k\in\mathbb{N}$
 $L(k)\leq(1-\eta)^{k}L(0).$ (20)
-(2) When $\sigma(\cdot)$ satisfies Assumption 5, under Assumption 4, we set
-$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
+(2) When $\sigma(\cdot)$ satisfies Assumption 4.3, under Assumption 3.2, we set
+$m=\Omega\left(\frac{1}{1-\eta}\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
 and $\eta\in(0,1)$, then with probability at least $1-\delta$ over the random initialization for all $k\in\mathbb{N}$
 $L(k)\leq(1-\eta)^{k}L(0).$ (21)
-In fact, NGD can enjoy a second-order convergence rate when $\sigma(\cdot)$ satisfies Assumption 5.
-Corollary 1.
-Under the setting of Theorem 3(2), in addition to (40), we also have that
-$\left\|\begin{pmatrix}\bm{s}(t+1)\\ \bm{h}(t+1)\end{pmatrix}\right\|_{2}\leq\frac{CB^{4}}{\sqrt{m\lambda_{0}^{3}}}\left\|\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}\right\|_{2}^{2},$
-where $C$ is a universal constant and $B=\sqrt{2(d+2)\log(2m(d+2)/\delta)}+1$.
+In Theorem 4.7, the requirements of $m$ with $\text{ReLU}^{3}$ and smooth activation functions exhibit different dependencies on $\lambda_{0}$ and $d$.
+The discrepancy is primarily due to the distinct formulations presented in (16) and (18) of Lemma 4.5.
+Remark 4.8.
+We first compare our results with those of NGD for $L^{2}$ regression problems.
+Given that the convergence results are the same, our focus shifts to examining the necessary conditions for the width $m$.
+As demonstrated in [19] and [20], it is required that $m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}\delta^{3}}\right)$ for ReLU activation function and $m=\Omega\left(\max\left\{\frac{n^{4}}{\lambda_{0}^{4}},\frac{n^{2}d\log(n/\delta)}{\lambda_{0}^{2}}\right\}\right)$ for smooth activation function.
+Clearly, our result has a worse dependence on $d$, which is inevitable due to the involvement of derivatives in the loss function.
+Moreover, our requirement for $m$ appears to be almost independent of $n$, primarily because our loss function has been normalized.
+With smooth activation functions, in addition to the dependence on $d$, Theorem 4.7 (2) only requires that $m=\Omega(\lambda_{0}^{-3})$.
+However, [20] demands a more stringent condition, requiring that $m=\Omega(\lambda_{0}^{-4})$.
+Continuing our analysis, we contrast our results with those of GD for PINNs.
+Roughly speaking, [21] has shown that when $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, $m=\widetilde{\Omega}\left(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}}\right)$ and $\eta=\mathcal{O}(\lambda_{0})$, then the convergence result (9) holds.
+It is evident that our result, i.e, Theorem 4.7 (1), has a milder dependence on $n_{1},n_{2}$ and $\delta$.
+Furthermore, the learning rate and convergence rate are independent of $\lambda_{0}$, resulting in faster convergence.
+Comparing with our results in Section 3, the requirement for $m$ in Theorem 4.7 (1) is the same as in Theorem 3.8, when we make $\eta$ less close to $1$.
+On the other hand, since $\eta=\mathcal{O}(1)$ and the convergence rate only depends on $\eta$, NGD can lead to faster convergence than GD.
+Note that as $\eta$ approaches $1$, the width $m$ tends to infinity, thus, the convergence results in Theorem 4.7 become vacuous.
+In fact, when $\eta=1$, NGD can enjoy a second-order convergence rate even though $m$ is finite, provided that $\sigma(\cdot)$ satisfies Assumption 4.3.
+Corollary 4.9.
+Under Assumption 3.2 and Assumption 4.3, set $\eta=1$ and
+$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right),$
+then with probability at least $1-\delta$, we have
+$\left\|\begin{pmatrix}\bm{s}(t+1)\\ \bm{h}(t+1)\end{pmatrix}\right\|_{2}\leq\frac{CB^{4}}{\sqrt{m\lambda_{0}^{3}}}\left\|\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}\right\|_{2}^{2}$
+for all $t\in\mathbb{N}$, where $C$ is a universal constant and $B=\sqrt{2(d+2)\log(2m(d+2)/\delta)}+1$.
+Remark 4.10.
+[20] has demonstrated the second-order convergence for regression problems with smooth activation functions. Specifically, it is shown in [20] that
+$\|\bm{y}-\bm{u}(t+1)\|_{2}\lesssim\frac{n^{3/2}}{\sqrt{m}\lambda_{0}^{2}}\|\bm{y}-\bm{u}(t)\|_{2}^{2}.$
+Actually, when applying our method used in Corollary 4.9, we can get a more satisfactory result as follows.
+$\|\bm{y}-\bm{u}(t+1)\|_{2}\lesssim\frac{n^{3/2}}{\sqrt{m\lambda_{0}^{3}}}\|\bm{y}-\bm{u}(t)\|_{2}^{2}.$
+Instead of inducing on the convergence rate of the empirical loss function, as shown in Condition 1, we perform induction on the movements of the hidden weights as follows.
+Condition 1.
+At the $t$-th iteration, we have $\|\bm{w}_{r}(t)\|_{2}\leq B$ and
+$\|\bm{w}_{r}(t)-\bm{w}_{r}(0)\|_{2}\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}}$
+for all $r\in[m]$, where $C$ is a universal constant and $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1.$
+With Condition 1, we can directly derive the following convergence rate of the empirical loss function.
+Corollary 4.11.
+If Condition 1 holds for $t=0,\cdots,k$ and $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, then
+$L(t)\leq(1-\eta)^{t}L(0),$
+holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 4.5 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ is in (16) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ is in (18) when $\sigma$ satisfies Assumption 4.3.
+5. Experimental Results
+Comparison to Existing Optimizers.
+We report the relative $L^{2}$-error of our NGD optimizer, the SGD optimizer, the Adam optimizer, and the L-BFGS optimizer in Table 1.
+The experiments are conducted on four problems, the 1D Poisson equation with reference silution $u_{ref}=\sin(x)$, the 2D Poisson equation with reference solution $u_{ref}=\sin(\pi x)\sin(\pi y)$, the 1D Heat equation with reference solution $u_{ref}=e^{-\frac{\pi^{2}t}{4}}\sin(\pi x)$ and the 2D Helmholtz equation with wave number $k=4$, and the reference solution $u_{ref}=\sin(\pi x)\sin(k\pi y)$.
+We implement all numerical experiments on a single NVIDIA RTX 4090 GPU and all codes are conducted by the Pytorch framework.
+The four problems, as well as the configurations used in these examples are listed in Appendix A.
+Table 1. Relative $L^{2}$-error of Different Optimizers.
+SGD Adam L-BFGS NGD
+1D Poisson 1.28e-01 ± 4.31e-02 6.46e-02 ± 1.43e-02 2.63e-04 ± 8.95e-05 1.67e-05 ± 9.07e-06
+2D Poisson 1.45e-01 ± 7.34e-02 5.32e-03 ± 9.79e-04 3.17e-03 ± 8.66e-04 1.12e-04 ± 6.99e-05
+1D Heat 5.43e-01 ± 9.98e-02 6.91e-03 ± 1.31e-03 4.98e-03 ± 1.83e-03 3.42e-04 ± 7.52e-05
+2D Helmholtz 8.48e+00 ± 6.37e+00 1.06e+00 ± 8.11e-01 3.35e+00 ± 1.94e+00 6.67e-03 ± 1.89e-03
+Learning Rate Study.
+We first examine the behavior of convergence at different learning rates to demonstrate that the training process remains stable and achieves consistent convergence.
+Table 2 presents the convergence performance of the NGD method with different learning rates on the 2D Poisson equation.
+The experimental results demonstrate that NGD maintains stable convergence across a wide range of learning rates without significant degradation in final accuracy.
+This characteristic, which markedly outperforms conventional optimization approaches, clearly illustrates the strong robustness of the NGD method to hyperparameter selection.
+Table 2. Relative $L^{2}$-error Comparison Across Different Learning Rates $\eta$.
+learning rate $\eta$ 1.0 0.5 0.1 0.05 0.01 0.005 0.001
+error 1.97e-03 1.18e-03 3.24e-04 1.87e-04 1.12e-04 1.22e-04 1.68e-04
+Network Width Study.
+In addition, a comparative analysis of the model performance is performed with progressively increasing network widths.
+The experiments are conducted on the 1D Poisson equation with $f(x)=16sin(4x)$, and the exact solution is given by $u(x)=sin(4x)$.
+The NGD optimizer is trained for 200 epochs.
+The results empirically validate the hypothesis that wider architectures improve representational capacity, leading to improved accuracy while maintaining training efficiency.
+Table 3 presents the variation of $L^{2}$-error with respect to network width.
+The results demonstrate that increasing network width leads to significant accuracy improvements.
+This trend validates that wider architectures exhibit enhanced function approximation capabilities.
+Table 3. Relative $L^{2}$-error Comparison Across Different Network Width $m$.
+$m$ 20 40 80 160 320 640 1280 2560
+error 1.59-03 7.21e-04 5.18e-04 3.8e-04 3.08e-04 2.76e-04 1.78e-04 7.05e-05
+6. Limitations
+The computational cost of NGD is mainly on the $(J\cdot J^{\top})^{-1}$ with the Jacobian matrix $J$ is of size $n\times p$, where $n=n_{1}+n_{2}$ is the training data size and $p=m(d+2)$ is the number of trainable parameters.
+So NGD will be quite expensive for large amount of training data, which may arise in high-dimensional PDEs.
+As a result, several cost-effective variants have been proposed, such as K-FAC [9] and mini-batch NGD.
+We only proved the convergence results for the full-batch NGD in this paper, and it would be interesting to investigate the convergence of these methods for PINNs in future works.
 7. Conclusion and Discussion
-In this paper, we have improved the learning rate of gradient descent for both $L^{2}$ regression problems and PINNs, indicating that gradient descent enjoys better convergence.
+In this paper, we have improved the conditions required for the convergence of gradient descent for PINNs, showing that gradient descent actually achieves a better convergence rate.
 Furthermore, we demonstrate that natural gradient descent can find the global optima of two-layer PINNs with $\text{ReLU}^{3}$ or smooth activation functions for a class of second-order linear PDEs.
-Compared to gradient descent, natural gradient descent possesses faster convergence rate and the maximal learning rate is $\mathcal{O}(1)$.
-Despite this, natural gradient descent is quite expensive in terms of cumputation and memory in training neural networks.
-Therefore, several cost-effective variants have been proposed, such as K-FAC [9] and mini-batch natural gradient descent [16].
-It would be interesting to investigate the convergence of these methods for PINNs.
-Additionally, generalizing the convergence analysis to deep neural networks is an important direction for future work.
+Compared to gradient descent, natural gradient descent exhibits a faster convergence rate and its maximal learning rate is $\mathcal{O}(1)$.
+Additionally, extending the convergence analysis to deep neural networks and studying the generalization bounds of trained PINNs are important directions for future research.
 Appendix
+8. Experimental implementation
+In this section, we provide several examples to demonstrate the superiority of the natural gradient descent(NGD) approach.
+The configurations used in these examples are listed in Table 4.
+We report the relative $L^{2}$-error of the NGD optimizer, the SGD optimizer, the Adam optimizer, and the L-BFGS optimizer in Table 1.
+The relative $L^{2}$-error is defined as follows:
+$L^{2}=\frac{||\hat{u}-u_{\mathrm{ref}}||_{2}}{||u_{\mathrm{ref}}||_{2}},$ (22)
+where $\hat{u}$ denotes the predicted solution and $u_{\mathrm{ref}}$ represents the reference solution.
+Table 4. Configurations of Different Equations.
+$N_{f}$ $N_{b}$ batch size hidden layers hidden neurons activation function
+1D Poisson 500 2 100 1 128 $\tanh(\cdot)$
+2D Poisson 1000 200 100 1 128 $\tanh(\cdot)$
+1D Heat 1000 200 100 1 128 $\tanh(\cdot)$
+2D Helmholtz 1000 200 100 1 128 $\tanh(\cdot)$
+8.1. 1D Poisson Equation
+First, we begin with a toy example of the 1D Poisson equation to display the performance of the NGD method. The equation is defined in the domain $\Omega=[0,\pi]$,
+$\left.\left\{\begin{array}[]{c}-u^{\prime\prime}(x)=f(x),\quad x\in\Omega,\\ u(x)=0,\quad x\in\partial\Omega.\end{array}\right.\right.$ (23)
+The true solution is set as $u(x)=\sin(x)$, which allows us to derive the corresponding force term $f(x)=\sin(x)$.
+We randomly sample $N_{f}=500$ points in the domain $\Omega$.
+For the neural network architecture, we employ a single hidden layer model with $128$ units and $\tanh(\cdot)$ activation functions across all computations.
+The NGD optimizer is trained for $100$ epochs, while the LBFGS optimizer is run for $1$ epoch with a maximum of $500$ iterations per epoch.
+All other optimizers are run for $10,000$ epochs for comprehensive comparison.
+The relative $L^{2}$-error is $051.67e-$ 05 for the NGD optimizer.
+Figure 1 shows the predicted solution for the 1D Poisson equation alongside the reference solution.
+The prediction is in excellent agreement with the reference solution, highlighting the superior performance of the NGD method.
+Figure 2 depicts the loss decay during the training process, we can see that the NGD method achieves a quite small loss at the very beginning.
+Figure 1. Reference Solution and Predicted Solution for the 1D Poisson Equation.
+Figure 2. Loss Decay for the 1D Poisson Equation.
+8.2. 2D Poisson Equation
+We consider a 2D Poisson equation in the domain $\Omega=[0,1]\times[0,1]$,
+$\left.\left\{\begin{array}[]{c}-\frac{\partial^{2}u}{\partial x^{2}}-\frac{\partial^{2}u}{\partial y^{2}}=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=0,\quad(x,y)\in\partial\Omega.\end{array}\right.\right.$ (24)
+The true solution is given by $u(x,y)=\sin(\pi x)\sin(\pi y)$, and the force term $f(x,y)=2\pi^{2}\sin(\pi x)\sin(\pi y)$ is consequently derived.
+We sample $N_{b}=200$ random points on the boundary $\partial\Omega$ and $N_{f}=1,000$ random points within the domain $\Omega$.
+The neural network used for computation consists of two hidden layers, each containing $128$ units with $\tanh(\cdot)$ activation functions.
+We run the NGD method for $200$ epochs, while the L-BFGS method is trained for $1$ epoch with a maximum of $5,000$ iterations per epoch.
+All other optimization methods are trained for $20,000$ epochs.
+The resulting relative $L^{2}$-error is $041.12e-$ 04.
+Figure 3 illustrates the prediction of the 2D Poisson equation, along with the exact solution and the absolute error between them.
+It is clear that the predicted solution closely matches the reference solution, further demonstrating the superior performance of the NGD method.
+Figure 4 shows the loss decay during training, demonstrating that the NGD method converges significantly faster than other optimization methods.
+Figure 3. NGD Prediction and Analysis for the 2D Poisson Equation.
+Figure 4. Loss Decay for the 2D Poisson Equation.
+8.3. 1D Heat Equation
+We consider the 1D heat equation
+$\left.\left\{\begin{array}[]{l}\frac{\partial u(t,x)}{\partial t}=\frac{1}{4}\frac{\partial^{2}u(t,x)}{\partial x^{2}},\quad(t,x)\in\left[0,1\right]^{2},\\ u(0,x)=\sin(\pi x),\quad x\in\left[0,1\right],\\ u(t,x)=0,\quad(t,x)\in\left[0,1\right]\times\{0,1\}.\end{array}\right.\right.$ (25)
+The reference solution is analytically defined by $u(t,x)=\exp(-\frac{\pi^{2}t}{4})\sin(\pi x)$.
+We generate $N_{b}=200$ random sampling points for the boundary and initial conditions and $N_{f}=1,000$ random points in the domain $\Omega$ to evaluate the PDE residual.
+The neural network used for all computations consists of $1$ hidden layer, each containing $128$ neurons with $\tanh(\cdot)$ activation functions.
+To train the model, we run the NGD method for $200$ epochs and the L-BFGS method for $1$ epoch with a maximum of $17$ iterations per epoch, and other optimizers are trained for $10,000$ epochs.
+The resulting relative $L^{2}$-error is $043.42e-$ 04.
+Figure 5 provides a visual comparison between the predicted and exact solutions for the 1D heat equation, along with the corresponding absolute error distribution.
+The high degree of accuracy in the predicted solution demonstrates the effectiveness of the NGD method, showing its ability to capture the solution with remarkable precision.
+Figure 6 shows the loss curve over the course of training for the 1D heat equation.
+Notably, the NGD method rapidly reduces the loss, reaching a low value in the training process, demonstrating its efficiency in optimization.
+Figure 5. NGD Prediction and Analysis for the 1D Heat Equation.
+Figure 6. Loss Decay for the 1D Heat Equation.
+8.4. 2D Helmholtz Equation
+We deal with the 2D helmholtz equation on the domain $\Omega=[0,1]\times[0,1]$ given by
+$\left\{\begin{array}[]{l}\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}+k^{2}u(x,y)=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=0,\quad(x,y)\in\partial\Omega.\end{array}\right.$ (26)
+The reference solution for $k=4$ is $u(x,y)=\sin(\pi x)\sin(4\pi y)$, and the force term $f(x,y)$ can be easily computed.
+To evaluate the performance of the NGD approach on the 2D Helmholtz equation, we generate $N_{b}=200$ random boundary points on $\partial\Omega$ and $N_{f}=1,000$ random points inside the domain $\Omega$.
+The neural network employed consists of $1$ hidden layers with $128$ neurons per layer, utilizing $\tanh(\cdot)$ activation functions.
+Training is carried out for $200$ epochs using the NGD method and $1$ epoch with a maximum of $15$ iterations per epoch.
+All other optimizers are run for $20,000$ epochs for comparison.
+The computed relative $L^{2}$-error is $036.67e-$ 03, which is $3$ orders of magnitude lower than those of the remaining optimizers.
+Figure 7 illustrates the predicted solution along with the exact reference solution and the absolute error distribution.
+The results indicate that the NGD method effectively captures the oscillatory nature of the Helmholtz equation, achieving a high level of accuracy.
+Figure 8 shows the evolution of the loss function during training for the 2D Helmholtz equation.
+In particular, the NGD method demonstrates rapid convergence, achieving a low loss value at the end of the training process.
+Figure 7. NGD Prediction and Analysis for the 2D Helmholtz Equation.
+Figure 8. Loss Decay for the 2D Helmholtz Equation.
 9. Proof of Section 3
@@
 for all $i\in[n]$.
-Note that the event happens if and only if $|\bm{w}_{r}(0)^{T}\bm{x}_{i}|<R$, thus by the anti-concentration inequality of Gaussian distribution, we have
-$P(A_{ir})=P_{z\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2}^{2})}(|z|<R)=P_{z\sim\mathcal{N}(0,1)}\left(|z|<\frac{R}{\|\bm{x}_{i}\|_{2}}\right)\leq\frac{2R}{\sqrt{2\pi}\|\bm{x}_{i}\|_{2}}\leq\frac{2R}{\sqrt{2\pi}},$ (42)
-as $\|\bm{x}_{i}\|_{2}\geq 1$ for all $i\in[n]$.
-6.1. Proof of Lemma 2
-Proof.
-Recall that
-$\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2}$ (43)
-$=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}(\bm{x}_{i}^{T}\bm{x}_{j})^{2}\left(\frac{1}{m}\sum\limits_{r=1}^{m}(I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\})\right)^{2}.$
-To prove the measurability, we will show that $I\left\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\right\}$ can be approximated by the sequence $\{I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0,\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\}_{k\in\mathbb{N}}$, where $\tilde{\bm{w}}_{k}$ is a rational vector in $\mathbb{Q}^{d+1}$ for each $k\in\mathbb{N}$, with $\mathbb{Q}$ denoting the set of all rational numbers.
-Therefore, taking the supremum of $\bm{w}$ within the ball is equivalent to taking the supremum over all rational vectors within this ball, which implies that the random variable is measurable.
-We first focus on $I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}$ and let $\bm{w}_{r}=(\bm{w}_{r1}^{T},w_{r0})^{T},\bm{x}_{i}=(\bm{x}_{i1}^{T},1)^{T}$ with $w_{r0}\in\mathbb{R}$ and $\bm{w}_{r1},\bm{x}_{i1}\in\mathbb{R}^{d}$.
-For each $k\in\mathbb{N}$, we can take $\bm{\beta}_{k}\in\mathbb{Q}^{d}$ such that $\|\bm{\beta}_{k}-\bm{w}_{r1}\|_{2}\leq\frac{1}{2k}$.
-Then choose $t_{k}\in\mathbb{Q}$ such that $t\in(w_{r0}+\frac{1}{2k},w_{r0}+\frac{1}{k}]$ and let $\tilde{\bm{w}}_{k}=(\bm{\beta}_{k}^{T},t_{k})^{T}$.
-From the construction, we have
-$I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0\}=I\{\bm{\beta}_{k}^{T}\bm{x}_{i1}+t_{k}\geq 0\}=I\{\bm{w}_{r}^{T}\bm{x}_{i}+t_{k}-w_{r0}+(\bm{\beta}_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}\geq 0\},$
-as $\bm{w}_{r}^{T}\bm{x}_{i}=\bm{w}_{r1}^{T}\bm{x}_{i1}+w_{r0}$.
-Define $r_{k}=t_{k}-w_{r0}+(\beta_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}$, then $r_{k}>0$ and $r_{k}\to 0$, as $|(\beta_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}|\leq\|\beta_{k}-\bm{w}_{r1}\|_{2}\leq\frac{1}{2k}$ and $t-w_{r0}\in(\frac{1}{2k},\frac{1}{k}]$.
-Since the function $u\rightarrow I\{u\geq x\}$ is right-continuous for any $x\in\mathbb{R}$.
-Thus $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0\}=I\{\bm{w}_{k}^{T}\bm{x}_{i}+r_{k}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{i}\geq 0\}$ as $k\to\infty$.
-Meanwhile, when $k$ is large enough, $\tilde{\bm{w}}_{k}$ is in the small ball centered at $\bm{w}_{r}(0)$, i.e., $\|\tilde{\bm{w}}_{k}-\bm{w}_{r}(0)\|_{2}<R$.
-Similarly, we have that $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{j}\geq 0\}$ as $k\to\infty$.
-Thus, we can deduce that $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0,\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{i}\geq 0,\bm{w}_{k}^{T}\bm{x}_{j}\geq 0\}$ as $k\to\infty$.
-Therefore,
-$\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2}=\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\bm{w}_{1}\in\mathbb{Q}^{d+1}\\ \cdots\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R,\bm{w}_{m}\in\mathbb{Q}^{d+1}\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2},$
-which leads to the measurability.
-From the definition of $A_{ir}$, i.e., (41), and the fact $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, we can deduce that
-$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\}|\leq I\{A_{ir}\}.$
-Thus, when both $A_{ir}$ and $A_{jr}$ do not happen, we have
-$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|$
-$\leq|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\geq 0\}|+|I\{\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\geq 0\}|$
-$\leq I\{A_{ir}\}+I\{A_{jr}\}=0,$
-which imples
-$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|\leq I\{A_{ir}\lor A_{jr}\}.$ (44)
-Therefore, combining (43) and (44) yields that
-$\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots,\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(w)_{ij}-H(0)_{ij}|^{2}$ (45)
-$\leq 4\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots,\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\left(\frac{1}{m}\sum\limits_{r=1}^{m}(I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\})\right)^{2}$
-$\leq 4\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\left(\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\right)^{2}.$
-It remains only to bound the term $\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}$. Note that
-$\mathbb{E}[I\{A_{ir}\lor A_{jr}\}]\leq P(A_{ir})+P(A_{jr})\leq\frac{4R}{\sqrt{2\pi}}$
-and then
-$Var(I\{A_{ir}\lor A_{jr}\})\leq\mathbb{E}[I\{A_{ir}\lor A_{jr}\}\leq\frac{4R}{\sqrt{2\pi}}.$
-Thus, applying Bernstein’s inequality for the random variable $I\{A_{ir}\lor A_{jr}\}-\mathbb{E}[I\{A_{ir}\lor A_{jr}\}]$ yields that with probability at least $1-e^{-t}$,
-$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\leq\frac{4R}{\sqrt{2\pi}}+\sqrt{\frac{2t}{m}\frac{4R}{\sqrt{2\pi}}}+\frac{t}{3m}.$
-Choosing $t=mR$ and taking a union bound for $i,j\in[n]$ yields that with probability at least $1-n^{2}e^{-mR}$,
-$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\leq 4R$ (46)
-holds for all $i,j\in[n]$.
-Plugging (46) into (45) leads to the conclusion. ∎
-6.2. Proof of Lemma 3
-Proof.
-First, we can decompose $u_{i}(k+1)-u_{i}(k)$ as follows.
-$u_{i}(k+1)-u_{i}(k)$ $=u_{i}(k+1)-u_{i}(k)-\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle+\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$
-$:=I_{1}^{i}(k)+I_{2}^{i}(k).$
-For $I_{2}^{i}(k)$, we have
-$I_{2}^{i}(k)$ $=\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$
-$=\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle$
-$=\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\rangle$
-$=\sum\limits_{r=1}^{m}(-\eta)\sum\limits_{j=1}^{n}(u_{j}-y_{j})\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\frac{\partial u_{j}(k)}{\partial\bm{w}_{r}}\rangle$
-$=\eta\sum\limits_{j=1}^{n}(\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\frac{\partial u_{j}(k)}{\partial\bm{w}_{r}}\rangle)(y_{j}-u_{j})$
-$=\eta[\bm{H}(k)]_{i}(y-u(k)),$
-where $[\bm{H}(k)]_{i}$ denotes the $i$-row of the matrix $\bm{H}(k)$.
-Thus,
-$u_{i}(k+1)-u_{i}(k)=\eta[\bm{H}(k)]_{i}(y-u(k))+I_{1}^{i}(k)$
-and then
-$(y_{i}-u_{i}(k))-(y_{i}-u_{i}(k+1))=\eta[\bm{H}(k)]_{i}(y-u(k))+I_{1}^{i}(k),$
-which yields
-$(\bm{y}-\bm{u}(k))-(\bm{y}-\bm{u}(k+1))=\eta\bm{H}(k)(\bm{y}-\bm{u}(k))+I_{1}(k).$
-A simple transformation leads to the conclusion
-$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k).$
-6.3. Proof of Theorem 1
-Proof.
-The proof follows a similar induction to that in [10]. Our induction hypothesis is the following condition.
-Condition 1.
-At the $t$-th iteration, we have
-$\|\bm{y}-\bm{u}(t)\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{t}\|\bm{y}-\bm{u}(0)\|_{2}^{2}.$ (47)
-Suppose that Condition 1 holds for $t=0,\cdots,k$, then Corollary 4.1 in [10] implies that for every $r\in[m]$ and $t=0,\cdots,k$
-$\|\bm{w}_{r}(t+1)-\bm{w}_{r}(0)\|_{2}\leq\frac{4\sqrt{n}\|\bm{y}-\bm{u}(0)\|_{2}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}}.$ (48)
-In the following, we aim to show that Condition 1 also holds for $t=k+1$, thus the conclusion of Theorem 1 holds directly.
-Let $S_{i}=\{r\in[m]:I\{A_{ir}\}=0\}$ and $S_{i}^{\perp}=[m]\textbackslash S_{i}$.
-Recall that
-$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k).$
-As Condition 1 holds for $t=0,\cdots,k$, from (48), we have that for any $r\in[m]$,
-$\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}.$
-By taking $R=\frac{\lambda_{0}}{32n}$ in Lemma 2 and $R^{{}^{\prime}}\leq R$, we have that
-$\|\bm{H}(k)-\bm{H}(0)\|_{2}\leq\frac{\lambda_{0}}{4},$
-which implies that $\lambda_{min}(\bm{H}(k))\geq\lambda_{min}(\bm{H}(0))-\frac{\lambda_{0}}{4}\geq\frac{\lambda_{0}}{2}$ and
-$\|\bm{H}(k)\|_{2}\leq\|\bm{H}(0)\|_{2}+\frac{\lambda_{0}}{4}\leq\|\bm{H}^{\infty}\|_{2}+\frac{\lambda_{0}}{2}\leq\frac{3}{2}\|\bm{H}^{\infty}\|_{2}.$
-Thus, when $\eta\leq\frac{2}{3\|H^{\infty}\|_{2}}$, $\bm{I}-\eta\bm{H}(k)$ is positive definite and meanwhile $\|\bm{I}-\eta\bm{H}(k)\|_{2}\leq 1-\frac{\eta\lambda_{0}}{2}$.
-Combining with the recursion formula (10) yields that
-$\|\bm{y}-\bm{u}(k+1)\|_{2}^{2}$ $=\|(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k)\|_{2}^{2}$ (49)
-$=\|(I-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}-2\langle(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k)),\bm{I}_{1}(k)\rangle$
-$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\|\bm{I}_{1}(k)\|_{2}\|(I-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))\|_{2}$
-$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\|\bm{I}_{1}(k)\|_{2}\|\bm{y}-\bm{u}(k)\|_{2}.$
-Then, it remains only to bound the term $\|\bm{I}_{1}(k)\|_{2}$.
-Recall that for $i\in[n]$,
-$I_{1}^{i}(k)=\sum\limits_{r=1}^{m}\frac{a_{r}}{\sqrt{m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\right].$
-Note that for $r\in S_{i}$, we can deduce that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{i}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$, as $\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$ and $\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$.
-Thus, for $r\in S_{i}$, we have
-$\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
-$=[(\bm{w}_{r}(k+1)^{T}\bm{x}i)I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{i}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}]$
-$\quad-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
-$=[(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}]$
-$\quad-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
-$=0.$
-It implies that
-$I_{1}^{i}(k)=\sum\limits_{r\in S_{i}^{\perp}}\frac{a_{r}}{\sqrt{m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\right].$
-Therefore,
-$|I_{1}^{i}(k)|$ $\leq\frac{1}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}2\sqrt{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}$ (50)
-$=\frac{2\sqrt{2}}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}\|-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\|_{2}$
-$\leq\frac{4}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}\frac{\eta\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{\sqrt{m}}$
-$=\frac{4\eta\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\},$
-where the second inequality follows from that
-$\frac{\partial L(k)}{\partial\bm{w}_{r}}$ $=\sum\limits_{i=1}^{n}(u_{i}(k)-y_{i})\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}}$
-$=\sum\limits_{i=1}^{n}(u_{i}(k)-y_{i})\frac{a_{r}}{\sqrt{m}}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\bm{x}_{i}$
-and then
-$\left\|\frac{\partial L(k)}{\partial\bm{w}_{r}}\right\|_{2}\leq\frac{\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{\sqrt{m}},$
-as $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$.
-Since $S_{i}=\{r\in[m]:I\{A_{ir}\}=0\}$, it follows that $I\{r\in S_{i}^{\perp}\}=I\{A_{ir}\}$, then applying Berstein’s inequality yields that with probability at least $1-ne^{-mR}$,
-$\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\}\leq 4R,$ (51)
-holds for all $i\in[n]$.
-Thus from (50) and (51), we have
-$\|I_{1}(k)\|_{2}$ $=\sqrt{\sum\limits_{i=1}^{n}|I_{1}^{i}(k)|^{2}}$ (52)
-$\leq 4\eta\sqrt{n}\|y-u(k)\|_{2}\sqrt{\sum\limits_{i=1}^{n}(\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\})^{2}}$
-$\leq 8\eta nR\|y-u(k)\|_{2}$
-$=\frac{\eta\lambda_{0}}{8}\|y-u(k)\|_{2},$
-where the last inequality is due to $R=\frac{\lambda_{0}}{64n}$.
-Plugging (52) into (49) yields that
-$\|y-u(k+1)\|_{2}^{2}$ $\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\|\bm{I}_{1}(k)\|_{2}\|\bm{y}-\bm{u}(k)\|_{2}$
-$\leq\left[(1-\frac{\eta\lambda_{0}}{2})^{2}+\frac{\eta^{2}\lambda_{0}^{2}}{64}+2(1-\frac{\eta\lambda_{0}}{2})\frac{\eta\lambda_{0}}{8}\right]\|\bm{y}-\bm{u}(k)\|_{2}^{2}$
-$\leq\left[1-\eta\lambda_{0}+\frac{\eta^{2}\lambda_{0}^{2}}{4}+\frac{\eta^{2}\lambda_{0}^{2}}{64}+\frac{\eta\lambda_{0}}{4}\right]\|\bm{y}-\bm{u}(k)\|_{2}^{2}$
-$\leq(1-\frac{\eta\lambda_{0}}{2})\|\bm{y}-\bm{u}(k)\|_{2}^{2},$
-where the last inequality follows from the fact that
-$\eta\leq\frac{2}{3\|\bm{H}^{\infty}\|_{2}}\leq\frac{2}{3\lambda_{0}},$
-which imples that
-$\frac{\eta^{2}\lambda_{0}^{2}}{4}+\frac{\eta^{2}\lambda_{0}^{2}}{64}\leq\frac{\eta\lambda_{0}}{4}.$
-Finally, we need consider the requirement for $m$.
-First, Lemma 1 shows that $m=\Omega\left(\frac{n^{2}}{\lambda_{0}^{2}}\log\left(\frac{n}{\delta}\right)\right)$.
-Second, from $R^{{}^{\prime}}\leq R=\frac{\lambda_{0}}{32n}$ and $R^{{}^{\prime}}=\frac{4\sqrt{n}\|\bm{y}-\bm{u}(0)\|_{2}}{\sqrt{m}\lambda_{0}}$, we know
-$m=\Omega\left(\frac{n^{3}\|\bm{y}-\bm{u}(0)\|_{2}^{2}}{\lambda_{0}^{4}}\right).$
-Then from the estimation in Lemma 10 for the initial prediction $\|\bm{y}-\bm{u}(0)\|_{2}$, we obtain that
-$m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}}\log\left(\frac{n}{\delta}\right)\right).$
-7. Proof of Section 3
-Before the proofs, we first recall that
-$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{1}}}\left(\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}\right)$ (53)
+Note that the event happens if and only if $|\bm{w}_{r}(0)^{T}\bm{x}_{i}|<\|\bm{x}_{i}\|_{2}R$, thus by the anti-concentration inequality of Gaussian distribution, we have
+$P(A_{ir})=P_{z\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2}^{2})}(|z|<R)=P_{z\sim\mathcal{N}(0,1)}\left(|z|<R\right)\leq\frac{2R}{\sqrt{2\pi}}.$ (28)
+Let $S_{i}=\{r\in[m]:I\{A_{ir}\}=0\}$ and $S_{i}^{\perp}=[m]\backslash S_{i}$.
+Then, we need to recall that
+$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{1}}}\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}\right]$ (29)
 and
 $\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{2}}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j}.$ (30)
-7.1. Proof of Lemma 4
+9.1. Proof of Lemma 3.3
 Proof.
@@
 Note that the entries of $\bm{H}(0)-\bm{H}^{\infty}$ have three forms as follows.
-$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle,$ (55)
-$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle$ (56)
+$\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0)}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right],$ (31)
+$\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]$ (32)
 and
-$\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle.$ (57)
+$\sum\limits_{r=1}^{m}\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right].$ (33)
 For the first form (30), to simplify the analysis, we let
@@
 then
-$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle=\frac{1}{n_{1}m}\sum\limits_{r=1}^{m}(X_{r}(ij)-\mathbb{E}X_{r}(ij)).$
+$\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{p}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{p}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]=\frac{1}{n_{1}m}\sum\limits_{r=1}^{m}\left[X_{r}(ij)-\mathbb{E}X_{r}(ij)\right].$
 Note that $|X_{r}(ij)|\lesssim 1+\|\bm{w}_{r}(0)\|_{2}^{4}$, thus
-$\|X_{r}(ij)\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{4}\right\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{2}\right\|_{\psi_{1}}^{2}\lesssim d^{2}.$
+$\left\|X_{r}(ij)\right\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{4}\right\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{2}\right\|_{\psi_{1}}^{2}\lesssim d^{2}.$
+Here, for more details on the Orlicz norm, see the remarks after Lemma D.1.
 For the centered random variable, the property of $\psi_{\frac{1}{2}}$ quasi-norm implies that
 $\|X_{r}(ij)-\mathbb{E}[X_{r}(ij)]\|_{\psi_{\frac{1}{2}}}\lesssim\|X_{r}(ij)\|_{\psi_{\frac{1}{2}}}+\|\mathbb{E}[X_{r}(ij)]\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}.$
-Therefore, applying Lemma 8 yields that with probability at least $1-\delta$,
-$|\sum\limits_{r=1}^{m}\frac{1}{m}(X_{r}(ij)-\mathbb{E}[X_{r}(ij)])|\lesssim\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2},$
-which means
-$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}w_{r}}\rangle|\lesssim\frac{d^{2}}{n_{1}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{1}m}(\log(\frac{1}{\delta}))^{2}.$ (58)
+Therefore, applying Lemma D.1 yields that with probability at least $1-\delta$,
+$\left|\sum\limits_{r=1}^{m}\frac{1}{m}\left[X_{r}(ij)-\mathbb{E}X_{r}(ij)\right]\right|\lesssim\frac{d^{2}}{\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{m}\left(\log\left(\frac{1}{\delta}\right)\right)^{2},$
+which directly yields that
+$\left|\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{p}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{p}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right|\lesssim\frac{d^{2}}{n_{1}\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{n_{1}m}\left(\log\left(\frac{1}{\delta}\right)\right)^{2}.$ (34)
 Similarly, for the second form (31) and third form (32), we can deduce that
-$\left\|\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}$
+$\left\|\left\langle\frac{\partial s_{i}(\bm{w}(0)}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}$
 and
-$\left\|\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{n_{2}m}.$
-Thus applying Lemma 8 yields that with probability at least $1-\delta$,
-$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}\log(\frac{1}{\delta})$ (59)
+$\left\|\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{n_{2}m}.$
+Thus applying Lemma D.1 yields that with probability at least $1-\delta$,
+$\left|\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right|\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}\log\left(\frac{1}{\delta}\right)$ (35)
 and with probability at least $1-\delta$,
-$|\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{n_{2}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{2}m}\log(\frac{1}{\delta}).$ (60)
+$\left|\sum\limits_{r=1}^{m}\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right|\lesssim\frac{d^{2}}{n_{2}\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{n_{2}m}\log\left(\frac{1}{\delta}\right).$ (36)
 Combining (33), (34) and (35), we can deduce that with probability at least $1-\delta$,
@@
 $\leq\|\bm{H}(0)-\bm{H}^{\infty}\|_{F}^{2}$
-$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})+\frac{d^{4}}{m^{2}}(\log(\frac{n_{1}+n_{2}}{\delta}))^{4}$
-$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta}).$
-Thus when $\sqrt{\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})}\lesssim\frac{\lambda_{0}}{4}$, i.e.,
-$m=\Omega(\frac{d^{4}}{\lambda_{0}^{2}}\log(\frac{n_{1}+n_{2}}{\delta})),$
+$\lesssim\frac{d^{4}}{m}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{d^{4}}{m^{2}}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)^{4}$
+$\lesssim\frac{d^{4}}{m}\log\left(\frac{n_{1}+n_{2}}{\delta}\right).$
+Thus when $\sqrt{\frac{d^{4}}{m}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)}\lesssim\frac{\lambda_{0}}{4}$, i.e.,
+$m=\Omega\left(\frac{d^{4}}{\lambda_{0}^{2}}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right),$
 we have $\lambda_{min}(\bm{H}(0))\geq\frac{3}{4}\lambda_{0}$.
-7.2. Proof of Lemma 5
+9.2. Proof of Lemma 3.5
 Proof.
 We first reformulate the term $\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ in (28) as follows.
-$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}w_{r0}(k)x_{p0}\\ w_{r0}(k)\bm{x}_{p1}-2\bm{w}_{r1}(k)\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}.$
-Similar to Lemma 4, it suffices to bound $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}$, which can in turn allows us to bound each entry of $\bm{H}(\bm{w})-\bm{H}(0)$.
-For $i\in[n_{1}]$ and $j\in[n_{1}]$, we have
-$H_{ij}(\bm{w})$ $=\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}}{\partial\bm{w}_{r}},\frac{\partial s_{j}}{\partial\bm{w}_{r}}\rangle$
+$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{1}}}\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}w_{r0}x_{p0}\\ w_{r0}\bm{x}_{p1}-2\bm{w}_{r1}\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}\right].$
+It suffices to bound $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}$, which can in turn allows us to bound each entry of $\bm{H}(\bm{w})-\bm{H}(0)$.
+For $i\in[n_{1}]$ and $j\in[n_{1}]$, we have that
+$H_{ij}(\bm{w})$ $=\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\right\rangle$
 $=\frac{1}{n_{1}m}\sum\limits_{r=1}^{m}\left\langle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}w_{r0}x_{i0}\\ w_{r0}\bm{x}_{i1}-2\bm{w}_{r1}\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{i},\right.$
@@
 In fact, when bounding the corresponding terms for $H_{ij}(\bm{w})-H_{ij}(0)$, the first four classes can be grouped into one category.
-They are of the form $f_{1}(\bm{w})f_{2}(\bm{w})f_{3}(\bm{w})f_{4}(\bm{w})$, where for each $1\leq i\leq 4$, $f_{i}(\bm{w})$ is Lipschitz continuous with respect to $\|\cdot\|_{2}$ and $|f_{i}(\bm{w})|\lesssim\|\bm{w}\|_{2}$ (Note that $\sigma^{{}^{\prime}}(\cdot)=(\sigma^{{}^{\prime\prime}}(\cdot))^{2}$).
+They are of the form $f_{1}(\bm{w})f_{2}(\bm{w})f_{3}(\bm{w})f_{4}(\bm{w})$, where for each $i$ ($1\leq i\leq 4$), $f_{i}(\bm{w})$ is Lipschitz continuous with respect to $\|\cdot\|_{2}$ and $|f_{i}(\bm{w})|\lesssim\|\bm{w}\|_{2}$ (Note that $\sigma^{{}^{\prime}}(\cdot)=(\sigma^{{}^{\prime\prime}}(\cdot))^{2}$).
 On the other hand, when $\|\bm{w}_{1}-\bm{w}_{2}\|_{2}\leq R\leq 1$, we can deduce that
 $|f_{1}(\bm{w}_{1})f_{2}(\bm{w}_{1})f_{3}(\bm{w}_{1})f_{4}(\bm{w}_{1})-f_{1}(\bm{w}_{2})f_{2}(\bm{w}_{2})f_{3}(\bm{w}_{2})f_{4}(\bm{w}_{2})|\lesssim R(\|\bm{w}_{1}\|_{2}^{3}+1).$
-Thus, for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the first four classes, we can deduce that they are less than $CR(\|\bm{w}_{r}(0)\|_{2}^{3}+1)$, where $C$ is a universal constant.
+Thus, for the terms in $H_{ij}(\bm{w})-H_{ij}(0)$ that belong to the first four classes, we can deduce that they are less than $CR(\|\bm{w}_{r}(0)\|_{2}^{3}+1)$, where $C$ is a universal constant.
 For the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are both involving $\sigma^{{}^{\prime\prime\prime}}$ that is not Lipschitz continuous.
@@
 $|\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{j})|=|I\{\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|\leq I\{A_{jr}\},$
-where the definition of $A_{jr}$ is same as (44) for the regression problems.
-Thus, we can deduce that for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are less than
-$C\left[(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{3}+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1)\right],$
+where the event $A_{jr}$ has been defined in (36).
+Thus, we can deduce that for the terms in $H_{ij}(\bm{w})-H_{ij}(0)$ that belong to the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are less than
+$C\left[(I\{A_{ir}\}+I\{A_{jr}\})(\|\bm{w}_{r}(0)\|_{2}^{3}+1)+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1)\right],$
 where $C$ is a universal constant.
@@
 $|\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{j})\|\bm{w}_{r1}(0)\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j}|$
-$\lesssim I\{A_{ir}\lor A_{jr}\}\|\bm{w}_{r}(0)\|_{2}^{4}+R\|\bm{w}_{r}(0)\|_{2}^{3}+R.$
-Combining the upper bounds for the six classes, we have that
-$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{3})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right]$ (61)
-$\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right],$
+$\lesssim I\{A_{ir}\lor A_{jr}\}\|\bm{w}_{r}(0)\|_{2}^{4}+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1).$
+Combining the upper bounds for the terms in the six classes, we have that
+$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}\left(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{3}\right)+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})(\|\bm{w}_{r}(0)\|_{2}^{4}+\|\bm{w}_{r}(0)\|_{2}^{3}+1)+R\right]$ (37)
+$\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}\left(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}\right)+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})(\|\bm{w}_{r}(0)\|_{2}^{4}+1)+R\right],$
 where the last inequality follows from that $\|\bm{w}_{r}(0)\|_{2}^{3}\lesssim\|\bm{w}_{r}(0)\|_{2}^{4}+1$ due to Young’s inequality for products.
@@
 Since
-$P(|w_{ri}(0)|^{2}\geq 2\log(\frac{2}{\delta}))\leq\delta$
+$P\left(|w_{ri}(0)|^{2}\geq 2\log\left(\frac{2}{\delta}\right)\right)\leq\delta$
 and then
-$P(\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log(\frac{2(d+2)}{\delta}))\leq\delta,$
-we obtain that
-$P(\exists r\in[m],\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\frac{2m(d+2)}{\delta})\leq\delta.$ (62)
-Let $M=2(d+2)\log(\frac{2m(d+2)}{\delta})$, then
-$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}$
-$=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}\leq M\}+\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}$
-$\leq\frac{M^{2}}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}.$
+$P\left(\|\bm{w}_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\left(\frac{2(d+2)}{\delta}\right)\right)\leq\delta.$
+This implies that
+$P\left(\exists r\in[m],\|\bm{w}_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)\right)\leq\delta.$ (38)
+Let $M=2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)$, then
+$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}$
+$=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}$
+$\leq\frac{M^{2}}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}.$
 Applying Bernstein’s inequality for the first term yields that with probability at least $1-e^{-mR}$,
 $\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\leq 4R.$
-Moreover, with probability at least $1-\delta$, $I\{\|w_{r}(0)\|_{2}^{2}>M\}=0$ holds for all $r\in[m]$.
+Moreover, from (37), we have that with probability at least $1-\delta$, $I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}=0$ holds for all $r\in[m]$.
 Thus from (36), with probability at least $1-\delta-n_{1}e^{-mR}$, we have that for any $i\in[n_{1}]$ and $j\in[n_{1}]$,
@@
 $\lesssim\frac{1}{n_{1}}M^{2}R.$
-For $i\in[n_{1}],j\in[n_{1}+2,n_{2}]$ and $i\in[n_{1}+1,n_{2}],j\in[n_{2}]$, from the form of $\frac{\partial h_{j}}{\partial\bm{w}_{r}}$, i.e.,
-$\frac{\partial h_{j}}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j},$
-we can obtain similar results for the terms $\langle\frac{\partial s_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$ and $\langle\frac{\partial h_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$.
-With all results above, we have that with probability at least $1-\delta-(n_{1}+n_{2})e^{-mR}$,
+For $i\in[n_{1}],j\in[n_{1}+2,n_{2}]$ and $i\in[n_{1}+1,n_{2}],j\in[n_{2}]$, from the form of $\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}$, i.e.,
+$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j},$
+we can obtain similar results for the terms $\left\langle\frac{\partial s_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\right\rangle$ and $\left\langle\frac{\partial h_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\right\rangle$.
+With all results above, we have that with probability at least $1-\delta-n_{1}e^{-mR}$,
 $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}\lesssim M^{2}R.$
-7.3. Proof of Lemma 6
+9.3. Proof of Lemma B.1
+Indeed, the stringent requirement of the learning rate in [10] stems from an inadequate decomposition method for the residual.
+Specifically, in [21], the decomposition for the residual in the $(k+1)$-th iteration is same as the one in [10], i.e.,
+$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\left[\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}-\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right],$ (39)
+which leads to the requirements that $\eta=\mathcal{O}(\lambda_{0})$ and $m=Poly(n_{1},n_{2},1/\delta)$.
+Thus, it requires a new approach to achieve the improvements for $\eta$ and $m$.
+In fact, we can derive the following recursion formula.
+Lemma 9.1.
+For all $k\in\mathbb{N}$, we have
+$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k),$ (40)
+where
+$\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$
+and for $p\in[n_{1}]$,
+$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle,$ (41)
+for $j\in[n_{2}]$,
+$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle.$ (42)
+In the recursion formula (39), $\bm{I}_{1}(k)$ serves as a residual term.
+From the proof, we can see that $\|\bm{I}_{1}(k)\|_{2}=\mathcal{O}(1/\sqrt{m})$ and thus, as $m$ becomes large enough, only the term $\bm{I}-\eta\bm{H}(k)$ is significant.
+This observation is the reason for the requirement of $\eta$.
 Proof.
-Similar to the proof of Lemma 3, we have
-$s_{p}(k+1)-s_{p}(k)$ $=\left[s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (63)
+First, we have
+$s_{p}(k+1)-s_{p}(k)$ $=\left[s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle\right]+\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle$ (43)
 $:=I_{1}^{p}(k)+I_{2}^{p}(k).$
-For the second term $I_{2}^{p}(k)$, we have that
-$I_{2}^{p}(k)$ $=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (64)
-$=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},-\eta\frac{\partial L(k)}{\partial\bm{w}}\rangle$
-$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial L(k)}{\partial\bm{w}_{r}}\rangle$
-$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\sum\limits_{t=1}^{n_{1}}s_{t}(k)\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}+\lambda\sum\limits_{j=1}^{n_{2}}h_{j}(k)\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\rangle$
-$=-\eta[\bm{H}(k)]_{p}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$
+For the second term $I_{2}^{p}(k)$, from the updating rule of gradient descent, we have that
+$I_{2}^{p}(k)$ $=\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle$ (44)
+$=\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},-\eta\frac{\partial L(k)}{\partial\bm{w}}\right\rangle$
+$=-\sum\limits_{r=1}^{m}\eta\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial L(k)}{\partial\bm{w}_{r}}\right\rangle$
+$=-\sum\limits_{r=1}^{m}\eta\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\sum\limits_{t=1}^{n_{1}}s_{t}(k)\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}+\sum\limits_{j=1}^{n_{2}}h_{j}(k)\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\right\rangle$
+$=-\eta\left[\sum\limits_{t=1}^{n_{1}}\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}\right\rangle s_{t}(k)+\sum\limits_{j=1}^{n_{2}}\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\right\rangle h_{j}(k)\right]$
+$=-\eta[\bm{H}(k)]_{p}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix},$
 where $[\bm{H}(k)]_{p}$ denotes the $p$-row of $\bm{H}(k)$.
 Similarly, for $h(k)$, we have
-$h_{j}(k+1)-h_{j}(k)$ $=\left[h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (65)
+$h_{j}(k+1)-h_{j}(k)$ $=\left[h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle\right]+\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle$ (45)
 $:=I_{1}^{n_{1}+j}(k)+I_{2}^{n_{1}+j}(k)$
 and
-$I_{2}^{n_{1}+j}(k)=-\eta[\bm{H}(k)]_{n_{1}+j}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$ (66)
+$I_{2}^{n_{1}+j}(k)=-\eta[\bm{H}(k)]_{n_{1}+j}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}.$ (46)
 Combining (42), (43), (44) and (45) yields that
-$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}$ $=\bm{I}_{1}(k)+\bm{I}_{2}(k)$
-$=\bm{I}_{1}(k)-\eta\bm{H}(k)\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$
+$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}-\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}$ $=\bm{I}_{1}(k)+\bm{I}_{2}(k)$
+$=\bm{I}_{1}(k)-\eta\bm{H}(k)\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}.$
 A simple transformation directly leads to
-$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
-7.4. Proof of Theorem 2
-Proof.
-The proof strategy is similar to that for Theorem 1. Our induction hypothesis is the following boundedness of the hidden weights and convergence rate of the empirical loss.
+$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k),$
+which is exactly (39), the new recursion formula we need to prove. ∎
+9.4. Proof of Theorem 3.7
+Similar to [10] and [21], we prove Theorem 3.7 by induction. Our induction hypothesis is the following convergence rate of the empirical loss and upper bounds for the weights.
 Condition 2.
@@
 $L(t)\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{t}L(0),$ (47)
-where $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1$.
-From (62), we know that with probability at least $1-\delta$, $\|\bm{w}_{r}(0)\|_{2}\leq\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}$ holds for any $r\in[m]$.
-Thus, if we can prove that $\bm{w}_{r}(t)$ is closed enough to $\bm{w}_{r}(0)$, then $\|\bm{w}_{r}(t)\|_{2}\leq B$.
-Corollary 2 (Lemma 4.1 in [19]).
+where $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1$ and $L(k)$ is an abbreviation of $L(\bm{w}(k))$.
+From the update formula of gradient descent, we can directly derive the following corollary, which indicates that under over-parameterization, the weights are closed to their initializations.
+Corollary 9.2.
 If Condition 2 holds for $t=0,\cdots,k$, then we have for every $r\in[m]$,
@@
 where $C$ is a universal constant.
-Corollary 2 implies that when $m$ is large enough, we have $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq 1$ and then $\|\bm{w}_{r}(k+1)\|_{2}\leq B$.
+Proof Sketch: Assume that Condition 2 holds for $t=0,\cdots,k$, it suffices to demonstrate that Condition 2 also holds for $t=k+1.$
+From the recursion formula (39), we have that
+$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$ (49)
+$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k)\right\|_{2}^{2}$
+$\leq\left\|\bm{I}-\eta\bm{H}(k)\right\|_{2}^{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\left\|\bm{I}_{1}(k)\right\|_{2}^{2}+2\left\|\bm{I}-\eta\bm{H}(k)\right\|_{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}\left\|\bm{I}_{1}(k)\right\|_{2},$
+where the inequality follows from the Cauchy’s inequality.
+Combining Corollary B.2 with Lemma 3.5, we can deduce that when $m$ is large enough, we have $\|\bm{H}(k)-\bm{H}(0)\|_{2}\leq\lambda_{0}/4$.
+Thus, $\lambda_{min}(\bm{H}(k))\geq\lambda_{0}/2$ and $\bm{I}-\eta\bm{H}(k)$ is positive definite when $\eta=\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$.
+On the other hand, with Corollary B.2, we can derive that $\|\bm{I}_{1}(k)\|_{2}=\mathcal{O}(\eta\sqrt{L}(k)/\sqrt{m})$.
+Plugging these results into (48), we have
+$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$ (50)
+$=\left(\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}+\mathcal{O}\left(\frac{\eta^{2}}{m}\right)+\mathcal{O}\left(\frac{\eta}{\sqrt{m}}\right)\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}$
+$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2},$
+where the last inequality holds when $m$ is large enough.
+Now we come to prove Theorem 3.7.
+Proof.
+Corollary B.2 implies that when $m$ is large enough, we have $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq 1$ and then $\|\bm{w}_{r}(k+1)\|_{2}\leq B$.
 Thus, in induction, we only need to prove that (46) also holds for $t=k+1$, which relies on the recursion formula (49).
-Recall that
-$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
-From Corollary 2 and Lemma 5, taking $CM^{2}R<\frac{\lambda_{0}}{4}$ in (24) and $R^{{}^{\prime}}\leq R$ in (68) yields that $\lambda_{min}(\bm{H}(k))\geq\frac{\lambda_{0}}{2}$ and
+Recall that the recursion formula is
+$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k).$
+From Corollary B.2 and Lemma 3.5, taking $CM^{2}R<\frac{\lambda_{0}}{4}$ in (8) and $R^{{}^{\prime}}\leq R$ in (47) yields that $\lambda_{min}(\bm{H}(k))\geq\lambda_{min}(\bm{H}(0))-\frac{\lambda_{0}}{4}\geq\frac{\lambda_{0}}{2}$ and
 $\|\bm{H}(k)\|_{2}\leq\|\bm{H}(0)\|_{2}+\frac{\lambda_{0}}{4}\leq\|\bm{H}^{\infty}\|_{2}+\frac{\lambda_{0}}{2}\leq\frac{3}{2}\|\bm{H}^{\infty}\|_{2}.$
@@
 Combining these facts with the recursion formula, we have that
-$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$ (69)
-$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left\langle(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},\bm{I}_{1}(k)\right\rangle$
-$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}.$
-Thus, it remains to bound $\|\bm{I}_{1}(k)\|_{2}$.
+$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$ (51)
+$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left\langle(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix},\bm{I}_{1}(k)\right\rangle$
+$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}.$
+Thus, it remains only to bound $\|\bm{I}_{1}(k)\|_{2}$.
 For $\bm{I}_{1}(k)$, recall that $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}}(k),I_{1}^{n_{1}+1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$ and for $p\in[n_{1}]$,
-$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle,$
+$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle,$
 for $j\in[n_{2}]$,
-$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$
+$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle.$
 Recall that
-$s_{p}(k)=\frac{1}{\sqrt{n_{1}}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}x_{p})w_{r0}(k)-a_{r}\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}x_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}-f(x_{p})\right)$
+$s_{p}(k)=\frac{1}{\sqrt{n_{1}}}\left(\frac{1}{\sqrt{m}}\left(\sum\limits_{r=1}^{m}a_{r}\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}x_{p})w_{r0}(k)-a_{r}\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}x_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\right)-f(x_{p})\right)$
 and
-$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+2}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}$
-$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}(k)\end{pmatrix}].$
+$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+2}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}\right.$
+$\quad\left.-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}(k)\end{pmatrix}\right].$
 Define $\chi_{pr}^{1}(k):=\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$ and $\chi_{pr}^{2}(k):=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$, i.e., $\chi_{pr}^{1}(k)$ and $\chi_{pr}^{2}(k)$ are related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ respectively.
 Then define
-$\hat{\chi}_{pr}^{1}(k)=\chi_{pr}^{1}(k+1)-\chi_{pr}^{1}(k)-\langle\frac{\partial\chi_{pr}^{1}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle$
+$\hat{\chi}_{pr}^{1}(k)=\chi_{pr}^{1}(k+1)-\chi_{pr}^{1}(k)-\left\langle\frac{\partial\chi_{pr}^{1}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\right\rangle$
 and
-$\hat{\chi}_{pr}^{2}(k)=\chi_{pr}^{2}(k+1)-\chi_{pr}^{2}(k)-\langle\frac{\partial\chi_{pr}^{2}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle.$
+$\hat{\chi}_{pr}^{2}(k)=\chi_{pr}^{2}(k+1)-\chi_{pr}^{2}(k)-\left\langle\frac{\partial\chi_{pr}^{2}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\right\rangle.$
 At this time, we have
 $I_{1}^{p}(k)=\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}a_{r}\left[\hat{\chi}_{pr}^{1}(k)-\hat{\chi}_{pr}^{2}(k)\right].$
-The purpose of defining them in this way is to enable us to handle the terms related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ separately.
-Similar to that for regression problems with ReLU activation function, we first define some events. For $p\in[n_{1}]$, define
+The purpose of defining $\hat{\chi}_{pr}^{1}(k)$ and $\hat{\chi}_{pr}^{1}(k)$ in this way is to enable us to handle the terms related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ separately.
+We first recall some definitions. For $p\in[n_{1}]$,
 $A_{p,r}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}\}$
-and let $S_{p}^{\perp}=[n_{1}]\textbackslash S_{p}$.
+and $S_{p}=\{r\in[m]:I\{A_{p,r}=0\}\}$, $S_{p}^{\perp}=[n_{1}]\backslash S_{p}$.
 In the following, we are going to show that $|\hat{\chi}_{pr}^{1}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for every $r\in[m]$ and $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for $r\in S_{p}$, $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2})$ for $r\in S_{p}^{\perp}$.
-Thus, we can prove that $\|\bm{I}_{1}\|_{2}=\mathcal{O}\left(\frac{\sqrt{L(k)}}{\sqrt{m}}\right)$.
-Then combining with (69) leads to a milder requirement for $\eta$.
-For $\hat{\chi}_{pr}^{1}(k)$, we have that
+Thus, we can prove that $\|\bm{I}_{1}(k)\|_{2}=\mathcal{O}\left(\frac{\sqrt{L(k)}}{\sqrt{m}}\right)$.
+Then combining with (69) leads to the conclusion.
+For $\hat{\chi}_{pr}^{1}(k)$, from its definition, we have that
 $\hat{\chi}_{pr}^{1}(k)$ $=\sigma^{{}^{\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})w_{r0}(k+1)-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$
@@
 $\quad+\left[\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})(w_{r0}(k+1)-w_{r0}(k))\right].$
-It implies that
+This implies that
 $|\hat{\chi}_{pr}^{1}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}.$
@@
 $\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p}).$
-For $r\in S_{p}$, as $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq R,\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R$, we have that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$, which yields that
+For $r\in S_{p}$, since $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq R,\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R$, we have that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$, which yields that
 $\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})$ (55)
@@
 Combining (53), (54) and (55), we can deduce that for $r\in S_{p}$,
-$|\hat{\chi}_{pr}^{2}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2},$
-for $r\in S_{p}^{\perp}$,
+$|\hat{\chi}_{pr}^{2}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}$
+and for $r\in S_{p}^{\perp}$,
 $|\hat{\chi}_{pr}^{2}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}+B^{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
@@
 For $j\in[n_{2}]$, we consider $I_{1}^{n_{1}+j}(k)$, which can be written as follows.
-$I_{1}^{n_{1}+j}(k)$ $=h_{j}(k+1)-h_{j}(k)-\langle\bm{w}(k+1)-\bm{w}(k),\frac{\partial h_{j}(k)}{\partial\bm{w}}\rangle$
+$I_{1}^{n_{1}+j}(k)$ $=h_{j}(k+1)-h_{j}(k)-\left\langle\bm{w}(k+1)-\bm{w}(k),\frac{\partial h_{j}(k)}{\partial\bm{w}}\right\rangle$
 $=\sum\limits_{r=1}^{m}\frac{a_{r}}{\sqrt{n_{2}m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{y}_{j})-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})\right].$
@@
 $\leq\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{2}m}}.$
-Recall that
-$A_{pr}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}^{T}\bm{x}_{p}\geq 0\}\}$
-and
-$P(A_{p,r})\leq\frac{R}{\sqrt{2\pi}},\ S_{p}=\{r\in[m]:I\{A_{p,r}\}=0\}.$
+Note that
+$P(A_{p,r})\leq\frac{2R}{\sqrt{2\pi}},\ S_{p}=\{r\in[m]:I\{A_{p,r}\}=0\}.$
 Thus, from Bernstein’s inequality, we have that with probability at least $1-e^{-mR}$,
 $\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{p}^{\perp}\}=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{pr}\}\lesssim 4R.$
-Then the inequality holds for all $p\in[n_{1}]$ with probability at least $1-n_{1}e^{-mR}$. Thus from (78), we have that for every $p\in[n_{1}]$
+Then the inequality holds for all $p\in[n_{1}]$ with probability at least $1-n_{1}e^{-mR}$. Plugging this into (59), we can conclude that for every $p\in[n_{1}]$
 $|I_{1}^{p}(k)|\lesssim\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{1}m}}+\frac{\eta B^{4}\sqrt{L(k)}}{\sqrt{n_{1}}}R.$ (62)
@@
 $\lesssim\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{m}}+\eta B^{4}\sqrt{L(k)}R.$
-Plugging it into (69) yields that
-$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$
-$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}$
-$\leq[(1-\frac{\eta\lambda_{0}}{2})^{2}+C^{2}(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)^{2}+2C(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)]\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}$
-$\leq(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2},$
-where the last inequality requires that
+Plugging this into (50) yields that
+$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$
+$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}$
+$\leq\left[\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}+C^{2}\left(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R\right)^{2}+2C\left(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R\right)\right]\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}$
+$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2},$
+where $C$ is a universal constant and the last inequality requires that
 $\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}\lesssim\eta\lambda_{0},\ \eta B^{4}R\lesssim\eta\lambda_{0}.$
-Recall that we also require that $CM^{2}R<\frac{\lambda_{0}}{4}$ and
-$R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}<R.$
-Finally, with $R=\mathcal{O}(\frac{\lambda_{0}}{B^{4}})$ and Lemma 11 for the upper bound of $L(0)$, $m$ needs to satisfies that
-$m=\Omega\left(\frac{B^{12}L(0)}{\lambda_{0}^{4}}\right)=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
+Recall that we also require $CM^{2}R<\frac{\lambda_{0}}{4}$ for $R$ in (8) and
+$R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}<R$
+for $R^{{}^{\prime}}$ in (47) to make sure $\|\bm{H}(k)-\bm{H}(0)\|_{2}\leq\frac{\lambda_{0}}{4}$.
+Finally, with $R=\mathcal{O}(\frac{\lambda_{0}}{M^{2}})$ and the upper bound of $L(0)$, $m$ needs to satisfies that
+$m=\Omega\left(\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right)=\Omega\left(\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
 10. Proof of Section 4
-8.1. Proof of Lemma 7
+10.1. Proof of Lemma 4.4
 Proof.
 Recall that
-$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}x_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}$
-$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}]$
+$\bm{H}(\bm{w})=\bm{D}^{T}\bm{D},\quad\bm{D}=\left[\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right],$
+and $\bm{H}^{\infty}=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\bm{H}(\bm{w})$.
+We denote $\varphi(\bm{x};\bm{w})=\sigma^{{}^{\prime}}(\bm{w}^{T}\bm{x})w_{0}-\sigma^{{}^{\prime\prime}}(\bm{w}^{T}\bm{x})\|\bm{w}_{1}\|_{2}^{2}$, where $\bm{w}=(w_{0},\bm{w}_{1}^{T})^{T}$, $w_{0}\in\mathbb{R},\bm{w}_{1}\in\mathbb{R}^{d}$, then
+$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{1}{\sqrt{n}_{1}}\frac{a_{r}}{\sqrt{m}}\frac{\partial\varphi(\bm{x}_{p};\bm{w}_{r})}{\partial\bm{w}_{r}}.$
+Similarly, we denote $\psi(\bm{y};\bm{w})=\sigma(\bm{w}^{T}\bm{y})$, then
+$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{1}{\sqrt{n}_{2}}\frac{a_{r}}{\sqrt{m}}\frac{\partial\psi(\bm{y}_{j},\bm{w}_{r})}{\partial\bm{w}_{r}}.$
+With the notations, we can deduce that
+$H_{p,j}^{\infty}=\left\{\begin{aligned} \frac{1}{n_{1}}\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left\langle\frac{\partial\varphi(\bm{x}_{p};\bm{w})}{\partial\bm{w}},\frac{\partial\varphi(\bm{x}_{j};\bm{w})}{\partial\bm{w}}\right\rangle&,&1\leq p\leq n_{1},1\leq j\leq n_{1},\\ \frac{1}{\sqrt{n_{1}n_{2}}}\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left\langle\frac{\partial\varphi(\bm{x}_{p};\bm{w})}{\partial\bm{w}},\frac{\partial\psi(\bm{y}_{j};\bm{w})}{\partial\bm{w}}\right\rangle&,&1\leq p\leq n_{1},n_{1}+1\leq j\leq n_{1}+n_{2},\\ \frac{1}{n_{2}}\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left\langle\frac{\partial\psi(\bm{y}_{p};\bm{w})}{\partial\bm{w}},\frac{\partial\psi(\bm{y}_{j};\bm{w})}{\partial\bm{w}}\right\rangle&,&n_{1}+1\leq p\leq n_{1}+n_{2},n_{1}+1\leq j\leq n_{1}+n_{2},\end{aligned}\right.$
+where $H_{p,j}^{\infty}$ is the $(p,j)$-th entry of $\bm{H}^{\infty}$.
+The proof of this lemma requires tools from functional analysis.
+Let $\mathcal{H}$ be a Hilbert space of integrable $(d+2)$-dimensional vector fields on $\mathbb{R}^{d+2}$, i.e., $f\in\mathcal{H}$ if $\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}[\|f(\bm{w})\|_{2}^{2}]<\infty$.
+The inner product for any two elements $f,g$ in $\mathcal{H}$ is $\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}[\langle f(\bm{w}),g(\bm{w})\rangle]$.
+Thus, proving $\bm{H}^{\infty}$ is strictly positive definite is equivalent to show that
+$\frac{\partial\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}},\frac{\partial\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}}\in\mathcal{H}$
+are linearly independent. Suppose that there are $\alpha_{1},\cdots,\alpha_{n_{1}},\beta_{1},\cdots,\beta_{n_{2}}\in\mathbb{R}$ such that
+$\alpha_{1}\frac{\partial\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\alpha_{n_{1}}\frac{\partial\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}}+\beta_{1}\frac{\partial\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\beta_{n_{2}}\frac{\partial\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}}=0\ in\ \mathcal{H}.$
+This implies that
+$\alpha_{1}\frac{\partial\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\alpha_{n_{1}}\frac{\partial\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}}+\beta_{1}\frac{\partial\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\beta_{n_{2}}\frac{\partial\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}}=0$ (63)
+holds for all $\bm{w}\in\mathbb{R}^{d+1}$, as $\sigma(\cdot)$ is smooth.
+First, we compute the derivatives of $\varphi$ and $\psi$ with respect to $\bm{w}$. For the $k$-th derivative of $\psi(\bm{y};\bm{w})$ with respect to $\bm{w}$, we have
+$\frac{\partial^{k}\psi(\bm{y};\bm{w})}{\partial\bm{w}^{k}}=\sigma^{(k)}(\bm{w}^{T}\bm{y})\bm{y}^{\otimes(k)},$
+where $\otimes$ denotes the tensor product.
+For $\varphi(\bm{x};\bm{w})$, let $\varphi_{0}(\bm{x};\bm{w})=\sigma^{{}^{\prime}}(\bm{w}^{T}\bm{x})w_{0}$ and $\varphi_{i}(\bm{x};\bm{w})=\sigma^{{}^{\prime\prime}}(\bm{w}^{T}\bm{x})w_{i}^{2}$ for $1\leq i\leq d$.
+Then
+$\varphi(\bm{x};\bm{w})=\varphi_{0}(\bm{x};\bm{w})-\sum\limits_{i=1}^{d}\varphi_{i}(\bm{x};\bm{w}).$
+For the $k$-th derivative of $\varphi_{0}(\bm{x};\bm{w})$ with respect to $\bm{w}$, analogous to the Leibniz rule for the $k$-th derivative of the product of two scalar functions, we have
+$\frac{\partial^{k}\varphi_{0}(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ $=\sigma^{(k+1)}(\bm{w}^{T}\bm{x})w_{0}\bm{x}^{\otimes(k)}+\sum\limits_{i=1}^{k}\bm{x}^{\otimes(i-1)}\otimes\bm{e}_{0}\otimes\bm{x}^{\otimes(k-i)}\sigma^{(k)}(\bm{w}^{T}\bm{x})$ (64)
+$=\sigma^{(k+1)}(\bm{w}^{T}\bm{x})w_{0}\bm{x}^{\otimes(k)}+\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}^{\otimes(k)},$
+where $\bm{e}_{0}=(1,0,\cdots,0)^{T}\in\mathbb{R}^{d+1}$ and in the second equality, $\bm{e}_{0}^{(i)}$ denotes that $\bm{e}_{0}$ is placed at the $i$-th position.
+Similarly, for $\varphi_{i}(\bm{x};\bm{w})$ where $1\leq i\leq d$, taking $i=1$ as example, we have
+$\frac{\partial^{k}\varphi_{1}(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ $=\sigma^{(k+2)}(\bm{w}^{T}\bm{x})w_{1}^{2}\bm{x}^{\otimes(k)}+2kw_{1}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{1}^{(i)}\otimes\bm{x}^{\otimes(k-1)}$ (65)
+$\quad+k(k-1)\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{1\leq i<j\leq k}\bm{e}_{1}^{(i)}\otimes\bm{e}_{1}^{(j)}\otimes\bm{x}^{\otimes(k-2)},$
+where $\bm{e}_{i}\in\mathbb{R}^{d+1}$ is a vector with the $(i+1)$-th component equal to $1$ and all other components equal to 00, and $\bm{e}_{1}^{(i)}$ indicates that $\bm{e}_{1}$ is placed at the $i$-th position.
+By combining the derivatives of $\varphi_{0}(\bm{x};\bm{w}),\cdots,\varphi_{d}(\bm{x};\bm{w})$ from equations (64) and (65), we can compute the $k$-th derivative of $w_{0}\sigma^{{}^{\prime}}(\bm{w}^{T}\bm{x})-\sum\limits_{i=1}^{d}w_{i}^{2}\sigma^{{}^{\prime\prime}}(\bm{w}^{T}\bm{x})$ as follows:
+$\frac{\partial^{k}\varphi(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ $=\frac{\partial^{k}\varphi_{0}(x;\bm{w})}{\partial\bm{w}^{k}}-\sum\limits_{i=1}^{d}\frac{\partial^{k}\varphi_{i}(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ (66)
+$=w_{0}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\bm{x}^{\otimes(k)}+\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}^{\otimes(k)}$
+$\quad-\sum\limits_{t=1}^{d}\left[w_{t}^{2}\sigma^{(k+2)}(\bm{w}^{T}\bm{x})\bm{x}^{\otimes(k)}+2kw_{t}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{t}^{(i)}\otimes\bm{x}^{\otimes(k-1)}\right.$
+$\quad\left.+k(k-1)\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{1\leq i<j\leq k}\bm{e}_{t}^{(i)}\otimes\bm{e}_{t}^{(j)}\otimes\bm{x}^{\otimes(k-2)}\right].$
+Note that when any two points in $\{\bm{x}_{1},\cdots,\bm{x}_{n_{1}},\bm{y}_{1},\cdots,\bm{y}_{n_{2}}\}$ are non-parallel, the tensors
+$\bm{x}_{1}^{\otimes(n_{1}+n_{2})},\cdots,\bm{x}_{n_{1}}^{\otimes(n_{1}+n_{2})},\bm{y}_{1}^{\otimes(n_{1}+n_{2})},\cdots,\bm{y}_{n_{2}}^{\otimes(n_{1}+n_{2})}$
+are linearly independent (see Lemma G.6 in [11]). This observation motivates us to take the $(k-1)$-th derivative of both sides of equation (63) with respect to $\bm{w}$, yielding
+$\alpha_{1}\frac{\partial^{k}\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}^{k}}+\cdots+\alpha_{n_{1}}\frac{\partial^{k}\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}^{k}}+\beta_{1}\frac{\partial^{k}\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}^{k}}+\cdots+\beta_{n_{2}}\frac{\partial^{k}\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}^{k}}=0.$ (67)
+Since this equation holds for all $\bm{w}\in\mathbb{R}^{d+1}$, we specifically consider $\bm{w}=(w_{0},\bm{0}_{d})$, where $w_{0}$ is to be determined. Under this condition, equation (67) becomes
+$w_{0}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k+1)}(w_{0}x_{p}^{0})\bm{x}_{p}^{\otimes(k)}\right]+\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k)}(w_{0}x_{p}^{0})\bm{z}_{p}\right]+\sum\limits_{j=1}^{n_{2}}\beta_{j}\left[\sigma^{(k)}(w_{0}y_{j}^{0})\bm{y}_{j}^{\otimes(k)}\right]=0,$ (68)
+where the tensor $\bm{z}_{p}$ is defined as
+$\bm{z}_{p}=\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}_{p}^{\otimes(k)}-k(k-1)\sum\limits_{t=1}^{d}\sum\limits_{1\leq i<j\leq k}\bm{e}_{t}^{(i)}\otimes\bm{e}_{t}^{(j)}\otimes\bm{x}_{p}^{\otimes(k-2)}.$ (69)
+By assumption, for any positive integer $n\geq 0$ we have $\lim\limits_{x\to+\infty}\frac{\sigma^{(n)}(x)}{\phi(x)}=c_{n}\neq 0$.
+We first consider the case where all input components satisfy $x_{1}^{0}=\cdots=x_{n_{1}}^{0}=y_{1}^{0}=\cdots=y_{n_{2}}^{0}=a>0$.
+Under this condition, equation (68) simplifies to:
+$w_{0}\sigma^{(k+1)}(w_{0}a)\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}\right]+\sigma^{(k)}(w_{0}a)\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{z}_{p}\right]+\sigma^{(k)}(w_{0}a)\left[\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}\right]=0.$ (70)
+Dividing both sides of equation (70) by $\phi(w_{0}a)$ yields
+$w_{0}\frac{\sigma^{(k+1)}(w_{0}a)}{\phi(w_{0}a)}\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}\right]+\frac{\sigma^{(k)}(w_{0}a)}{\phi(w_{0}a)}\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{z}_{p}\right]+\frac{\sigma^{(k)}(w_{0}a)}{\phi(w_{0}a)}\left[\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}\right]=0.$ (71)
+Now taking the limit as $w_{0}$ tends to positive infinity, we observe that $\frac{\sigma^{(k)}(w_{0}a)}{\phi(w_{0}a)}$ converges to a non-zero constant, while $w_{0}\frac{\sigma^{(k+1)}(w_{0}a)}{\phi(w_{0}a)}$ diverges to infinity.
+This asymptotic behavior leads to the following conclusions:
+$\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}=0,\ \sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{z}_{p}+\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}=0.$ (72)
+By the linear independence of the tensor products (established earlier), we can deduce that $\alpha_{p}=0$ for all $p=1,\cdots,n_{1}$, which subsequently implies $\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}=0$ and thus $\beta_{j}=0$ for all $j=1,\cdots,n_{2}$.
+When our previous assumption does not hold—that is, when $x_{1}^{0},\cdots,x_{n_{1}}^{0},y_{1}^{0},\cdots,y_{n_{2}}^{0}$ are not necessarily all equal—we proceed as follows:
+Case 1: All elements are strictly positive.
+Let $b=\min\{x_{1}^{0},\cdots,x_{n_{1}}^{0},y_{1}^{0},\cdots,y_{n_{2}}^{0}\}$. Dividing both sides of (68) by $\phi(w_{0}b)$, we observe that for any $x>b$,
+$\lim\limits_{w_{0}\to+\infty}\frac{\sigma^{(n)}(w_{0}x)}{\phi(w_{0}b)}=0.$
+Thus, the problem reduces to the previously considered case where all inputs are equal (to $b$).
+Due to linear independence, the coefficients $\alpha_{p}$ and $\beta_{j}$ corresponding to the minimal $b$ must vanish.
+Repeating this process iteratively, we conclude that all $\alpha_{p}$ and $\beta_{j}$ must be zero.
+Case 2: Some elements are zero.
+Since the $\bm{x}_{p}$ are interior points (and thus non-zero), any zero-valued inputs must correspond to boundary or initial conditions.
+Let us assume without loss of generality that $y_{1}^{0},\cdots,y_{n_{2}}^{0}$ are all zero.
+1. If $\sigma^{(k)}(0)=0$, , then following our previous method, we can conclude that the coefficients corresponding to non-zero inputs vanish. Returning to equation (67), we have
+$\sum\limits_{j=1}^{n_{2}}\beta_{j}\sigma^{(k)}(\bm{w}^{T}\bm{y}_{j})\bm{y}_{j}^{\otimes(k)}=0.$ (73)
+By the independence of $\bm{y}_{1},\cdots,\bm{y}_{n_{2}}$, we can deduce that $\beta_{j}\sigma^{(k)}(\bm{w}^{T}\bm{y}_{j})=0$ holds for all $j\in[n_{2}]$.
+From the assumption, we can select $\bm{w}$ such that $\sigma^{(k)}(\bm{w}^{T}\bm{y}_{j})\neq 0$, and consequently, $\beta_{j}=0$ holds for all $j\in[n_{2}]$.
+2.
+If $\sigma^{(k)}(0)\neq 0$, let $b$ be the smallest strictly positive value among $x_{1}^{0},\cdots,x_{n_{1}}^{0},y_{1}^{0},\cdots,y_{n_{2}}^{0}$.
+Divide (70) by $\phi(w_{0}b/2)$.
+Since all other positive terms decay to zero as $w_{0}\to+\infty$, we obtain:
+$\lim\limits_{w_{0}\to\infty}\sum\limits_{j=1}^{n_{2}}\beta_{j}\sigma^{(k)}(0)\bm{y}_{j}^{\otimes(k)}/\phi(w_{0}b/2)=0.$ (74)
+This implies that $\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}=0$. By linear independence, all $\beta_{j}=0$. ∎
+10.2. Proof of Lemma 4.5
+Proof.
+Recall that
+$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}\right.$
+$\quad\left.-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}\right]$
 and
 $\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j}.$
-(1) For $p\in[n_{1}]$, define the event
-$A_{p,r}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}\}.$
-Similar to (44), we can deduce that for any $p\in[n_{1}]$,
-$P(A_{p,r})\leq\frac{2R}{\sqrt{2\pi}}.$
+(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function.
 From the form of $\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$, we can deduce that
@@
 $\lesssim\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+|I\{\bm{w}_{r}^{T}\bm{x}_{p}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}|(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right]$
-$\leq\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right],$
-where the second inequality follows from the fact $\|\bm{w}-\bm{w}_{r}(0)\|_{2}<R\leq 1$ and the definition of $A_{p,r}$.
+$\leq\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{pr}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right],$
+where the second inequality follows from the fact $\|\bm{w}-\bm{w}_{r}(0)\|_{2}<R\leq 1$ and the definition of $A_{pr}$ in (26).
 Similarly, we have that
@@
 $=\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}\right)$
-$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}(R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1))^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
-$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
+$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}\left(R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{pr}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right)^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
+$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{pr}\}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
 $=\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)$
-$\quad+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}\left(\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}+1\right)$
-$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{M^{2}}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\},$
+$\quad+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{pr}\}\left(\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}+1\right)$
+$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{M^{2}}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{pr}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\},$
 where $M=2(d+2)\log(2m(d+2)/\delta)$. Note that from (47), we have
@@
 On the other hand, applying Bernstein’s inequality yields that with probability at least $1-n_{1}e^{-mR}$,
-$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{p,r}\}<4R$
+$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{pr}\}<4R$
 holds for all $p\in[n_{1}]$.
@@
 holds with probability at least $1-\delta-n_{1}e^{-mR}$.
-(2) Note that when $\sigma$ satisfies the assumption 5, $\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}$ are all Lipschitz continuous. Thus we can obtain that
+(2) Note that when $\sigma$ satisfies Assumption 4.3, $\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}$ are all Lipschitz continuous and bounded.
+Thus, we can obtain that
 $\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}R(\|\bm{w}_{r}(0)\|_{2}^{2}+\|\bm{w}_{r}(0)\|_{2}+1)\lesssim\frac{1}{\sqrt{n_{1}m}}R(\|\bm{w}_{r}(0)\|_{2}^{2}+1),$ (77)
@@
 $\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
-$\lesssim R^{2}\left[d^{2}+\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2}\right],$
-where the last inequality follows from the fact that $\|\|\bm{w}_{r}(0)\|_{2}^{4}\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}$ and Lemma 8. ∎
-8.2. Proof of Theorem 3
-We prove Theorem 3 by induction. Our induction is the following condition for the hidden weights.
+$\lesssim R^{2}\left[d^{2}+\frac{d^{2}}{\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{m}\left(\log\left(\frac{1}{\delta}\right)\right)^{2}\right],$
+where the last inequality follows from the fact that $\left\|\|\bm{w}_{r}(0)\|_{2}^{4}\right\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}$ and Lemma D.4. ∎
+10.3. Proof of Theorem 4.7
+For the sake of completeness in the proof, we restate Condition 1 and Corollary 4.11 from the main text, and label them as Condition 3 and Corollary C.1, respectively.
 Condition 3.
@@
 for all $r\in[m]$, where $C$ is a universal constant and $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1.$
-Instead of inducing on the convergence rate of the empirical loss function, as shown in Condition 2, we perform induction on the movements of the hidden weights.
-Because with Condition 3, we can directly derive the following convergence rate of the empirical loss function.
-Corollary 3.
-If Condition 3 holds for $t=0,\cdots,k$ and $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$, then
+Corollary 10.1.
+If Condition 3 holds for $t=0,\cdots,k$ and $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, then
 $L(t)\leq(1-\eta)^{t}L(0),$
-holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 7 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ in (38) when $\sigma$ satisfies Asssumption 5.
-Thanks to Corollary 3, it is sufficient to prove that Condition 3 also holds for $t=k+1$.
-For readability, we defer the proof of Corollary 3 to the end of this section.
-In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary 3 leads to Theorem 3.
-Proof of Theorem 3.
-We let $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function and let $R^{{}^{\prime\prime}}=CdR$ in (37) when $\sigma$ satisfies Assumption 5.
-First, set $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$, then from Lemma 7 we have $\|\bm{J}(t)-\bm{J}(0)\|_{2}\leq\frac{\sqrt{\lambda_{0}}}{3}$, thus $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and $\lambda_{min}(\bm{H}(t))\geq\frac{4\lambda_{0}}{9}$ for $t=0,\cdots,k$.
-From the updating role of NGD, we have
+holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 4.5 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (16) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ in (18) when $\sigma$ satisfies Assumption 4.3.
+Thanks to Corollary C.1, it is sufficient to prove that Condition 3 also holds for $t=k+1$.
+For readability, we defer the proof of Corollary C.1 to the end of this section.
+In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary C.1 leads to Theorem 4.7.
+Sketch Proof of Theorem 4.7.
+First, let $\bm{u}(t)=\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}$, then from the updating formula of NGD (11), we have
+$\bm{u}(t+1)-\bm{u}(t)$ (79)
+$=\bm{u}\left(\bm{w}(t)-\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(\bm{w}(t))\right)-\bm{u}(\bm{w}(t))$
+[…]
+[…]
+[…]
+$:=\bm{I}_{1}(t)+\bm{I}_{2}(t),$
+where the second equality is from the fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
+In the proof, we assume that Condition 3 holds for $t=0,\cdots,k$.
+Then from Corollary C.1, to prove Theorem 4.7, it suffices to demonstrate that this condition also holds for $t=k+1$.
+Here, we primarily explain the process from Condition 3 to Corollary C.1, while other content is placed in the following full proof of Theorem 4.7.
+Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging this into (79) yields that
+$\bm{u}(t+1)=(1-\eta)\bm{u}(t)+\bm{I}_{2}(t).$ (80)
+From the above equation, we can see the difference between NGD and GD. Recall that the iteration formula for GD is
+$\bm{u}(t+1)=(1-\eta\bm{H}(t))\bm{u}(t)+\bm{I}_{1}(t).$
+Precisely because of this, the convergence rate of GD is inevitably influenced by $\lambda_{0}$ , whereas that of NGD is not.
+From the stability of the Jacobian matrix, we can deduce that $\|\bm{I}_{2}(t)\|_{2}=\mathcal{O}(\eta\|\bm{u}(t)\|_{2}/\sqrt{m})$. Plugging this into (80) yields that
+$\|\bm{u}(t+1)\|_{2}^{2}$ (81)
+$\leq\|(1-\eta)\bm{u}(t)\|_{2}^{2}+\|\bm{I}_{2}(t)\|_{2}^{2}+2(1-\eta)\|\bm{u}(t)\|_{2}\|\bm{I}_{2}(t)\|_{2}$
+$=\left((1-\eta)^{2}+\mathcal{O}\left(\frac{\eta^{2}}{m}\right)+2(1-\eta)\mathcal{O}\left(\frac{\eta}{\sqrt{m}}\right)\right)\|\bm{u}(t)\|_{2}^{2}$
+$\leq(1-\eta)\|\bm{u}(t)\|_{2}^{2},$
+where the last inequality holds if $m$ is large enough. ∎
+Full Proof of Theorem 4.7.
+Recall that we let $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (16) when $\sigma$ is the $\text{ReLU}^{3}$ activation function and let $R^{{}^{\prime\prime}}=CdR$ in (18) when $\sigma$ satisfies Assumption 4.3.
+First, we can set $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{3\lambda_{0}}}{6}$, since $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$.
+Then from Lemma 4.5 we have $\|\bm{J}(t)-\bm{J}(0)\|_{2}\leq\frac{\sqrt{3\lambda_{0}}}{6}$, thus
+$\sigma_{min}(\bm{J}(t))\geq\sigma_{min}(\bm{J}(0))-\|\bm{J}(t)-\bm{J}(0)\|_{2}\geq\frac{\sqrt{3\lambda_{0}}}{2}-\frac{\sqrt{3\lambda_{0}}}{6}=\frac{\sqrt{3\lambda_{0}}}{3}$
+and then $\lambda_{min}(\bm{H}(t))\geq\frac{\lambda_{0}}{3}$ for $t=0,\cdots,k$, where $\sigma_{min}(\cdot)$ denotes the least singular value.
+From the updating rule of NGD, we have
 $\bm{w}_{r}(t+1)=\bm{w}_{r}(t)-\eta\left[\bm{J}(t)^{T}\right]_{r}(\bm{H}(t))^{-1}\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix},$
 where
-$\left[\bm{J}(t)^{T}\right]_{r}=\left(\frac{\partial s_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial s_{n_{1}}(t)}{\partial\bm{w}_{r}},\frac{\partial h_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial h_{n_{2}}(t)}{\partial\bm{w}_{r}}\right).$
+$\left[\bm{J}(t)^{T}\right]_{r}=\left[\frac{\partial s_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial s_{n_{1}}(t)}{\partial\bm{w}_{r}},\frac{\partial h_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial h_{n_{2}}(t)}{\partial\bm{w}_{r}}\right].$
 Therefore, for $t=0,\cdots,k$ and any $r\in[m]$, we have
@@
 $\leq\eta\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\|\bm{H}(t)^{-1}\|_{2}\sqrt{L(t)}$
-$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\sqrt{L(t)}$
-$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{F}\sqrt{L(t)}$
-$=\frac{9\eta}{4\lambda_{0}}\sqrt{\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}}\sqrt{L(t)}$
-$\lesssim\frac{4\eta}{9\lambda_{0}}\sqrt{\frac{B^{4}+1}{m}}\sqrt{L(t)}$
+$\leq\frac{3\eta}{\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\sqrt{L(t)}$
+$\leq\frac{3\eta}{\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{F}\sqrt{L(t)}$
+$=\frac{3\eta}{\lambda_{0}}\sqrt{\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}}\sqrt{L(t)}$
+$\lesssim\frac{\eta}{\lambda_{0}}\sqrt{\frac{B^{4}+1}{m}}\sqrt{L(t)}$
 $\lesssim\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
 $\leq\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}(1-\eta)^{t/2}\sqrt{L(0)},$
-where the last inequality is due to Corollary 3.
+where the last inequality is due to Corollary C.1.
 Summing $t$ from 00 to $k$ yields that
@@
 $\leq C\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}\sum\limits_{t=0}^{k}(1-\eta)^{t/2}\sqrt{L(0)}$
-$\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}.$
-Note that although $C$ is a universal constant, it still depends on the activation function.
-Now, when $R^{{}^{\prime}}\leq 1$, we can deduce that Condition 3 also holds for $t=k+1$. Thus, it remains only to derive the requirement for $m$.
-Recall that we need $m$ to satisfy that $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$.
-(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CM\sqrt{R}\leq\frac{\sqrt{\lambda_{0}}}{3}$, wihch implies that $R\lesssim\frac{\lambda_{0}}{M^{2}}$.
-Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ imples that
-$m=\Omega\left(\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right).$
-From Lemma 11, we can deduce that
-$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
-(2) When $\sigma$ satisfies Assumption 5, we have that
-$R\lesssim\frac{\sqrt{\lambda_{0}}}{d},R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R.$
-From Lemma 5 in [23], we know
-$L(0)\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right).$
-Thus, we can deduce that
-$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
-Proof of Corollary 3.
-Similar as before, when $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$ , we have $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and then $\lambda_{min}(\bm{H}(t))\geq\frac{4}{9}\lambda_{0}$ for $t=0,\cdots,k$.
+$\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}},$
+where $C$ is a universal constant.
+Now, when $R^{{}^{\prime}}\leq 1$, we can deduce that $\|\bm{w}_{r}(k+1)\|_{2}\leq B$, implying that Condition 3 also holds for $t=k+1$. Thus, it remains only to derive the requirement for $m$.
+Recall that we need $m$ to satisfy that $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$.
+(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, in Corollary C.1, $R^{{}^{\prime\prime}}=CM\sqrt{R}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, implying that $R\lesssim\frac{(1-\eta)\lambda_{0}}{M^{2}}$.
+Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ implies that
+$m=\Omega\left(\frac{1}{(1-\eta)^{2}}\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right).$
+From Lemma D.4 for the estimation of $L(0)$, i.e.,
+$L(0)\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
+we can deduce that
+$m=\Omega\left(\frac{1}{(1-\eta)^{2}}\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
+(2) When $\sigma$ satisfies Assumption 4.3, we have that
+$R\lesssim\frac{\sqrt{(1-\eta)\lambda_{0}}}{d},R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R.$
+From Lemma D.4, we can deduce that
+$m=\Omega\left(\frac{1}{1-\eta}\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
+Proof of Corollary C.1.
+Similar as before, when $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{3\lambda_{0}}}{6}$ , we have $\sigma_{min}(\bm{J}(t))\geq\frac{\sqrt{3\lambda_{0}}}{3}$ and then $\lambda_{min}(\bm{H}(t))\geq\frac{\lambda_{0}}{3}$ for $t=0,\cdots,k$.
 Let $\bm{u}(t)=\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}$, then
@@
 $:=\bm{I}_{1}(t)+\bm{I}_{2}(t),$
-where the second equality is from fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
-Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging it into (85) yields that
+where the second equality is from the fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
+Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging this into (83) yields that
 $\bm{u}(t+1)=(1-\eta)\bm{u}(t)+\bm{I}_{2}(t).$ (84)
@@
 […]
-$\lesssim\frac{\eta\sqrt{L(t)}}{\sqrt{\lambda_{0}}}R^{{}^{\prime\prime}}$
+$\lesssim\frac{\eta\sqrt{L(t)}}{\sqrt{\lambda_{0}}}R^{{}^{\prime\prime}},$
 where the last inequality follows from the fact that
 $\|\bm{w}_{r}(s)-\bm{w}_{r}(0)\|_{2}\leq s\|\bm{w}_{r}(t+1)-\bm{w}_{r}(0)\|_{2}+(1-s)\|\bm{w}_{r}(t)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$
-and Lemma 7.
+and Lemma 4.5.
 Plugging (85) into the recursion formula (84) yields that
@@
 $\leq(1-\eta)^{2}\|\bm{u}(t)\|_{2}^{2}+\|\bm{I}_{2}(t)\|_{2}^{2}+2(1-\eta)\|\bm{u}(t)\|_{2}\|\bm{I}_{2}(t)\|_{2}$
-$\leq\left[(1-\eta)^{2}+\frac{C^{2}\eta^{2}(R^{{}^{\prime\prime}})^{2}}{\lambda_{0}}+2(1-\eta)\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\right]\|u(t)\|_{2}^{2}$
+$\leq\left[(1-\eta)^{2}+\frac{C^{2}\eta^{2}(R^{{}^{\prime\prime}})^{2}}{\lambda_{0}}+2(1-\eta)\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\right]\|\bm{u}(t)\|_{2}^{2},$
+where $C$ is a universal constant.
+Then we can choose $R^{{}^{\prime\prime}}$ such that
+$\|\bm{I}_{2}(t)\|_{2}\leq\frac{C\eta\sqrt{L(t)}R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\leq C_{1}\eta\sqrt{L(t)}=C_{1}\eta\sqrt{\bm{u}(t)},$
+where $C$ is a universal constant and $C_{1}$ is a constant to be determined.
+Thus, we can deduce that
+$\|\bm{u}(t+1)\|_{2}^{2}$ $\leq\left[(1-\eta)^{2}+(C_{1}\eta)^{2}+2(1-\eta)C_{1}\eta\right]\|\bm{u}(t)\|_{2}^{2}$
+$=\left[(1-\eta)+\eta(\eta C_{1}^{2}+2(1-\eta)C_{1}+\eta-1)\right]\|\bm{u}(t)\|_{2}^{2}$
 $\leq(1-\eta)\|\bm{u}(t)\|_{2}^{2},$
-where the last inequality follows from the requirement
-$2\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\leq\frac{\eta}{4},$
-i.e., $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$.
-Therefore, we can deduce that $\|\bm{u}(t)\|_{2}^{2}\leq(1-\eta)^{t}\|\bm{u}(0)\|_{2}^{2}$ holds for $t=0,\cdots,k$.
-8.3. Proof of Corollary 1
+where in the last inequality is due to that we can choose $C_{1}$ such that $\eta C_{1}^{2}+2(1-\eta)C_{1}+\eta-1\leq 0$.
+Note that since $\eta\in(0,1)$, the quadratic equation $\eta x^{2}+2(1-\eta)x+\eta-1=0$ has one negative root and one positive root, denoted as $x_{0}$ and $x_{1}$ respectively.
+Therefore, the condition $C_{1}\leq x_{1}$ is sufficient to satisfy the requirement.
+The explicit form of $x_{1}$ can be written as:
+$x_{1}=\frac{2(\eta-1)+\sqrt{4(1-\eta)^{2}-4\eta(\eta-1)}}{2\eta}=\frac{\sqrt{1-\eta}}{1+\sqrt{1-\eta}}\geq\frac{\sqrt{1-\eta}}{2}.$
+Thus, $C_{1}=\frac{\sqrt{1-\eta}}{2}$ is sufficient to satisfy that $\eta C_{1}^{2}+2(1-\eta)C_{1}+\eta-1\leq 0$.
+From this, we can deduce that
+$R^{{}^{\prime\prime}}\lesssim C_{1}\sqrt{\lambda_{0}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}.$
+Therefore, we can conclude that $\|\bm{u}(t)\|_{2}^{2}\leq(1-\eta)^{t}\|\bm{u}(0)\|_{2}^{2}$ holds for $t=0,\cdots,k$.
+10.4. Proof of Corollary 4.9
 Proof.
+In the proof of Theorem 4.7, we have proved that Condition 3 holds for all $t\in\mathbb{N}$. Thus, it is sufficient to prove that Condition 3 can lead to the conclusion in Corollary 4.9.
 Setting $\eta=1$ in (84) yields that
@@
 […]
-As $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)$, then for any $r\in[m]$, we have $\|\bm{w}_{r}(s)\|_{2}\leq s\|\bm{w}_{r}(t+1)\|_{2}+(1-s)\|\bm{w}_{r}(t)\|_{2}\leq B$.
-When $\sigma$ is smooth, we can deduce that for any $r\in[m]$,
+Since $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)$, then for any $r\in[m]$, we have $\|\bm{w}_{r}(s)\|_{2}\leq s\|\bm{w}_{r}(t+1)\|_{2}+(1-s)\|\bm{w}_{r}(t)\|_{2}\leq B$.
+When $\sigma(\cdot)$ is smooth, we can deduce that for any $r\in[m]$,
 $\left\|\frac{\partial s_{p}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}(B^{2}+1)\|\bm{w}_{r}(s)-\bm{w}_{r}(t)\|_{2}\leq\frac{1}{\sqrt{n_{1}m}}(B^{2}+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}$
@@
 $\lesssim B^{4}\left(\frac{B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}\right)^{2}.$
-Plugging it into (89), we have
+Plugging this into (86), we have
 […]
-$\lesssim\frac{\sqrt{L(t)}}{\sqrt{\lambda_{0}}}\frac{\eta B^{4}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
+$\lesssim\frac{\sqrt{L(t)}}{\sqrt{\lambda_{0}}}\frac{B^{4}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
 $=\frac{B^{4}}{\sqrt{m\lambda_{0}^{3}}}L(t).$
@@
 $\left\|\begin{pmatrix}\bm{s}(t+1)\\ \bm{h}(t+1)\end{pmatrix}\right\|_{2}\leq\frac{CB^{4}}{\sqrt{m\lambda_{0}^{3}}}\left\|\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}\right\|_{2}^{2}$
-holds for $t\in\mathbb{N}$, where $C$ is a universal constant. ∎
+holds for $t\in\mathbb{N}$, where $C$ is a universal constant.
+In the proof above, we only require that $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}=CdR\leq\frac{\sqrt{3\lambda_{0}}}{6}$, leading to the requirement for $m$ that
+$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
 11. Auxiliary Lemmas
-Lemma 8 (Theorem 3.1 in [24]).
-If $X_{1},\cdots,X_{n}$ are independent mean zero random variables with $\|X_{i}\|_{\psi_{\alpha}}<\infty$ for all $1\leq i\leq n$ and some $\alpha>0$, then for any vector $a=(a_{1},\cdots,a_{n})\in\mathbb{R}^{n}$, the following holds ture:
-$P(|\sum\limits_{i=1}^{n}a_{i}X_{i}|\geq 2eC(\alpha)\|b\|_{2}\sqrt{t}+2eL_{n}^{*}(\alpha)t^{1/\alpha}\|b\|_{\beta(\alpha)})\leq 2e^{-t},\ for\ all\ t\geq 0,$
+Lemma 11.1 (Theorem 3.1 in [29]).
+If $X_{1},\cdots,X_{n}$ are independent mean zero random variables with $\|X_{i}\|_{\psi_{\alpha}}<\infty$ for all $1\leq i\leq n$ and some $\alpha>0$, then for any vector $a=(a_{1},\cdots,a_{n})\in\mathbb{R}^{n}$, the following holds true:
+$P\left(\left|\sum\limits_{i=1}^{n}a_{i}X_{i}\right|\geq 2eC(\alpha)\|b\|_{2}\sqrt{t}+2eL_{n}^{*}(\alpha)t^{1/\alpha}\|b\|_{\beta(\alpha)}\right)\leq 2e^{-t},\ for\ all\ t\geq 0,$
 where $b=(a_{1}\|X_{1}\|_{\psi_{\alpha}},\cdots,a_{n}\|X_{n}\|_{\psi_{\alpha}})\in\mathbb{R}^{n}$,
 $C(\alpha):=\max\{\sqrt{2},2^{1/\alpha}\}\left\{\begin{aligned} \sqrt{8}(2\pi)^{1/4}e^{1/24}(e^{2/e}/\alpha)^{1/\alpha}&,&if\ \alpha<1,\\ 4e+2(\log 2)^{1/\alpha}&,&if\ \alpha\geq 1.\end{aligned}\right.$
-and for $\beta(\alpha)=\infty$ when $\alpha\geq 1$ and $\beta(\alpha)=\alpha/(\alpha-1)$ when $\alpha>1$,
+and for $\beta(\alpha)=\infty$ when $\alpha\leq 1$ and $\beta(\alpha)=\alpha/(\alpha-1)$ when $\alpha>1$,
 $L_{n}(\alpha):=\frac{4^{1/\alpha}}{\sqrt{2}\|b\|_{2}}\times\left\{\begin{aligned} &\|b\|_{\beta(\alpha)},&if\ \alpha<1,\\ &4e\|b\|_{\beta(\alpha)}/C(\alpha),&if\alpha\geq 1.\end{aligned}\right.$
 and $L^{*}_{n}(\alpha)=L_{n}(\alpha)C(\alpha)\|b\|_{2}/\|b\|_{\beta(\alpha)}$.
-Lemma 9 (Bernstein inequality, Theorem 3.1.7 in [25]).
+In the following, we will provide some preliminary information about Orlicz norms.
+Let $f:[0,\infty)\rightarrow[0,\infty)$ be a non-decreasing function with $f(0)=0$. The $f$-Orlicz norm of a real-valued random variable $X$ is given by
+$\|X\|_{f}:=\inf\{C>0:\mathbb{E}\left[f\left(\frac{|X|}{C}\right)\right]\leq 1\}.$
+If $\|X\|_{\psi_{\alpha}}<\infty$, we say that $X$ is sub-Weibull of order $\alpha>0$, where
+$\psi_{\alpha}(x):=e^{x^{\alpha}}-1.$
+Note that when $\alpha\geq 1$, $\|\cdot\|_{\psi_{\alpha}}$ is a norm and when $10<\alpha<$ 1, $\|\cdot\|_{\psi_{\alpha}}$ is a quasi-norm.
+Moreover, since $(|a|+|b|)^{\alpha}\leq|a|^{\alpha}+|b|^{\alpha}$ holds for any $a,b\in\mathbb{R}$ and $10<\alpha<$ 1, we can deduce that
+$\mathbb{E}e^{\frac{|X+Y|^{\alpha}}{|C|^{\alpha}}}\leq\mathbb{E}e^{\frac{|X|^{\alpha}+|Y|^{\alpha}}{|C|^{\alpha}}}=\mathbb{E}e^{\frac{|X|^{\alpha}}{|C|^{\alpha}}}e^{\frac{|Y|^{\alpha}}{|C|^{\alpha}}}\leq\left(\mathbb{E}e^{\frac{2|X|^{\alpha}}{|C|^{\alpha}}}\right)^{1/2}\left(\mathbb{E}e^{\frac{2|Y|^{\alpha}}{|C|^{\alpha}}}\right)^{1/2}.$
+This implies that
+$\|X+Y\|_{\psi_{\alpha}}\leq 2^{1/\alpha}\max\{\|X\|_{\psi_{\alpha}},\|Y\|_{\psi_{\alpha}}\}\leq 2^{1/\alpha}(\|X\|_{\psi_{\alpha}}+\|Y\|_{\psi_{\alpha}}).$
+Furthermore, for $p,q>0$, we have $\||X|\|_{\psi_{p}}=\||X|^{p/q}\|_{\psi_{q}}^{q/p}$.
+And in the related proofs, we may frequently use the fact that for real-valued random variable $X\sim\mathcal{N}(0,1)$, we have $\|X\|_{\psi_{2}}\leq\sqrt{6}$ and $\|X^{2}\|_{\psi_{1}}=\|X\|_{\psi_{2}}^{2}\leq 6$.
+Lemma 11.2.
+If $\|X\|_{\psi_{\alpha}},\|Y\|_{\psi_{\beta}}<\infty$ with $\alpha,\beta>0$, then we have $\|XY\|_{\psi_{\gamma}}\leq\|X\|_{\psi_{\alpha}}\|Y\|_{\psi_{\beta}}$, where $\gamma$ satisfies that
+$\frac{1}{\gamma}=\frac{1}{\alpha}+\frac{1}{\beta}.$
+Proof.
+Without loss of generality, we can assume that $\|X\|_{\psi_{\alpha}}=\|Y\|_{\psi_{\beta}}=1$. To prove this, let us use Young’s inequality, which states that
+$xy\leq\frac{x^{p}}{p}+\frac{y^{q}}{q},for\ x,y\geq 0,p,q>1.$
+Let $p=\alpha/\gamma,q=\beta/\gamma$, then
+$\mathbb{E}[\exp(|XY|^{\gamma})]$ $\leq\mathbb{E}\left[\exp\left(\frac{|X|^{\gamma p}}{p}+\frac{|Y|^{\gamma q}}{q}\right)\right]$
+$=\mathbb{E}\left[\exp\left(\frac{|X|^{\alpha}}{p}\right)\exp\left(\frac{|Y|^{\beta}}{q}\right)\right]$
+$\leq\mathbb{E}\left[\frac{\exp(|X|^{\alpha})}{p}+\frac{\exp(|Y|^{\beta})}{q}\right]$
+$\leq\frac{2}{p}+\frac{2}{q}$
+$=2,$
+where the first and second inequality follow from Young’s inequality. From this, we have that $\|XY\|_{\psi_{\gamma}}\leq\|X\|_{\psi_{\alpha}}\|Y\|_{\psi_{\beta}}$.
+Lemma 11.3 (Bernstein inequality, Theorem 3.1.7 in [30]).
 Let $X_{i}$, $1\leq i\leq n$ be independent centered random variables a.s.
@@
 $P\left(S_{n}\geq\sqrt{\frac{2\sigma^{2}t}{n}}+\frac{ct}{3n}\right)\leq e^{-u}.$
-Lemma 10.
-For $10<\delta<$ 1, with probability at least $1-\delta$, we have
-$\|\bm{y}-\bm{u}(0)\|_{2}^{2}=\mathcal{O}\left(n\log\left(\frac{n}{\delta}\right)\right).$
-Proof.
-First, from Cauchy’s inequality, we have
-$\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ $=\sum\limits_{i=1}^{n}\left(y_{i}-\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
-$\leq\sum\limits_{i=1}^{n}2y_{i}^{2}+2\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
-$=2\sum\limits_{i=1}^{n}y_{i}^{2}+2\sum\limits_{i=1}^{n}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}.$
-Since $\bm{w}_{r}(0)^{T}\bm{x}_{i}\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2})$ and $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$, we can deduce that
-$\|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\|_{\psi_{2}}\leq\|\bm{w}_{r}(0)^{T}\bm{x}_{i}\|_{\psi_{2}}=\mathcal{O}(1)$
-holds for all $r\in[m]$ and $i\in[n]$.
-Thus for any fixed $i\in[n]$, applying Lemma 8 yields that with probability at least $1-2e^{-t}$,
-$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right|\leq C\sqrt{t},$
-where $C$ is a universal constant. By taking union bound on all $i\in[n]$, the above inequality holds for all $i\in[n]$ with probability at least $1-2ne^{-t}$.
-Let $2ne^{-t}=\delta$, i.e., $t=\log\left(\frac{2n}{\delta}\right)$, then we have
-$\|\bm{y}-\bm{u}(0)\|_{2}^{2}\lesssim n+nt\lesssim n\log\left(\frac{n}{\delta}\right).$
-Lemma 11.
+Lemma 11.4.
 For $10<\delta<$ 1, with probability at least $1-\delta$, we have that when $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$,
-$L(0)=\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}=\mathcal{O}\left(d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
+$L(0)=\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}=\mathcal{O}\left(d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
 Proof.
@@
 $L(0)$ $=\sum\limits_{p=1}^{n_{1}}\frac{1}{2}(s_{p}(0))^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{2}(h_{j}(0))^{2}$
-$\leq\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right)+\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}f^{2}(x_{p})$
+$\leq\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right)^{2}+\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}f^{2}(x_{p})$
 $\quad+\frac{1}{n_{2}}\sum\limits_{j=1}^{n_{2}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right)^{2}+\frac{1}{n_{2}}\sum\limits_{j=1}^{n_{2}}g^{2}(\bm{y}_{j}).$
 Note that
-$|a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}$
-and $|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}.$
-Since $\|\|\bm{w}_{r}(0)\|_{2}^{3}\|_{\psi_{\frac{2}{3}}}\leq(\|\|\bm{w}_{r}(0)\|_{2}^{2}\|_{\psi_{1}})^{\frac{3}{2}}\lesssim d^{3}$, from Lemma 8, we have that for fixed $i\in[n_{1}]$ and $j\in[n_{2}]$ with probability at least $1-2e^{-t}$,
-$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}$
+$\left|a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim\|\bm{w}_{r}(0)\|_{2}^{2}|\bm{w}_{r}(0)^{T}\bm{x}_{p}|$
+and $\left|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim\|\bm{w}_{r}(0)\|_{2}^{2}|\bm{w}_{r}(0)^{T}\bm{y}_{j}|.$
+Since $\left\|\|\bm{w}_{r}(0)\|_{2}^{2}\right\|_{\psi_{1}}=\mathcal{O}(d)$ and $\|\bm{w}_{r}(0)^{T}\bm{y}_{j}\|_{\psi_{2}},\|\bm{w}_{r}(0)^{T}\bm{x}_{p}\|_{\psi_{2}}=\mathcal{O}(1)$, from Lemma D.2, we have that
+$\|\|\bm{w}_{r}(0)\|_{2}^{2}|\bm{w}_{r}(0)^{T}\bm{x}_{p}|\|_{\psi_{\frac{2}{3}}}=\mathcal{O}(d),|\bm{w}_{r}(0)^{T}\bm{y}_{j}|\|_{\psi_{\frac{2}{3}}}=\mathcal{O}(d).$
+Applying Lemma D.1 yields that for fixed $p\in[n_{1}]$ and $j\in[n_{2}]$ with probability at least $1-2e^{-t}$,
+$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim d\sqrt{t}+\frac{d}{\sqrt{m}}t^{\frac{3}{2}}$
 and with probability at least $1-2e^{-t}$,
-$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}.$
-Then taking a union bound for all $i\in[n_{1}]$ and $j\in[n_{2}]$ with $2(n_{1}+n_{2})e^{-t}=\delta$ yields that
-$L(0)$ $\lesssim(d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}})^{2}$
-$\lesssim d^{6}t+\frac{d^{6}t^{3}}{m}$
-$=d^{6}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{1}{m}\log^{3}\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
-$\lesssim d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
-as $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$.
+$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim d\sqrt{t}+\frac{d}{\sqrt{m}}t^{\frac{3}{2}}.$
+Then taking a union bound for all $p\in[n_{1}]$ and $j\in[n_{2}]$ with $2(n_{1}+n_{2})e^{-t}=\delta$ yields that
+$L(0)$ $\lesssim\left(d\sqrt{t}+\frac{d}{\sqrt{m}}t^{\frac{3}{2}}\right)^{2}$
+$\lesssim d^{2}t+\frac{d^{2}t^{3}}{m}$
+$=d^{2}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{1}{m}\log^{3}\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
+$\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
+since $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$.

### v4

@@
 Assumption 4.3 holds for various commonly used activation functions, including logistic function $\sigma(z)=1/(1+e^{-z})$ (with $\phi(z)=e^{-z}$), softplus function $\sigma(z)=\log(1+e^{z})$ (with $\phi(z)=e^{-z}$), hyperbolic tangent function $\sigma(z)=(e^{z}-e^{-z})/(e^{z}+e^{-z})$ (with $\phi(z)=e^{-2z}$), swish function $\sigma(z)=z/(1+e^{-z})$ (with $\phi(z)=ze^{-z}$) and others.
+Although $\sin(z)$ and $\cos(z)$ do not satisfy the condition (16), it can be seen from Remark 10.2 in the appendix that the Gram matrices induced by them are also strictly positive definite.
 Compared to the $\text{ReLU}^{3}$ activation function, these smooth activation functions are more popular for PINNs because solving PDEs typically requires high-order derivatives.
@@
 This implies that $\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}=0$. By linear independence, all $\beta_{j}=0$. ∎
+Remark 10.1.
+The key point in the proof lies in the fact that the order of the PDE in the interior is higher than that of the initial and boundary conditions, allowing for a natural extension to broader classes of PDEs.
+For general PDEs, we may focus solely on the interior and boundary, assuming the interior is of second order and the boundary is of first order.
+Suppose the second-order interior term is taken at $x_{0}$, i.e., it has the form $\frac{\partial^{2}u}{\partial x_{0}^{2}}$, and the first-order boundary term is also taken at $x_{0}$.
+Since we can translate the coordinates, without loss of generality, we can assume that all $x_{0}$-components are positive.
+For the interior, taking the $k$-th derivative of $w_{0}^{2}\sigma^{(2)}(\bm{w}^{T}\bm{x})$ yields that
+$w_{0}^{2}\sigma^{(k+2)}(\bm{w}^{T}\bm{x})\bm{x}^{\otimes(k)}+2kw_{0}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{1}^{(i)}\otimes\bm{x}^{\otimes(k-1)}$
+$\quad+k(k-1)\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{1\leq i<j\leq k}\bm{e}_{0}^{(i)}\otimes\bm{e}_{0}^{(j)}\otimes\bm{x}^{\otimes(k-2)}.$
+For the boundary, taking the $k$-th derivative of $w_{0}\sigma^{(1)}(\bm{w}^{T}\bm{x})$ yields that
+$\sigma^{(k+1)}(\bm{w}^{T}\bm{x})w_{0}\bm{x}^{\otimes(k)}+\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}^{\otimes(k)}.$
+As before, we set $\bm{w}=(w_{0},\bm{0})$. Then, the equation (68) becomes
+$w_{0}^{2}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k+1)}(w_{0}x_{p}^{0})\bm{x}_{p}^{\otimes(k)}\right]+w_{0}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k)}(w_{0}x_{p}^{0})\bm{z}_{p}^{0}\right]+\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k)}(w_{0}x_{p}^{0})\bm{z}_{p}^{1}\right]$
+$\quad+w_{0}\sum\limits_{j=1}^{n_{2}}\beta_{j}\left[\sigma^{(k)}(w_{0}y_{j}^{0})\bm{y}_{j}^{\otimes(k)}\right]+\sum\limits_{j=1}^{n_{2}}\beta_{j}\left[\sigma^{(k)}(w_{0}y_{j}^{0})\bm{z}_{j}^{2}\right]=0,$
+where $\bm{z}_{p}^{0},\bm{z}_{p}^{1},\bm{z}_{j}^{2}$ are tensors of similar form $\bm{z}_{p}$ in (69), whose explicit definitions are omitted for simplicity.
+Dividing both sides by $w_{0}\phi(w_{0}x_{p}^{0})$ reduces it to the form considered earlier.
+We can therefore conclude that the Gram matrix is strictly positive definite.
+Indeed, since the orders of the interior and boundary terms in the partial differential equation differ, we can relax the conditions in Lemma 4.4 to simply requiring that no two samples in $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ are parallel and no two samples in $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ are parallel.
+In brief, we can set $k=n_{1}-1$ and $k=n_{2}-1$ in equation (67), and then use the method described above.
+Remark 10.2.
+For the activation functions $\sin(x)$ and $\cos(x)$, in equation (68), we may assume that $\sigma^{(k+1)}(x)=\sin(x),\sigma^{(k)}(x)=-\cos(x)$.
+Dividing both sides of equation (68) by $w_{0}$ and letting $w_{0}\to+\infty$, we can obtain
+$\lim\limits_{w_{0}\to+\infty}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sin(w_{0}x_{p}^{0})\bm{x}_{p}^{\otimes(k)}\right]=0.$
+We express the general form of the components of the tensor above as
+$\sum\limits_{p=1}^{n_{1}}\alpha_{p}c_{p}\sin(w_{0}x_{p}^{0})=\sum\limits_{i=1}^{n}a_{i}\sin(w_{0}b_{i}),$
+where the $b_{i}>0$ are distinct and $c_{p}$ denotes the components of the tensor $\bm{x}_{p}^{\otimes(k)}$.
+For simplicity, we denote $w_{0}$ as $x$.
+To prove that $\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}=0$, we need to show that any component of this tensor is zero, i.e., its general form satisfies $\sum\limits_{p=1}^{n_{1}}\alpha_{p}c_{p}=0$.
+This is equivalent to proving $\sum\limits_{i=1}^{n}a_{i}=0$.
+Let $f(x)=\sum\limits_{i=1}^{n}a_{i}\sin(b_{i}x)$, note that dividing both sides of equation (68) by $w_{0}$ yields that $f(x)=\mathcal{O}(1/x)$.
+Thus, we can consider the average energy of $f^{2}(x)$ over the interval $[T,T+L]$, i.e.,
+[…]
+Expanding this, we obtain
+[…]
+[…]
+$=\frac{1}{2}\sum\limits_{i=1}^{n}a_{i}^{2}-\frac{1}{L}\sum\limits_{i=1}^{n}\frac{\sin(2b_{i}(T+L))-\sin(2b_{i}T)}{4b_{i}}$
+$\quad+\frac{1}{L}\sum\limits_{i\neq j}a_{i}a_{j}\frac{\sin((b_{i}-b_{j})(T+L))-\sin((b_{i}-b_{j})T)}{2(b_{i}-b_{j})}$
+$\quad-\frac{1}{L}\sum\limits_{i\neq j}a_{i}a_{j}\frac{\sin((b_{i}+b_{j})(T+L))-\sin((b_{i}+b_{j})T)}{2(b_{i}+b_{j})}.$
+Taking the limits $L\to+\infty$ and $T\to+\infty$ in the above equation, the right-hand side tends to $\frac{1}{2}\sum\limits_{i=1}^{n}a_{i}^{2}$.
+Regarding the left-hand side, recall that $f(x)=\mathcal{O}(1/x)$, thus for any $\epsilon>0$, there exists $T_{0}$ such that for all $x>T_{0}$, $|f(x)|<\epsilon$.
+Therefore, for $T>T_{0}$ and any $L$, we have
+[…]
+By the arbitrariness of $\epsilon$, we can deduce that
+[…]
+Hence, we can deduce that $\frac{1}{2}\sum\limits_{i=1}^{n}a_{i}^{2}=0$, which implies that $\sum\limits_{i=1}^{n}a_{i}=0$, i.e., $\sum\limits_{p=1}^{n_{1}}\alpha_{p}c_{p}=0$. Finally, we obtain
+$\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}=0.$
+Applying the same approach as before, we conclude that $\alpha_{p}=0$ for all $p\in[n_{1}]$and $\beta_{j}=0$ for all $j\in[n_{2}]$.
 10.2. Proof of Lemma 4.5
@@
 10.3. Proof of Theorem 4.7
-For the sake of completeness in the proof, we restate Condition 1 and Corollary 4.11 from the main text, and label them as Condition 3 and Corollary C.1, respectively.
+For the sake of completeness in the proof, we restate Condition 1 and Corollary 4.11 from the main text, and label them as Condition 3 and Corollary 10.3, respectively.
 Condition 3.
@@
 holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 4.5 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (16) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ in (18) when $\sigma$ satisfies Assumption 4.3.
-Thanks to Corollary C.1, it is sufficient to prove that Condition 3 also holds for $t=k+1$.
-For readability, we defer the proof of Corollary C.1 to the end of this section.
-In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary C.1 leads to Theorem 4.7.
+Thanks to Corollary 10.3, it is sufficient to prove that Condition 3 also holds for $t=k+1$.
+For readability, we defer the proof of Corollary 10.3 to the end of this section.
+In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary 10.3 leads to Theorem 4.7.
 Sketch Proof of Theorem 4.7.
@@
 In the proof, we assume that Condition 3 holds for $t=0,\cdots,k$.
-Then from Corollary C.1, to prove Theorem 4.7, it suffices to demonstrate that this condition also holds for $t=k+1$.
-Here, we primarily explain the process from Condition 3 to Corollary C.1, while other content is placed in the following full proof of Theorem 4.7.
+Then from Corollary 10.3, to prove Theorem 4.7, it suffices to demonstrate that this condition also holds for $t=k+1$.
+Here, we primarily explain the process from Condition 3 to Corollary 10.3, while other content is placed in the following full proof of Theorem 4.7.
 Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging this into (79) yields that
@@
 $\leq\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}(1-\eta)^{t/2}\sqrt{L(0)},$
-where the last inequality is due to Corollary C.1.
+where the last inequality is due to Corollary 10.3.
 Summing $t$ from 00 to $k$ yields that
@@
 Recall that we need $m$ to satisfy that $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$.
-(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, in Corollary C.1, $R^{{}^{\prime\prime}}=CM\sqrt{R}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, implying that $R\lesssim\frac{(1-\eta)\lambda_{0}}{M^{2}}$.
+(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, in Corollary 10.3, $R^{{}^{\prime\prime}}=CM\sqrt{R}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, implying that $R\lesssim\frac{(1-\eta)\lambda_{0}}{M^{2}}$.
 Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ implies that
@@
 $m=\Omega\left(\frac{1}{1-\eta}\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
-Proof of Corollary C.1.
+Proof of Corollary 10.3.
 Similar as before, when $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{3\lambda_{0}}}{6}$ , we have $\sigma_{min}(\bm{J}(t))\geq\frac{\sqrt{3\lambda_{0}}}{3}$ and then $\lambda_{min}(\bm{H}(t))\geq\frac{\lambda_{0}}{3}$ for $t=0,\cdots,k$.